### v1

# Bound-state confinement after trap-expansion dynamics in integrable systems

Abstract
Integrable systems possess stable families of quasiparticles, which are composite objects (bound states) of elementary excitations. Motivated by recent quantum computer experiments, we investigate bound-state transport in the spin-$1/2$ anisotropic Heisenberg chain ($XXZ$ chain). Specifically, we consider the sudden vacuum expansion of a finite region $A$ prepared in a non-equilibrium state. In the hydrodynamic regime, if interactions are strong enough, bound states remain confined in the initial region. Bound-state confinement persists until the density of unbound excitations remains finite in the bulk of $A$. Since region $A$ is finite, at asymptotically long times bound states are “liberated” after the “evaporation” of all the unbound excitations. Fingerprints of confinement are visible in the space-time profiles of local spin-projection operators. To be specific, here we focus on the expansion of the $p$-Néel states, which are obtained by repetition of a unit cell with $p$ up spins followed by $p$ down spins. Upon increasing $p$, the bound-state content is enhanced. In the limit $p\to\infty$ one obtains the domain-wall initial state. We show that for $p<4$, only bound states with $n>p$ are confined at large chain anisotropy. For $p\gtrsim 4$, also bound states with $n=p$ are confined, consistent with the absence of transport in the limit $p\to\infty$. The scenario of bound-state confinement leads to a hierarchy of timescales at which bound states of different sizes are liberated, which is also reflected in the dynamics of the von Neumann entropy.
1 Introduction
Integrable quantum many-body systems provide an ideal playground to discover exotic behaviors both at equilibrium and out-of-equilibrium. As recognized already by Bethe [1], one of the hallmarks of interacting many-body systems is the formation of multiparticle bound states. In spin chains they are formed by elementary magnon-like excitations. Bound states play a pivotal role in the equilibrium and out-of-equilibrium integrable systems, as they enter prominently in the Thermodynamic Bethe Ansatz [2] approach. Since the birth of Bethe Ansatz, bound states attracted growing attention, also experimentally [3, 4, 5, 6, 7, 8]. Furthermore, bound states are key ingredients behind exotic transport behaviors, such as superdiffusion [9]. Very recently, they have been observed in quantum-computing platforms [10].
Here we investigate the dynamics of bound states after trap-expansion experiments in the Heisenberg spin chain, which is a prototypical integrable system [2]. Our out-of-equilibrium protocol is depicted in Fig. 1: A region $A$ of size $\ell$ is embedded in an infinite chain, and is prepared in a product state. The rest of the chain is in the ferromagnetic state with all the spins down. The ferromagnetic state is the vacuum state for the Bethe Ansatz quasiparticles. The full chain then evolves under the $XXZ$ Hamiltonian. Crucially, the $XXZ$ chain, for anisotropy $\Delta\geq 1$ possesses bound states of an arbitrary number of elementary excitations. During the dynamics, as the state of region $A$ is not an eigenstate of the model, region $A$ undergoes a nontrivial evolution, whereas the remaining part of the chain does not evolve because the ferromagnetic state is eigenstate of the $XXZ$ Hamiltonian. The dynamics of region $A$ is understood in terms of the thermodynamic excitations [2] of the model. In the limit $\ell\to\infty$, at long times local observables in the bulk of region $A$ are described by a thermodynamic macrostate, which is a Generalized Gibbs Ensemble [11] ($GGE$). The initial inhomogeneity that is present at the interface between region $A$ and the rest (see Fig. 1) “melts” during the dynamics. This is due to excitations leaving region $A$, where they are initially produced. At long times and large distances from the interface, the state of the system is described by a space-time dependent $GGE$, which can be obtained within the Generalized Hydrodynamics ($GHD$) approach [12, 13]. In contrast with free-fermion systems, the thermodynamic excitations of the $XXZ$ chain are interacting. However, only elastic scattering is allowed, due to integrability. During the scattering the identity of the quasiparticles is preserved. The net effect of the scatterings is to dress the “bare” velocities of the quasiparticles. The dressed velocities depend on the space-time density of quasiparticles, and hence on the initial state. Crucially, the dressing can be so dramatic that the sign of the velocities change. Typically, this is more striking for the bound states, because their bare velocities (see section 5) are typically smaller than the magnons, implying that the bound states are more affected by scatterings. As we will show, this leads to bound-state confinement, meaning that the bound states remain trapped in region $A$ where they are produced (see Fig. 1). This was observed already in Ref. [14] (see also [15]). Moreover, the sign change of the bound-state velocities was analyzed at the level of the space-time trajectories in Ref. [16], and it has striking effects on the entanglement dynamics [16, 17]. Besides confinement, the presence of bound states gives rise to several intriguing transport behaviors, such as nonanaliticities in the profile of local observables [18].
[…]
Here by combining the time-dependent Density Matrix Renormalization Group [19] ($tDMRG$) with state-of-the-art integrability methods [20, 21], we provide a thorough investigation of bound-state confinement in the $XXZ$ chain at anisotropy $\Delta>1$. We show that for generic initial states of region $A$, multiparticle bound states remain confined in $A$. Precisely, this happens provided that interactions, i.e., the anisotropy $\Delta$, are strong enough, and there is a finite density of magnon-like quasiparticles, i.e., unbound ones, in region $A$. Confinement is present at large $\Delta$, and it can be understood perturbatively in the limit $\Delta\to\infty$. Upon lowering $\Delta$ there is a threshold below which the lighter bound states, i.e., formed by two elementary quasiparticles, are not confined. This means that two-particle bound states are allowed to leave region $A$ (see Fig. 1). This is in agreement with the results of Ref. [14] and [15] for the case of Néel and Majumdar-Ghosh initial states. Moreover, for typical initial states of region $A$ there is a hierarchical emission of bound states, meaning that heavier bound states are “liberated” at lower $\Delta$. While confinement can be understood perturbatively, the de-confinement happens near $\Delta=1$, i.e., away from the perturbative regime. This means that the deconfinement pattern of the different bound states depends on the details of the initial state. Finally, it is important to stress that bound-state confinement happens at ballistic level, i.e., in the hydrodynamic limit $x,t\to\infty$ with $\zeta:=x/t$, with $x$ the distance from the interface between $A$ and $B$ (see Fig. 1).
Crucially, since the size $\ell$ of $A$ (see Fig. 1) is finite, at long times tmuch-greater-$t\gg\ell$ the density of magnon quasiparticles in $A$ vanishes. In turn, this implies that after the “evaporation” of all the magnons from region $A$, the system cannot support bound-state confinement. Thus, even for large $\Delta$ bound states cannot remain confined at arbitrarily long times (see Fig. 1 (c)) if region $A$ is finite. Still, the short-time behavior of the spin projectors $P_{\uparrow^{\otimes n}}$ on $n$ consecutive sites are reminiscent of confinement. Following Ref. [22], we use $P_{\uparrow^{\otimes n}}$ as witness for $n$-particles bound-states. We observe that if interactions are not strong enough to support confinement in the hydrodynamic limit, at short times the projectors exhibit a “halo” signal around the interface between $A$ and its complement. On the other hand, if confinement is present, the projectors exhibit a “bottleneck” at the interface between $A$ and $B$ before the “deconfinement” of the bound states happens at longer times.
Finally, we confirm the confinement scenario by considering a “cut” quench (see Fig. 1 (b)) in which the expansion is stopped after a time $t^{*}$. The stopping time is chosen such that the density of magnons in $A$ is finite, i.e., before the bound-state deconfinement starts. For $t>t^{*}$ only part $B$ of the chain is further evolved. If interactions are strong enough to permit bound-state confinement, the projectors $P_{\uparrow^{\otimes n}}$ do not show any signal at $t>t^{*}$ in region $B$.
In this work we focus on the vacuum expansion of the $p$-Néel states, which are “fattened” version of the Néel state. They are defined by repetition of a unit cell containing $p$ up spins followed by $p$ down spins (see Fig. 1 (a)). In the limit $p\to\infty$ the $p$-Néel states converge to the domain-wall state. Interestingly, for $\Delta>1$ transport is suppressed after the quench from the domain-wall state [23]. At $\Delta=1$ transport is neither ballistic nor diffusive, but a superdiffusive dynamics is expected [24]. An interesting feature of the $p$-Néel state is that by changing $p$, one can change the bound-state content. Precisely, states with larger $p$ contain more bound states. For $p=1$ one recovers the standard Néel state, which is somewhat special because for $p=1$ the $GGE$ describing the bulk of $A$ can be determined analytically [25]. This is not the case for $p>2$ [26], although the $GGE$ can be obtained at modest numerical cost by using the results of Ref. [20]. If region $A$ (see Fig. 1) is prepared in the Néel state (with $p=1$), at large enough $\Delta\gtrsim 1.4$, bound states are confined, as observed in Ref. [15]. Interestingly, for $2\leq p\lesssim 4$, only bound states with size $n>p$ are confined, whereas for $p\gtrsim 4$ bound states of size $n=p$ are also confined. Since bound states with $p=n$ dominate the behavior of local observables after the quench from the $p$-Néel states, this is compatible with the absence of transport in the $p\to\infty$ limit.
Finally, we show that bound-state confinement has striking consequences on the dynamics of entanglement entropy. First, within the quasiparticle picture for entanglement dynamics [27, 28, 29], entanglement is created in region $A$. Transport between $A$ and the rest does not generate new entanglement, but entanglement is simply “transported” from $A$ to the rest [17]. We focus on the full expansion of the Néel state. We numerically show that the hierarchical emission of bound states of different sizes is reflected in a multistage growth of the entanglement entropy.
The manuscript is organized as follows. In section 2 we introduce the $XXZ$ spin chain and the out-of-equilibrium protocols. In section 3 we review the Thermodynamic Bethe Ansatz of the model. In section 4 we discuss the Generalized Hydrodynamics ($GHD$) approach. In section 5 by using the $GHD$ approach, we investigate bound-state confinement. In particular, in section 5.1 we provide numerical $GHD$ results confirming bound-state confinement after the expansion of the $p$-Néel states. In section 6, we discuss numerical $tDMRG$ results for the trap-expansion dynamics. In particular, in section 6.1 we discuss the validity of the $GHD$ approach. In section 6.2 and section 6.3 we provide $tDMRG$ data for the full-expansion dynamics and the cut protocol (see Fig. 1 (a) and (b)), respectively. In section 7, we focus on the entanglement entropy after the full-expansion protocol. We conclude and discuss future directions in section 8.
2 Model and quantum transport protocols
Here we focus on the spin-$1/2$ anisotropic Heisenberg chain ($XXZ$ chain) defined by the Hamiltonian
$H=\sum_{j=1}^{L}\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}+\Delta(\sigma_{j}^{z}\sigma_{j+1}^{z}-1),$ (1)
where $\sigma_{j}^{\alpha}$ are Pauli matrices, $\Delta$ is the co-called anisotropy parameter, and $L$ is the length of the chain. We focus on the situation with $\Delta\geq 1$. We employ periodic boundary conditions. However, since our results do not depend on boundary conditions we employ open boundary conditions in the $tDMRG$ numerical simulations [19]. For both periodic and open boundary condition, the $XXZ$ is integrable by the Bethe Ansatz [2].
Here we consider the expansion of “fat” Néel states (see Fig. 1 (a)). They are obtained by repetition of a unit cell containing $2p$ spins, with $p$ spins being up and $p$ down. We denote these $p$-Néel states as $|N_{p}\rangle$. The generic $p$-Néel state reads as
$|N_{p}\rangle:=|\uparrow^{\otimes p}\downarrow^{\otimes p}\rangle^{\otimes L/(2p)},$ (2)
where $\otimes$ denotes the tensor product.
[…]
In protocol $(a)$ the full system undergoes unitary dynamics under the Hamiltonian (1). In the second protocol (Fig. 1 $(b)$) we evolve the full system up to a time $t^{*}$. As the dynamics progresses, the initial $p$-Néel state “melts”. For $t>t^{*}$ the evolution happens under the Hamiltonian (1) restricted to the region $B$ on the right of $|N_{p}\rangle$ (see Fig. 1), whereas part $A$ of the system is “frozen”. We choose $t^{*}$ large enough to ensure that the system is in the hydrodynamic limit but not too large, so that region $A$ contains a finite density of unbound particles. As we will discuss in section 5, the latter condition is crucial to observe bound-state confinement.
As anticipated, in both protocols $(a)$ and $(b)$ of Fig. 1, at $t>0$ part $A$ of the chain “melts” because $|N_{p}\rangle$ is not an eigenstate of (1). First, the dynamics of an infinite homogeneous chain prepared in $|N_{p}\rangle$ gives rise to a Generalized Gibbs Ensemble [11] ($GGE$) with extensive thermodynamic entropy. Provided that $\ell$ is large enough, local properties in the bulk of region $A$ are described by the $GGE$ arising in the homogeneous chain. Oppositely, at the interface between region $A$ and the rest, the dynamics leads to the formation of a nontrivial profile. Specifically, in the long-time limit, provided that tmuch-less-$t\ll\ell$, to ensure that the central region has a finite density of excitations, ballistic transport sets in. This means that expectation values of generic local observables ${\mathcal{O}}(x,t)$ (with $x$ measured from the interface between $A$ and rest) are functions of $\zeta:=x/t$, and not of space and time separately. The expectation values of ${\mathcal{O}}(\zeta)$ can be obtained (see section 4) by using the Generalized Hydrodynamics [13, 12] ($GHD$). Crucially, since the $XXZ$ chain is interacting, for sufficiently large $\Delta>1$ transport of bound states (of arbitrary size) of elementary excitations from region $A$ to $B$ is inhibited in the ballistic regime. This means that the expectation values ${\mathcal{O}}(\zeta)$ for $\zeta>0$ are fully determined by magnon-like excitations. This bound-state confinement happens for both the protocols of Fig. 1. As already stressed, for the protocol in Fig. 1 $(a)$ the dynamics at asymptotically long times leads to zero density of magnon-like excitations in the bulk of $A$. As we will discuss in section 5 this will lead to the deconfinement of bound states.
3 Thermodynamic Bethe Ansatz ($TBA$) for quenches from homogeneous states in the $XXZ$ chain
To derive bound-state confinement after the trap expansion of the $p$-Néel states (cf. (2)), it is necessary to construct the $GGE$ describing the steady state after the quench from the homogeneous $p$-Néel states. Here we provide a self-contained discussion of the derivation of the $GGE$ for a generic pre-quench initial state. We follow Ref. [20, 21].
We employ the framework of the Thermodynamic Bethe Ansatz [2] ($TBA$). The eigenstates of the $XXZ$ chain are identified by a set of complex numbers $\lambda_{j}$, which are called rapidities. Each eigenstate is written as |$|\{\lambda_{j}\}\rangle$, and the rapidities $\lambda_{j}$ label the different quasiparticles in the system. In the thermodynamic limit $L\to\infty$ at fixed density of quasiparticles, it is convenient to work with the rapidity density. Moreover, in the thermodynamic limit generic integrable systems exhibit different families of quasiparticles. They correspond to bound states of elementary, magnon-like, excitations. The corresponding rapidities form string patterns in the complex plane [1]. Then, any thermodynamic macrostate of the system is characterized by a set of rapidity densities $\{\rho_{j}(\lambda)\}_{j=1}^{\mathcal{N}}$, where the number of families ${\mathcal{N}}$ depends on the model under consideration. For the $XXZ$ chain at $\Delta>1$ one has ${\mathcal{N}}\to\infty$, meaning that bound states of arbitrary sized are present. Similar to the densities $\rho_{j}(\lambda)$, one defines the hole densities $\rho_{h,j}(\lambda)$. The particle and hole densities for any legitimate $TBA$ macrostate satisfy the $TBA$ equation [2] as
$\rho_{t,n}(\mu)=a_{n}(\mu)-a_{n,j}\star\rho_{j},$ (3)
where $\rho_{t,n}=\rho_{n}+\rho_{h,n}$ is the total density. The $\star$ symbol in (3) denotes the convolution, i.e.,
[…]
In (3), we defined $a_{n}$ as [2]
$a_{n}(\mu)=\frac{1}{2\pi}\frac{\sinh(n\eta)}{\cosh(\eta)-\cos(2\mu)},$ (5)
where the parameter $\eta$ is related to the anisotropy $\Delta$ as
$\eta:=\mathrm{arccosh}(\Delta).$ (6)
Moreover, the parameters $a_{n,m}$ in (3) are obtained as
$a_{n,m}(\lambda):=\partial_{\lambda}\Theta_{n,m}(\lambda)=(1-\delta_{nm})a_{|n-m|}(\lambda)+2a_{|n-m|+2}(\lambda)+\cdots+a_{n+m-2}(\lambda)+a_{n+m}(\lambda),$ (7)
where $a_{n}(\lambda)$ is defined in (5). Here $\Theta_{n,m}$ is related to the scattering phase $S_{n,m}(\lambda)$ between a $n$-particle and a $m$-particle bound state as $\Theta_{n,m}(\lambda)=-i\ln(S_{n,m})$. For the following, it is useful to notice that the “bare” velocities of the quasiparticles are given as
$v^{b}_{n}=\frac{\partial\varepsilon_{n}(\mu)}{\partial\mu}\left(\frac{\partial p_{n}}{\partial\mu}\right)^{-1},$ (8)
where the energy $\varepsilon_{n}$ and the quasimomentum $p_{n}$ are defined as
$\varepsilon_{n}(\mu)=\frac{\sinh(\eta)\sinh(n\eta)}{\cosh(n\eta)-\cos(2\mu)},\quad p_{n}(\mu)=-i\sum_{j=1}^{n}\ln\left[\frac{\sin\left(\mu+i\frac{\eta}{2}(n+1-2j)+i\frac{\eta}{2}\right)}{\sin\left(\mu+i\frac{\eta}{2}(n+1-2j)-i\frac{\eta}{2}\right)}\right].$ (9)
Notice that Eq. (3) allows to obtain only one of the $TBA$ densities, either $\rho_{n}(\lambda)$ or $\rho_{h,n}(\lambda)$. Crucially, to uniquely identify a thermodynamic macrostate, one has to determine both the particle and the hole densities. In the following, instead of $\rho_{n}(\lambda)$ and $\rho_{h,n}(\lambda)$ we will often consider the ratios $\eta_{n}(\lambda)$ and $\vartheta_{n}(\lambda)$ defined as
$\eta_{n}:=\frac{\rho_{h,n}}{\rho_{n}},\quad\vartheta_{n}(\lambda):=\frac{\rho_{n}(\lambda)}{\rho_{t,n}(\lambda)}.$ (10)
We now show that the information about the expectation values of local and quasilocal conserved quantities over the pre-quench initial state allows to fix uniquely the $TBA$ densities identifying the $GGE$ that describes the steady state after the quench.
A central object in the Algebraic Bethe Ansatz treatment [30] of the $XXZ$ chain is the spin-$j/2$ Lax operator $L_{j}(\mu)$ defined as
$L_{j}(\mu)=\frac{1}{\sinh(\eta)}\left(\begin{array}[]{cc}\sin(\mu+i\eta s^{z})&i\sinh(\eta)s^{-}\\ i\sinh(\eta)s^{+}&\sin(\mu-i\eta s^{z})\end{array}\right).$ (11)
where $\eta$ is defined in (6) and $\mu$ is a real parameter. For $j=1$, which corresponds to spin-$1/2$ representation, one has $s^{\alpha}=\sigma^{\alpha}/2$, with $\alpha=+,-,z$. Moreover, $\sigma^{+},\sigma^{-}$ are the standard combinations $\sigma^{\pm}:=\sigma^{x}\pm i\sigma^{y}$, with $\sigma^{\alpha}$ Pauli matrices. The Lax operator acts as a $2\times 2$ matrix over the physical Hilbert space $\mathcal{V}_{1}$ of the model. The entries of (11) live in the “auxiliary” Hilbert space $\mathcal{V}_{j}$ of dimension $j+1$. Thus, the global Hilbert space is $\mathcal{V}_{1}\otimes\mathcal{V}_{j}$. The auxiliary space is spanned by vectors $|n\rangle$ with $n=0,1,,\dots,j$, with $|0\rangle$ being a highest-weight vector. The operators $s^{\alpha}$ with $\alpha=\pm,z$ are $q$-deformed spin-$j$ operators, forming the $j$-representation of the $q$-deformed algebra $SU(2)$. They obey the commutation relations
$[s^{+},s^{-}]=[2s^{z}]_{q},\quad q^{\pm s^{z}}=q^{\pm 1}s^{\pm}q^{\pm s^{z}},\quad\mathrm{with}\,\,[x]_{q}:=\frac{\sinh(\eta x)}{\sinh(\eta)}.$ (12)
The action of the operators $s^{\pm},s^{z}$ on the states $|n\rangle$ is given as
$s^{z}|n\rangle=\left(\frac{j}{2}-n\right)|n\rangle$ (13)
$s^{+}|n\rangle=\sqrt{[j-n]_{q}[n+1]_{q}}|n+1\rangle$ (14)
$s^{-}|n+1\rangle=\sqrt{[j-n]_{q}[n+1]_{q}}|n\rangle.$ (15)
Thus, the eigenvalues of $s^{z}$ are $-j/2,\dots,j/2$.
As usual in Algebraic Bethe Ansatz [30], one can construct a set of conserved quantities from the transfer matrix $T_{j}(\mu)$, which is defined as
$T_{j}(\mu)=\mathrm{Tr}_{\mathcal{V}_{j}}\left[L_{j}^{(1)}(\mu)L_{j}^{(2)}(\mu)\cdots L_{j}^{(L)}(\mu)\right],$ (16)
where the trace is over the auxiliary space. In (16) the superscript $k$ in $L_{j}^{(k)}(\mu)$ is to stress that it acts in the space $\mathcal{V}_{1}\otimes\mathcal{V}_{j}$, where $\mathcal{V}_{1}$ is the spin-$1/2$ Hilbert space of site $k$ of the chain. Moreover, in (16), $T_{0}(\mu)$ is given as
$T_{0}(\mu)=\left(\frac{\sin(\mu)}{\sinh(\eta)}\right)^{L}\mathds{1}.$ (17)
A complete set of local and quasilocal conserved charges $X_{j}(\mu)$, which commute with $H$, is defined as [21]
$X_{j}(\mu)=\frac{1}{L}\frac{1}{2\pi i}\partial_{\mu}\ln\left(\frac{T^{[+]}_{j}(\mu))}{T^{[j+1]}_{0}(\mu)}\right),$ (18)
where we introduced the notation
$F_{j}^{[\pm k]}(\mu):=F_{j}\left(\mu\pm ik\frac{\eta}{2}\right),$ (19)
for a generic operator $F_{j}(\mu)$, and $T_{j}(\mu)$ is defined in (16). The transfer matrices $T_{j}(\mu)$ satisfy the commutation relations
$[T_{j}(\mu),T_{j^{\prime}}(\mu^{\prime})]=0,\quad\forall j,j^{\prime}\in\mathbb{Z}_{\geq 0},\,\forall\mu,\mu^{\prime}\in\mathbb{C}.$ (20)
One can also check that the transfer matrices $T_{j}$ in (22) satisfy a Hirota equation (or the so-called $T$-system) as [21]
$T_{j}^{[+]}T_{j}^{[-]}=T_{0}^{[j+1]}T_{0}^{[-j-1]}+T_{j-1}T_{j+1}.$ (21)
Importantly, one can also verify the $T$-$Q$ relation as [31]
$T_{1}Q=T_{0}^{[+]}Q^{*[-2]}+T_{0}^{[-]}Q^{[2]},$ (22)
where $Q(\mu)$ is the so-called Baxter $Q$-operator. The spectrum of the $Q$ operator is given as $Q(\mu):=\prod_{k=1}^{M}\sin(\mu-\mu_{k})$, with $\mu_{k}$ solutions of the Bethe equations. Crucially, the $Q$ operator commutes with the transfer matrices $T_{j}(\mu)$, and it can be used to construct a solution of the Hirota equations (21) (see Ref. [20]). In particular, by expanding the solution of (21) in terms of the $Q$ operator, it has been shown that in the thermodynamic limit $L\to\infty$ one has that [21]
$\lim_{L\gg 1}\frac{T_{j}^{[+]}(\mu)}{T_{0}^{[j+1]}(\mu)}=\frac{Q^{[-j]}(\mu)}{Q^{[+j]}\mu}=\prod_{k=1}^{M_{j}}\frac{\sin(\mu-\mu_{k}-j\eta/2)}{\sin(\mu-\mu_{k}+j\eta/2)}=\prod_{k=1}^{M_{j}}S_{j}(\mu-\mu_{k}),$ (23)
where we defined the scattering matrix $S_{j}(\mu-\mu_{k})$ between a magnon and a $j$-particle bound state, and $M_{j}$ is the number of $j$-particle bound states in the Bethe state. Importantly, we can now employ Eq. (23) to compute the expectation value of the local and quasilocal conserved quantities $X_{j}(\mu)$ (cf. (18)). Specifically, from (18) and (23) we obtain
$X_{j}(\mu)=\frac{1}{L}\frac{1}{2\pi i}\sum_{k=1}^{M_{j}}\partial_{\mu}\ln(S_{j}(\mu-\mu_{k})),$ (24)
where $S_{j}(\mu-\mu_{k})$ is defined in (23). Notice also that $\partial_{\mu}\ln(S_{j}(\mu-\mu_{k}))$ is defined in (7). Following the standard $TBA$ approach [2], in the thermodynamic limit we can replace the sum over the rapidities $\mu_{k}$ with an integral weighted with the root density $\rho_{k}$ as
[…]
where the $\star$ symbol denotes convolution (cf. (4)). Crucially, it can be shown that [20] the kernel $G_{j,k}$ is the Green’s function for the discrete Laplacian operator $\Box$, which acts on a generic function $f_{j}$ as
$\Box f_{j}:=f_{j}^{[+]}+f_{j}^{[-]}-f_{j+1}-f_{j-1},\quad\left(\Box G\right)_{j,k}=\delta_{j,k}\delta(\lambda),$ (26)
where $\delta(\lambda)$ is a Dirac delta. Now, we can invert (25) by applying the Laplacian on both sides, to obtain
$\rho_{n}(\mu)=\Box X_{n}(\mu),$ (27)
where $\Box X_{n}(\mu)$ is defined in (26). Eq. (27) states that one can obtain the densities $\rho_{n}$ describing a thermodynamic macrostate (such as the $GGE$ arising after a quench) starting from the expectation values of the quasilocal charges $X_{n}$. By using the $TBA$ equation (3), one can obtain the analog of (27) for the hole densities $\rho_{h,n}$ as
$\rho_{h,n}=a_{n}-X_{n}^{[+]}-X_{n}^{[-]},$ (28)
where we employ the definition in (19), $X_{n}$ is the expectation value of the generic quasilocal conserved quantity on the initial state, and $a_{n}$ is the same as in (5).
To proceed, one has to determine the expectation value of both members of (18) over the initial state $\psi_{0}$. A major complication is that Eq. (18) requires the inverse of the transfer matrix $T_{j}^{[+]}(\mu)$, which is a daunting task in general. Fortunately, in the large $L$ limit one can show that [21]
$\lim_{L\gg 1}\frac{T^{[-]}_{j}(\mu)T_{j}^{[+]}(\mu)}{{T_{0}^{[-j-1]}}(\mu)T_{0}^{[j+1]}(\lambda)}=1.$ (29)
Eq. (29) provides an inversion formula. This is crucial to rewrite the logarithmic derivative in (18) as
$X_{j}(\mu)\simeq\frac{1}{L}\frac{1}{2\pi i}\frac{T_{j}^{[-]}(\mu)}{T_{0}^{[-j-1]}(\mu)}\partial_{\mu}\frac{T_{j}^{[+]}(\mu)}{T_{0}^{[j+1]}(\mu)},$ (30)
which holds in the large $L$ limit.
We are now ready to compute the expectation value of $X_{j}$ over the initial state $\psi_{0}$. Here we consider initial states $\psi_{0}$ that are obtained by repeating a unit cell of length $q=2p$, as it is the case for the $p$-Néel states (2). This allows us to rewrite (30) as
$X_{j}^{\psi_{0}}(\mu)=\frac{1}{L}\frac{1}{2\pi i}\mathrm{Tr}_{V_{j}}\left[\partial_{x}\left.\left(\langle\psi_{0}^{q}|\prod_{k=1}^{N_{q}}\frac{L_{j}^{(k),[-]}(\mu)}{L_{0}^{(k),[-j-1]}(\mu)}\frac{L_{j}^{(k),[+]}(\mu+x)}{L_{0}^{(k),[j+1]}(\mu+x)}|\psi_{0}^{q}\rangle\right)\right|_{x=0}\right]^{L/L_{q}},$ (31)
where we denoted with $X_{j}^{\psi_{0}}$ the expectation value of the charges $X_{j}$ over the initial state, and with $|\psi_{0}^{q}\rangle$ the unit cell that is used to build $|\psi_{0}\rangle=|\psi_{0}^{q}\rangle^{\otimes L/L_{q}}$. In (31) the Lax operators $L_{j}^{(k),[\pm]}$ are obtained from the original ones in (11) by using (19), and the trace is over the auxiliary Hilbert space. Notice that the expectation value over the state $|\psi^{q}_{0}\rangle$ of the unit cell involves the physical Hilbert space. Finally, in (31) we shifted the argument of the second term in the brackets to bring the derivative outside of the expectation value. Eq. (31), together with Eq. (27) and Eq. (28), is the main tool to determine the $GGE$. It is useful to observe that Eq. (31) can be rewritten as
$X_{j}^{\psi_{0}}(\mu)=\frac{1}{L}\frac{1}{2\pi i}\left[\mathrm{Tr}_{V_{j}}\left.\left(\partial_{x}\mathbb{T}_{j}^{\psi_{0}^{q}}(x,\mu)\right|_{x=0}\right)\right],$ (32)
where we defined $\mathbb{T}^{\psi_{0}^{q}}_{j}$ as
$\mathbb{T}_{j}^{\psi_{0}^{q}}(x,\mu):=\langle\psi_{0}^{q}|\mathbb{L}_{j}^{(N_{q})}(x,\mu)\mathbb{L}^{(N_{q}-1)}\cdots\mathbb{L}^{(1)}_{j}(x,\mu)|\psi_{0}^{q}\rangle,$ (33)
with
$\mathbb{L}_{j}^{(k)}(x,\mu)=\frac{L_{j}^{(k),[-]}(\mu)L_{j}^{(k),[+]}(\mu+x)}{L_{0}^{(k),[-j-1]}(\mu)L_{0}^{(k),[j+1]}(\mu+x)}.$ (34)
4 Generalized Hydrodynamics ($GHD$) for quenches from inhomogeneous initial states
Here we introduce the framework of Generalized Hydrodynamics [13, 12] ($GHD$) to describe the expansion dynamics (see Fig. 1). This will allow us to characterize the bound-state confinement in section 5.
In the standard $GHD$ setup two semi-infinite homogeneous but different initial states are joined together at $t=0$ at the origin $x=0$. At $t>0$ the full system evolves under a homogeneous integrable Hamiltonian. For typical initial states, at long times and large distances with the ratio $\zeta:=x/t$ fixed, the behavior of the system is described by a $\zeta$-dependent $GGE$. This $GGE$ is fully characterized by a set of $TBA$ densities $\rho_{n,\zeta}(\lambda)$ and $\rho_{h,n,\zeta}(\lambda)$. The $TBA$ densities satisfy the $TBA$ equations (3) for any $\zeta$ and the $GHD$ equation, which reads as [12, 13, 17]
$[\zeta-v_{n,\zeta}(\lambda)]\partial_{\zeta}\vartheta_{n,\zeta}(\lambda)=0,\quad\mathrm{with}\,\zeta:=x/t.$ (35)
Here $\vartheta_{n,\zeta}$ are the filling functions (cf. (36)) defined as
$\vartheta_{n,\zeta}(\lambda)=\frac{\rho_{n,\zeta}(\lambda)}{\rho_{t,n,\zeta}(\lambda)},$ (36)
with $\rho_{t,n,\zeta}=\rho_{n,\zeta}+\rho_{h,n,\zeta}$ the total density. In (35), $v_{n,\zeta}$ are the dressed velocities, which are defined as
$v_{n,\zeta}=\frac{[\partial_{\lambda}\varepsilon_{n,\zeta}]^{\mathrm{dr}}(\lambda)}{[\partial_{\lambda}p_{n,\zeta}]^{\mathrm{dr}}(\lambda)},$ (37)
where $\varepsilon_{n,\zeta}(\lambda)$ and $p_{n,\zeta}(\lambda)$ are respectively the bare energy and quasimomentum of the quasiparticles, defined in (9). Given a generic “bare” quantity $\tau^{\mathrm{b}}_{n,\zeta}(\lambda)$, its dressed version $\tau_{n,\zeta}^{\mathrm{dr}}$ is obtained by solving the integral equation as [32, 30]
[…]
where $\partial_{\lambda}\Theta_{n,m}(\lambda)$ is defined in (7), and the filling function $\vartheta_{n,\zeta}(\lambda)$ is given in (36). It is clear from the definition of the momentum $p_{n,\zeta}$ as [2]
$\partial_{\lambda}p_{n,\zeta}(\lambda)=2\pi\rho_{t,n,\zeta}(\lambda),$ (39)
that the dressing of $\partial_{\lambda}p_{n}$ is obtained by considering $\tau^{\mathrm{dr}}_{n,\zeta}=2\pi\rho_{t,n,\zeta}(\lambda)$ in (38).
To solve (35) one has to provide the boundary conditions at $\zeta=\pm\infty$. Let us consider the prototypical setup in which the initial state is obtained by joining two semi-infinite chains prepared in two macroscopically different states. This is the so-called bipartite quench protocol [17]. Let us denote by $\vartheta_{n,-\infty}$ and $\vartheta_{n,+\infty}$ the two thermodynamic macrostates, i.e., the $TBA$ filling functions, describing the steady state arising at infinite time after the quench from the left and right homogeneous states. The functions $\vartheta_{n,\pm\infty}$, provide the boundary conditions in (35). Before proceeding, one can observe that Eq. (35) can be solved for each $\zeta$ separately. It is straightforward to check that the solution of (35) can be written as
$\vartheta_{n,\zeta}(\lambda)=\Theta_{\mathrm{H}}(v_{n,\zeta}-\zeta)[\vartheta_{n,-\infty}(\lambda)-\vartheta_{n,+\infty}(\lambda)]+\vartheta_{n,+\infty}(\lambda).$ (40)
where $\Theta_{\mathrm{H}}(x)$ is the Heaviside step function. The dressed velocity $v_{n,\zeta}$ in (40) is obtained by using (38) and (37) (39). Notice that since $v_{n,\zeta}$ depends on $\vartheta_{n,\zeta}$ via (38), Eq. (40) is not an explicit solution of (35). To obtain $\vartheta_{n,\zeta}$ one can use an iterative strategy, starting from (40) fixing $v_{n,\zeta}=v^{b}_{n,\zeta}$, with $v^{b}_{n,\zeta}$ the “bare” velocities (cf. (8)). Then, one substitutes Eq. (40) in (38) to obtain the dressed $\partial_{\lambda}\varepsilon_{n,\zeta}$. By using the $TBA$ equation (3) one obtains $\rho_{t,n,\zeta}$ and the new dressed velocities via (37). Finally, these steps are iterated until convergence.
Notice that although the setup of the bipartite quench is different from the one described in Fig. 1 (a), the two protocols give the same dynamics provided that $\ell$ (see Fig. 1) is large enough to ensure that the bulk of region $A$ is in the hydrodynamic limit and $1\ll v_{\mathrm{max}}t\ll\ell$ is large, but short enough to neglect the fact that region $A$ is finite. Here $v_{\mathrm{max}}$ is the maximum velocity in the system, i.e, $v_{\mathrm{max}}=\max_{n,\zeta,\lambda}v_{n,\zeta}(\lambda)$.
In the following we focus on the macrostate with $\zeta=0$, which describes the interface between the two chains. We also consider the situation in which the right chain is prepared in the vacuum state (see Fig. 1). This corresponds to $\vartheta_{n,+\infty}=0$ for any $n$. Now, Eq. (40) becomes
$\vartheta_{n,0}(\lambda)=\Theta_{\mathrm{H}}(v_{n,0})\vartheta_{n,-\infty}(\lambda).$ (41)
In the following we neglect the subscript $\zeta$ in $\vartheta_{n,\zeta}$ and $v_{n,\zeta}$ since we always consider $\zeta=0$.
5 Bound-state confinement after trap-expansion experiments
Here we discuss the confinement of the bound states for the expansion protocol of Fig. 1 for generic initial state of part $A$ (see Fig. 1). The state of $A$ corresponds to fixing $\vartheta_{n,-\infty}$ in (41). Let us summarize our main results. We show that for moderately large $\Delta$ ballistic transport of bound states is inhibited, for quite generic initial states of part $A$. Specifically, in the hydrodynamic limit only quasiparticles with $n=1$, i.e., unbound quasiparticles, can be present at $\zeta\geq 0$. This means that all the bound states with $n>1$ remain confined within region $A$ in Fig. 1 (a).
Bound-state confinement can be understood perturbatively in the large $\Delta$ limit, and it relies on the behavior of the scattering phases (7). Indeed, the only assumption on the initial state of region $A$ is that it gives rise to $TBA$ macrostates with ${\mathcal{O}}(1)$ density of quasiparticles with $n=1$. Although bound-state confinement is established in the large $\Delta$ limit, it survives up to $\Delta\approx 1$. Precisely, for the Néel state bound state confinement survives up to [15] $\Delta=\Delta^{*}\approx 1.4$. Upon lowering $\Delta$ below $\Delta^{*}$, bound states with $n=2$ start to leave region $A$. For $\Delta\ll\Delta^{*}$ bound states with $n>2$ are not confined. However, while the threshold $\Delta^{*}$ for confinement of the two-particle bound states is understood perturbatively, this is not the case for the larger bound states. Indeed, the deconfinement of bound states with $n>2$ happens in the nonperturbative regime at $\Delta\approx 1$, and is sensitive to the details of the initial state. Crucially, since region $A$ is finite (see Fig. 1), the condition that the density of quasiparticles with $n=1$ is finite in $A$, which ensures the bound-state confinement, is necessarily violated at long times. This implies that at times tproportional-$t\propto\ell$ excitations with $n>1$ can leave region $A$, i.e., they are not confined anymore.
Bound-state confinement manifests itself in the fact that the dressed velocities $v_{n}(\lambda)$ at $\zeta=0$ are negative for any $\lambda$ and for any $n>1$. To proceed, let us assume that only bound states with $n>n_{a}$, for some $n_{a}>1$, are confined. This implies that Eq. (38) for $\tau_{n}=\partial_{\lambda}\varepsilon_{n}$ becomes
$\partial_{\lambda}\varepsilon^{\mathrm{dr}}_{n}(\lambda)=\partial_{\lambda}\varepsilon^{\mathrm{b}}_{n}(\lambda)-\sum_{m=1}^{n_{\mathrm{a}}}\int\!\!\frac{d\mu}{2\pi}\partial_{\lambda}\Theta_{n,m}(\lambda-\mu)\vartheta_{m,-\infty}(\mu)\Theta_{\mathrm{H}}(\partial_{\mu}\varepsilon^{\mathrm{dr}}(\mu))\partial_{\lambda}\varepsilon_{m}^{\mathrm{dr}}(\mu),$ (42)
where $n\leq n_{\mathrm{a}}$, and $n_{\mathrm{a}}$ is such that $\partial_{\lambda}\varepsilon^{\mathrm{dr}}_{n}(\lambda)<0$ for any $\lambda$ and for any $n>n_{\mathrm{a}}$. Notice that in (42) we replaced $v_{n,0}$ with $\partial_{\lambda}\varepsilon_{n}^{\mathrm{dr}}$ in the argument of the Heaviside theta function. This is possible because $\rho_{t,n}(\lambda)$ (cf. (37)) is positive. Also, notice that in the integrand in (42) one has that $\partial_{\lambda}\Theta_{n,m}$ and $\vartheta_{n,-\infty}$ are both positive functions. For (42) to be consistent with confinement of the bound states with $n>n_{a}$, we require that the solutions of (42) satisfy the condition
$\partial_{\lambda}\varepsilon_{n}^{\mathrm{dr}}(\lambda)<0,\forall n>n_{\mathrm{a}},\forall\lambda,$ (43)
Clearly, while the dressing (38) for the $XXZ$ with $\Delta>1$ is a system of infinite equations since $n\in[1,\infty)$, the system (42) is finite. Still, given the solutions of (42), one has to check the validity of (43) for any $n>n_{a}$.
Before proceeding, it is interesting to observe that $n_{\mathrm{a}}>0$, meaning that not all the species of quasiparticles can be confined at the same time, at least in the hydrodynamic limit. Indeed, $n_{\mathrm{a}}=0$ implies (cf. (42)) that $\partial_{\lambda}\varepsilon^{\mathrm{dr}}(\lambda)=\partial_{\lambda}\varepsilon^{\mathrm{b}}(\lambda)$, and $\partial_{\lambda}\varepsilon^{\mathrm{b}}(\lambda)$ is not negative for all values of $\lambda$ (cf. (8)). This implies that the suppression of ballistic transport observed for instance in the domain-wall quench [23] cannot be attributed to the bound-state confinement.
The conditions under which the systems (42) and (43) admit a solution can be derived in the large $\Delta$ limit. First, let us observe that $0\leq\vartheta_{n,-\infty}(\lambda)\leq 1$ for any $\lambda$. Moreover, in the large $\Delta$ limit, at the leading order in powers of $\Delta$, one has that $\partial_{\lambda}\Theta_{n,m}(\lambda-\mu)={\mathcal{O}}(1)$ for any $n,m$. Precisely, for large $\Delta$ one has that
$\partial_{\lambda}\Theta_{n,m}(\lambda)\to\frac{1}{\pi}(n+m-1),\quad\mathrm{for}\,\,n,m\neq 1,$ (44)
and $\partial_{\lambda}\Theta_{n,m}(\lambda)\to(n+m)/\pi$ otherwise. Finally, it is straightforward to check that the bare velocities $v_{n}^{\mathrm{b}}$ (cf. (8)) at the leading order in the large $\Delta$ limit are
$v_{n}^{\mathrm{b}}(\lambda)=2z^{n-1}\sin(2\lambda)+{\mathcal{O}}(z^{n+1}),\quad\mathrm{with}\,\,z:=e^{-\eta}.$ (45)
Here we defined $z=\Delta-\sqrt{\Delta^{2}-1}$, and for large $\Delta$ one has $z\approx 1/(2\Delta)$. Moreover, unlike the dressed velocities, the bare ones do not depend on the pre-quench initial state, so (45) holds irrespective of the initial state of $A$ (see Fig. 1). Crucially, while the bare velocities of the quasiparticles with $n=1$ are ${\mathcal{O}}(1)$, the bound-state velocities for $\Delta>1$ decay exponentially with the bound-state size $n$. This hierarchy in the bare velocities is responsible for the bound-state confinement.
Let us consider the case with $\vartheta_{1,-\infty}(\lambda)={\mathcal{O}}(1)$, i.e., the situation in which in region $A$ there is a finite density of unbound quasiparticles. This happens for typical states, such as the Néel state and the Majumdar-Ghosh state [15]. We anticipate that the $p$-Néel states with $p>1$ do not satisfy this condition. Let us also assume that $n_{\mathrm{a}}=1$, i.e., bound states with $n\geq 2$ remain confined within $A$ during the expansion dynamics.
Now, let us consider Eq. (42) with $n=2$. It is clear that the second term in the right-hand side dominates for large $\Delta$ because $\partial_{\lambda}\varepsilon_{n}^{\mathrm{b}}={\mathcal{O}}(z)$ for $n=2$, whereas the first term in (42) is ${\mathcal{O}}(z)$. Clearly, the same argument applies also for $n>2$, implying that all the bound states with $n\geq 2$ are confined for sufficiently large $\Delta$. Interestingly, as the only dependence on $\lambda$ in the second term in (42) is in $\Theta_{n,m}$, which becomes independent of $\lambda$ for large $\Delta$, the dressed velocities of the confined bound states become “flat” as functions of $\lambda$.
Upon lowering $\Delta$, the two terms in (42) become of the same order, and quasiparticles with $n=2$ start to be transported in $B$ (see Fig. 1). Notice that while bound-state confinement can be established perturbatively, and even the threshold $\Delta^{*}$ at which bound states with $n=2$ are not confined can be calculated from the large $\Delta$ expansion, it is difficult to characterize analytically the scenario for $\Delta\to 1$. Typically, we observe that there is a “cascade” effect. Precisely, upon lowering $\Delta$, transport of bound states with progressively larger $n$ become allowed (see section 5.1).
The scenario outlined so far relies on the fact that at any time during the expansion $\vartheta_{1,-\infty}={\mathcal{O}}(1)$. This condition remains true at arbitrarily long times if region $A$ (cf. Fig. 1) is infinite. Oppositely, if region $A$ is finite, at long times one has that $\vartheta_{1,-\infty}$ vanishes. This implies that the second term in (42) becomes smaller. Thus, at long times the bound states with $n=2$ can leave region $A$ after the “evaporation” of the unbound particles. Moreover, the condition $\vartheta_{1,-\infty}$ is not valid for generic initial states. For instance, it is violated by all the $p$-Néel states with generic $p>1$. By using the framework of section 3, one can check that $\vartheta_{n,-\infty}={\mathcal{O}}(z^{2})$ for $n<p$, whereas $\vartheta_{n,-\infty}={\mathcal{O}}(1)$ for $n\geq p$. This has striking consequences on transport properties and bound-state confinement. Indeed, we anticipate that the fact that $\vartheta_{n,-\infty}={\mathcal{O}}(z^{2})$ implies a strong suppression of bound-state transport for $p\geq 4$. Let us first consider the case with $p=2$. Now, Eq. (42) suggests that bound states with $n=2$ are not confined. Indeed, the term with $m=1$ in the right-hand-side of (42) is ${\mathcal{O}}(z^{2})$ and it is negligible compared to $\partial_{\lambda}\varepsilon_{2}^{\mathrm{b}}={\mathcal{O}}(z)$. This means that unlike the case with $p=1$ the two-particle bound states cannot be confined by the magnons. Now, if we consider the case with $p=4$ we observe that since $\partial_{\lambda}\varepsilon_{4}^{\mathrm{b}}={\mathcal{O}}(z^{3})$, it is clear that the second term in (42) dominates because of the scatterings with the quasiparticles with $n\leq 4$, which contribute with terms with smaller powers of $z$. This implies that upon increasing $p$ the bound states that dominate the thermodynamic properties, i.e., the bound states with $n=p$, become confined, and transport is suppressed.
5.1 Bound-state confinement: Numerical GHD results
Figure 2: Group velocities $v_{n}$ of the $n$-particle bound states calculated over the $GGE$ with $\zeta:=x/t=0$, which in the hydrodynamic limit describes the interface between region $A$ and the rest (Fig. 1). At $t=0$ region $A$ is prepared in the $p$-Néel states (cf. (2)). The different panels correspond to different values of $p$ and $\Delta$. Panel (a) and (b) are for $p=1$. Panels (c-f) show results for $p=2-5$. In all the panels one has that $v_{1}(\lambda)$ is positive in an extended region in $\lambda$, signaling that magnon-like excitations are never confined. (a) The case of $p=1$, i.e., the Néel state. At $\Delta\lesssim 1.4$ $v_{2}<0$ for any $\lambda$, i.e., the two-particle bound states are confined. (b) Same as in (a) for $\Delta=1.01$. While bound states with $n=2$ are not confined, $v_{3}\leq 0$ for any $\lambda$, which means that three-particle bound states are confined. (c) The $p$-Néel state with $p=2$. Bound states with $n\geq 3$ are confined. At $\Delta\approx 1.09$ three-particle bound states are not confined anymore. (d) The case with $p=3$. At $\Delta\approx 1.02$ the quasiparticles with $n=4$ are not confined. (e) The case with $p=4$. Now, even quasiparticles with $n=4$ remain confined, at least up to $\Delta\approx 1.07$. (f) The Néel state with $p=5$. Similar to (e), quasiparticles with $n=p$ are confined.
Although qualitative features of bound-state confinement are understood perturbatively, to characterize quantitatively bound-state confinement it is necessary to determine numerically the dressed velocities $v_{n}(\lambda)$. Here we discuss numerical results confirming the scenario outlined in section 5. Precisely, we numerically solve the $GHD$ equation (35) to obtain the dressed velocities $v_{n}$ for $\zeta=0$. The numerical solution of (35) is standard. By using a Gauss discretization, Eq. (35) becomes an infinite system of nonlinear equations, which can be truncated considering only bound states of size $n\leq n_{\mathrm{max}}$. Finally, the truncated system can be solved with standard methods. Here we consider $n_{\mathrm{max}}=40$.
Our results are reported in Fig. 2. We plot $v_{n}(\lambda)$ for $1\leq n\leq 5$ versus $\lambda$. We consider as initial states the $p$-Néel states (cf. (2)) with $1\leq p\leq 5$. For $p=1$ the quench is the popular Néel quench. At large $\Delta$ one has that $v_{n}<0$ for any $\lambda$ and for any $n>1$, meaning that all the bound states are confined within region $A$. Fig. 2 (a) shows that for $\Delta\approx 1.4$ the faster bound states with $n=2$ start to be transmitted [15]. Upon lowering $\Delta$, at $\Delta\approx 1.01$, the bound states with $n=3$ are not confined, as shown in Fig. 2 (b). For $p=2$, as anticipated in section 5, the bound states with $n=2$ are not confined. We checked the absence of confinement up to $\Delta=5$, although we do not report the results in Fig. 2. On the other hand, Fig. 2 (c) shows that for $p=2$ at $\Delta\approx 1.1$ the three-particle bound states are still confined in $A$. A similar scenario occurs for $p=3$, i.e., the bound states with $n\leq 4$ are not confined, whereas the bound states with $n>4$ are all confined up to $\Delta\approx 1.02$. However, the scenario changes at $p=4$. As it is clear from Fig. 2 (e), for $p\geq 4$ the group velocities $v_{p}(\lambda)$ are negative for any $\lambda$. This holds true for $p=5$ as well (see Fig. 2 (f)). Since the thermodynamic properties of the system are dominated by the bound states with $n=p$, this means that transport becomes more and more suppressed upon increasing $p$.
6 Numerical tDMRG results
Here we discuss $tDMRG$ data [19] for the vacuum expansion protocols introduced in Fig. 1. Our results support the bound-state confinement scenario outlined in section 5. As a preliminary check, in section 6.1 we discuss $tDMRG$ data for full-expansion protocol $(a)$ in Fig. 1, considering times tmuch-less-$t\ll\ell$, such that the fact that region $A$ is of finite size can be neglected. We benchmark the results of section 4 for several $p$-Néel initial states. In section 6.2 we focus on the confinement of bound states for the full-expansion protocol (see Fig. 1 (a)). In section 6.3 we discuss the cut quench protocol (see Fig. 1 (b)).
Figure 3: Releasing the $p$-Néel states in the vacuum (see Fig. 1). The figure shows $tDMRG$ data for the local magnetization $\langle\sigma_{x}^{z}\rangle$ plotted as a function of $\zeta=x/t$. Here $x$ is distance from the interface between region $A$ and the rest (Fig. 1). The dashed line is the $GHD$ result valid in the limit $x,t\to\infty$ at fixed $\zeta$. We show results for $p\leq 3$ and $\Delta=1,2$. The different symbols denote different times $t$. Notice that for $p=3$ and $\Delta=2$ $t,x$ are not large enough to reach the hydrodynamic limit.
Figure 4: Releasing the Néel states in the vacuum (see Fig. 1). We plot the expectation value $\langle P_{x,\uparrow\uparrow}\rangle$ (cf. (47)) versus the rescaled time $\zeta=x/t$, with $x$ the distance from the interface between region $A$ and the rest (see Fig. 1). The size of region $A$ is $\ell=40$ sites. Panels (a) and (b) show the expansion dynamics for $\Delta=1$ and $\Delta=2$, respectively. The different symbols denote different times. We only show results up to $t\lesssim 15$, to ensure that in the bulk of region $A$ one has $\vartheta_{1,-\infty}={\mathcal{O}}(1)$, which implies bound-state confinement (see section 5). Notice that at $\zeta=0$ for $\Delta=1$ one has a finite value for $\langle P_{x,\uparrow\uparrow}\rangle$, whereas at $\Delta=2$ one has $\langle P_{x,\uparrow\uparrow}\rangle\approx 0$, reflecting bound-state confinement.
6.1 Validity of GHD
Before investigating bound-state confinement, it is crucial to assess the validity of the $GHD$ description for the protocols discussed in Fig. 1. This is summarized in Fig. 3. We consider the expansion of the $p$-Néel states with $1\leq p\leq 3$. The different panels show $tDMRG$ data for the local magnetization $\sigma_{x}^{z}$ as a function of $\zeta=x/t$. For each $p$ we show results for $\Delta=1,2$. Here $x$ is the distance from the right edge of region $A$ (see Fig. 1). In our $tDMRG$ simulations we consider times tmuch-less-$t\ll\ell$ to avoid effects due to the finite size of $A$. In Fig. 3 we provide $tDMRG$ data for several times up to $t\approx 20$ (different symbols in the Figure). The $tDMRG$ data exhibit collapse when plotted as function of $x/t$, confirming that $t\approx 20$ is large enough to ensure the hydrodynamic limit for all the values of $p$ and $\Delta$ considered, except for $p=3$ with $\Delta=2$ (see Fig. 3 (f)). For $p=3$ and $\Delta=2$ the evolution in the bulk of region $A$ is slower compared to the other cases. Indeed, at $\zeta\leq-1$, the magnetization profile exhibits strong oscillations, which are reminiscent of the structure of the initial state. In all the panels, the dashed curve are the $GHD$ results obtained by solving numerically the $GHD$ equation (35). The numerical results were obtained by truncating the infinite system of equations (35) keeping only the bound states with $n\leq n_{\mathrm{max}}=40$. The $GHD$ equation (35) allows to obtain the occupation functions $\vartheta_{\zeta=0,n}(\lambda)$. The numerical results for $\vartheta_{0,n}(\lambda)$ can be substituted in the $TBA$ equations (3), after using that $\rho_{t,n}(\lambda)=\vartheta_{n}(\lambda)\rho_{n}(\lambda)$. Eq. (3) allows one to obtain the quasiparticle densities $\rho_{n}(\lambda)$, which encode information about local observables. For instance, the local magnetization is given as
[…]
For all the quenches that we explore in panels (a-e) of Fig. 3 the $GHD$ results are in satisfactory agreement with the $tDMRG$ data, despite the limitation to times $t\lesssim 20$. For the expansion of the $p$-Néel state with $p=3$ and $\Delta=2$ (panel (f) in Fig. 3) deviations from the $GHD$ prediction are large. Also, notice that for $\zeta>0$ one has that $\langle\sigma_{x}^{z}\rangle\approx-1$, signaling that transport in general is strongly suppressed for $p=3$.
As anticipated in section 5 for the $p$-Néel state with $p=1$, the bound states with $n=2$ remain confined in region $A$, as long as the system is in the hydrodynamic limit, and the $GHD$ description is valid. Specifically, this holds for $x,t\to\infty$ with $x/t$ fixed, with $x$ being the distance from the right edge of region $A$ (see Fig. 1). Moreover, for the bound-state confinement to happen, the condition $\vartheta_{1,-\infty}={\mathcal{O}}(1)$ has to be satisfied. This is the case for $p=1$, at least if tmuch-less-$t\ll\ell$.
Confinement of the the two-particle bound states is discussed in Fig. 4. The Figure shows $P_{x,\uparrow\uparrow}$ (cf. (47)) plotted versus $\zeta=x/t$. The projectors $P_{x,\uparrow^{\otimes m}}$ are defined as
$P_{x,\uparrow^{\otimes m}}:=\frac{1}{2^{m}}\prod_{j=0}^{m-1}(\mathds{1}+\sigma_{x+j}^{z}).$ (47)
Although the projectors (47) are not bona fide witnesses of bound states, we anticipate that $P_{x,\uparrow^{\otimes m}}$ is sensitive to the presence of a $m$-particle bound state at position $x$. Panels (a) and (b) show $tDRMG$ data for $P_{x,\uparrow\uparrow}$ for the $XXZ$ Hamiltonian with $\Delta=1$ and $\Delta=2$, respectively. In each panel the different symbols correspond to different times. For both $\Delta=1$ and $\Delta=2$ the $tDMRG$ data exhibit scaling as a function of $\zeta$. Crucially, while for $\Delta=1$ (Fig. 4 (a)) $\langle P_{x,\uparrow\uparrow}\rangle$ is nonzero for $\zeta>0$, one has $\langle P_{x,\uparrow\uparrow}\rangle\approx 0$ for any $\zeta>0$ for $\Delta=2$ (see Fig. 4). This reflects that while for $\Delta=1$ the bound states with $n=2$ are not confined, for $\Delta=2$ they remain confined in region $A$ during the expansion. Still, at $\zeta=0$, $\langle P_{x,\uparrow\uparrow}\rangle\approx 0.025$ for $\Delta=1$, which suggests that the density of two-particle bound states that are not confined is “small”, meaning that it could be challenging to observe it, for instance experimentally.
Figure 5: Releasing the Néel state in the vacuum. Regionn $A$ in Fig. 1 of size $\ell=20$ is prepared in the Néel state at $t=0$. The remaining part of the chain is in the vacuum state, i.e., the ferromagnetic state. At $t>0$ the chain is let to evolve under the $XXZ$ chain Hamiltonian with $\Delta=2$. Panel (a) shows $tDMRG$ data for the local magnetization $\langle\sigma^{z}_{x}\rangle$ at generic position $x$ and at different times. Notice the reflection at the boundary of the chain at long times. Panels (b) and (c) show the projectors $P_{x,\uparrow\uparrow}$ and $P_{x,\uparrow\uparrow\uparrow}$.
Figure 6: Releasing the Néel state in the vacuum. Region $A$ (with $\ell=20$ sites) of a chain is prepared in the Néel state at $t=0$. The remaining part of the chain is in the vacuum state, i.e., the ferromagnetic state. At $t>0$ the chain is let to evolve under the $XXZ$ chain Hamiltonian with $\Delta=2$. Panel (a) shows $tDMRG$ data for the local magnetization $\langle\sigma^{z}_{x}\rangle$ at generic position $x$ and at different times. Notice the reflection at the boundary of the chain at long times. Panels (b) and (c) show the projectors $P_{x,\uparrow\uparrow}$ and $P_{x,\uparrow\uparrow\uparrow}$, respectively (cf. (47)). The black lines in the panels are the maximum bare velocities $\max_{\lambda}v_{n}^{b}(\lambda)$ (cf. (8)) for $n=1$ (in panel (a)), $n=2$ (panel (b)), and $n=3$ (panel (c)).
6.2 Full expansion dynamics
In section 6.1 we established the validity of the $GHD$ description for the quenches from the $p$-Néel states. Moreover, Fig. 3 suggests that $t\approx 20$ is enough to access the hydrodynamic limit, which is essential to observe bound-state confinement. We now focus on the full expansion dynamics, i.e., considering times tmuch-greater-$t\gg\ell$. We show $tDMRG$ data for the expansion of the $p$-Néel states with $p=1$ for $\Delta=1$ and $\Delta=2$ in Fig. 5 and Fig. 6, respectively. The protocol is reported in Fig. 1. Region $A$ at center of the chain is prepared in the Néel state. Here we consider $\ell=20$. The reason for choosing $\ell=20$ is that it allows to simulate the dynamics using $tDMRG$ up to times $t\approx 100$. Indeed, entanglement, which is the major limitation of $tDMRG$, is produced in the bulk of region $A$ only, whereas the expansion does not contribute significantly to the entanglement growth. This also reflects that the $GHD$ dynamics preserves the thermodynamic entropy [14, 16]. As a consequence, the largest entanglement produced during the dynamics is bounded by $\ell$.
The different panels in Fig. 5 and Fig. 6 show different observables. Precisely, we consider the local magnetization $\sigma_{x}^{z}$, and the local projectors $P_{x,\uparrow\uparrow},P_{x,\uparrow\uparrow\uparrow}$ (cf. (47)). As anticipated, and has already observed in Ref. [22] and Ref. [10], $P_{x,\uparrow^{\otimes m}}$ is sensitive to bound states of $m$ spins starting from position $x$.
Let us first discuss the dynamics with the $XXZ$ chain at $\Delta=1$ (see Fig. 5). In Fig. 5 we plot the density of local magnetization $\langle\sigma_{x}^{z}\rangle$ as a function of position $x$ and time $t$. After the expansion starts, the central region “melts” as quasiparticles are emitted at its edges. First, several “jets” of excitations are visibile, traveling at different velocities. The fastest quasiparticles correspond to the magnon-like excitations, i.e., with $n=1$, whereas the jets moving at smaller velocities correspond to bound states with $n=2$ and $n=3$ quasiparticles. The dashed lines are the bare maximum group velocities of the quasiparticles (cf. (8)). Since the initial state is expanded in the vacuum, the density of quasiparticles at the edges of the jets is small and interactions are “weak”, which explains why the bare velocities describe the dynamics of the jets. In Fig. 5 (b) we show the dynamics of the projector $P_{x,\uparrow\uparrow}$. Now, the outermost jet in Fig. 5 (a) is not present, confirming that it is due to the quasiparticles with $n=1$. Oppositely, the two innermost jets of Fig. 5, corresponding to $n=2,3$ are still present. Similar behavior was observed in the expansion of few-spin initial states in Ref. [22]. Importantly, one should observe that at short times $t\lesssim 20$, $P_{x,\uparrow\uparrow}$ exhibits a “halo” feature at the edges of region $A$. This can be attributed to the fact that for $\Delta=1$ the two-particle bound states are not confined in the hydrodynamic limit. Notice, however, that the halo is quite weak, reflecting that the density of two-particle bound states at the interface between $A$ and the rest is small. A similar scenario is observed for $P_{x,\uparrow\uparrow\uparrow}$ (see Fig. 5 (c)). Now, the three-particle bound states behave as if they were confined up to $t\approx 20$. Notice that the fact (see Fig. 2 (b)) that $v_{3}\leq 0$ at $\Delta\approx 1.01$ suggests that the density $\rho_{3}(\lambda)$ of three-particle bound states at the interface between $A$ and the vacuum could be small. This implies that even though bound states with $n=3$ are not confined at $\Delta=1$, this could be difficult to observe.
Let us now discuss the same protocol in the $XXZ$ chain with $\Delta=2$ (see Fig. 6). The qualitative behavior is the same as in Fig. 5. However, the projector $P_{x,\uparrow\uparrow}$ at short times $t\lesssim 20$ exhibits a “bottleneck” feature. In contrast with $\Delta=1$, at $\Delta=2$ the two-particle bound states are confined (see Fig. 2 (c)). Moreover, in the hydrodynamic regime the density $\rho_{2}(\lambda)$ is zero for $\zeta=0$, and it is presumably small also at $\zeta\lesssim 0$. This explains the bottleneck in Fig. 6 At long times after all the magnons have left region $A$, the two-particle bound states are allowed to leave region $A$, as discussed in section 5, and as it is confirmed in Fig. 6 (b). A similar dynamics happens for the three-particle bound states, as shown in Fig. 6 (c).
Figure 7: Releasing the Néel state in the vacuum. Same setup as in Fig. 5 and Fig. 6. We show the magnetization profile $\langle\sigma_{x}^{z}\rangle$ as a function of the position $x$ in the chain. The different curves correspond to different times. Notice the different peaks, which correspond to different families of excitations. Precisely, the outermost peaks correpsond to magnon-like excitations, whereas the innermost ones to the bound states with $n\geq 2$.
Let us now discuss the space-time profiles of local observables. In Fig. 7 we show the profile of the local magnetization $\langle\sigma_{x}^{z}\rangle$ after the expansion of the Néel state for the $XXZ$ chain with $\Delta=1$ and $\Delta=2$, respectively. Both panels in Fig. 7 confirm that for $\Delta=1$ and $\Delta=2$ the bound states with $n=2$ and $n=3$ are not confined at long times, as already observed in Fig. 5 and Fig. 6. The $tDMRG$ data for $\Delta=1$ (left panel in Fig. 7) show a rapid melting of the profile. As the dynamics proceeds, $\langle\sigma_{x}^{z}\rangle$ exhibits a multi-peak structure, each peak corresponding to a different bound-state family. For $\Delta=1$ the peaks are blended together, reflecting that the bound states of different types start to leave region $A$ at the same time. In particular, already at $t=40$ the bound states with $n=2$ are moving towards the edges of region $A$. In the inset we show the dynamics of $P_{x,\uparrow\uparrow}$, which is sensitive to quasiparticles with $n\geq 2$ only. The inset shows a rapid melting of $P_{x,\uparrow\uparrow}$. In contrast, for $\Delta=2$ (right panel in Fig. 7) the dynamics is slower, and the peaks corresponding to different $n$ are well separated. Notice that at $t=80$ the lump around $x=100$ corresponds to the three-particle bound states, which remain almost locked at the center of the chain. This observation suggests that for $\Delta=2$ the trap-expansion protocol can be used to “distillate” states with large bound-state content. The inset shows the dynamics of $P_{x,\uparrow\uparrow}$. Notice that for $t\lesssim 10$ the bound states with $n=2$ are still within region $A$. Importantly, for $\ell=20$ the density of quasiparticles with $n=1$ in the bulk of $A$ starts to be small already at $t\approx 10$, implying that bound-state confinement is not supported anymore (see section 5).
Figure 8: Releasing the Néel state in the vacuum. At $t=0$ the central part $A$ (see Fig. 1) of a spin chain containing $\ell=20$ sites is prepared in the Néel state. At $t>0$ the chain is let to evolve under the $XXZ$ chain Hamiltonian with $\Delta=2$. For $0<t\leq 20$ the full chain evolves. At $t>20$ only the $B$ part (see Fig. 1) is evolved. Panels (a)-(d) show $tDMRG$ data for the local magnetization $\langle\sigma_{x}^{z}\rangle$ as a function of the position $x$ in the chain. Panels (e)-(h) show results for the projector $P_{x,\uparrow\uparrow}$. Notice that at $\Delta\approx 1.4$, $P_{x,\uparrow\uparrow}$ vanishes, signalling bound-state confinement.
6.3 Dynamics after the cut protocol
To investigate bound-state confinement, it is also convenient to employ the protocol in Fig. 1 (b). The full chain is now evolved with the $XXZ$ Hamiltonian up to a time $t^{*}$. Here $t^{*}$ is chosen large enough to ensure the validity of the hydrodynamic limit. As it is clear from section 6.1, $t\approx 20$ is sufficient to ensure the hydrodynamic limit. Moreover, $t^{*}$ is chosen short enough compared to $\ell$ to have $\vartheta_{1,-\infty}={\mathcal{O}}(1)$ in the bulk of $A$, which is essential to have bound-state confinement. Here we choose $t^{*}=20$ and $\ell=40$. We consider the $p$-Néel states with $p=1-3$.
6.3.1 $p$-Néel state with $p=1$
Our results for the Néel state (i.e., for $p=1$) are shown in Fig. 8. Panels $(a-d)$ shows the local magnetization $\sigma_{x}^{z}$, as a function of position $x$ and time $t$. As it is clear from the Figure, for $t>t^{*}$ (see Fig. 1), under the evolution with the Hamiltonian (1) restricted to $x>0$ the quasiparticles emitted from $A$ propagte ballistically in the vacuum. The dependence on $\Delta$ is weak. In the bottom row in Fig. 8 (panels (e-h) in the Figure) we report the dynamics of $P_{x,\uparrow\uparrow}$. For $\Delta=1$, $\langle P_{x,\uparrow\uparrow}\rangle$ exhibits a ballistic spreading. The nonzero signal reflects that the bound states with $n=2$ are not confined. $P_{x,\uparrow\uparrow}$ becomes smaller upon increasing $\Delta$. For instance, it is barely visible at $\Delta=2$. Notice that in the hydrodynamic limit one should expect that $\langle P_{x,\uparrow\uparrow}\rangle$ is exactly zero for $\Delta>\Delta^{*}\approx 1.4$ (see Fig. 2).
6.3.2 $p$-Néel state with $p=2$
Let us now discuss the cut protocol (see Fig. 1 (b)) for the $p$-Néel state with $p=2$. Our results are reported in Fig. 9. Panels (a)-(c) show $tDMRG$ data for the dynamics with the $XXZ$ chain Hamiltonian and $\Delta=1,2,4$. In the Figure we plot the local magnetization $\langle\sigma_{x}^{z}\rangle$ as a function of position $x$ and time $t$. For $\Delta=1$ (Fig. 9 (a)) two distinct signals are visible. The two “jets” correspond to quasiparticles with $n=1$, i.e., free ones, and to the two-particle bound states. The two-particle jet is due to the fact that for $p=2$ the bound-state content of the chain is enhanced as compared to $p=1$, and two-particle bound states are not confined at $\Delta=1$. Notice also that the bound states exhibit smaller group velocities as compared with the quasiparticles with $n=1$. The dashed lines reported in Fig. 9 (a) are the maximum bare velocities $v_{n,\mathrm{max}}^{\mathrm{b}}=\max_{\lambda}(v_{n}^{\mathrm{b}}(\lambda))$ (cf. (8)).
In Fig. 9 (b) we show data for $\Delta=2$. Crucially, the jet corresponding to the two-particle bound states is still visible. This is in agreement with the results of section 5. Indeed, for $p=2$ one should expect that the quasiparticles with $n=2$ are never confined. On the other hand, as it is clear from Fig. 2 (c), the three-particle bound states are confined for $\Delta\gtrsim 1.1$. We should mention that, however, as the density of quasiparticles with $n=3$ is “small”, it is numerically challenging to observe the three-particle bound states confinement. Finally, in Fig. 9 (c) we show results for $\Delta=4$. Now, the jet corresponding to $n=1$ is weak, in contrast with the case with $p=1$. This is due to the fact that for $p=2$ the density $\rho_{1}(\lambda)$ vanishes, reflecting that thermodynamic properties of the system are dominated by the bound states with $n=2$. On the other hand, for $\Delta=4$ the maximum group velocity of the two-particle bound states is $v_{2,\mathrm{max}}^{\mathrm{b}}\approx 0.1$, and their expansion is barely visible.
Finally, in Fig. 10 we report the projector $P_{x,\uparrow\uparrow\uparrow}$ after the expansion of the $p$-Néel state with $p=2$. In Fig. 10 (a) and (b) we show results for $\Delta=1$ and $\Delta=2$, respectively. While for $\Delta=1$ ballistic spreading is visible, for $\Delta=2$, the projector $P_{x,\uparrow\uparrow\uparrow}$ is “frozen”.
Figure 9: Releasing the “fattened” $p$-Néel state with $p=2$ in the vacuum. The figure shows $tDMRG$ data for the local magnetization $\langle\sigma_{x}^{z}\rangle$ as a function of $x$ and the time $t$. The different panels show results for different values of $\Delta$. For $\Delta=1$ and $\Delta=2$ the two jets correspond to the magnon-like excitations and the two-particle bound states.
Figure 10: Releasing the $p$-Néel state with $p=2$ in the vacuum. The figure shows $\langle P_{x,\uparrow\uparrow\uparrow}\rangle$ as a function of time and position $x$. Panel (a) and (b) show $tDMRG$ data for the $XXZ$ chain with $\Delta=1$ and $\Delta=2$, respectively. Notice that for $\Delta=2$ the quasiparticles with $n=3$ are confined in $A$.
Figure 11: Releasing the Néel state in the vacuum. Dynamics of the entanglement entropy. The curve shows $tDMRG$ data for a chain with $L=200$ sites. The size of region $A$ (see Fig. 1) is $\ell=20$. The plot shows the entropy $S$ of subsystem $B$ (see Fig. 1). Since subsystem $A$ includes the Néel part, the growth of $S$ with time is due to the entanglement between the excitations that leave the Néel part and those that are left behind. At long times $S$ measures the entanglement between the “jets” of excitations in Fig. 5 and Fig. 6. Notice the multistage growth of the entropy. At short times the growth is dominated by the entangled magnon-like excitations. The second slope visible in the figure is attributed to the two-excitations bound states. The last stage of the growth is due to the three-particle bound states.
7 Bound-state confinement and dynamics of the entanglement entropy
Here we show that bound-state confinement has striking signatures in the dynamics of the von Neumann entropy. We focus on the full expansion of the Néel state (see Fig. 1 (a)). Fig. 11 shows the von Neumann entropy $S$ of region $B$ (see Fig. 1 (a)). The dotted curve is the result for $\Delta=1$, whereas the full line shows the result for $\Delta=2$. First, for both values of $\Delta$ the von Neumann entropy saturates at long time. This reflects that the entanglement dynamics is due to the creation of pairs of entangled quasiparticles in region $A$ [29]. This is the content of the so-called quasiparticle picture [33, 28, 29, 34]. The two quasiparticles forming the entangled pairs travel with opposite group velocities. As they propagate, they entangle larger regions. Precisely, the von Neumann entropy of a region is proportional to the number of entangled pairs that are shared between the region and the rest. The contribution of a shared pair characterized by rapidities $(-\lambda,\lambda)$ is the contribution of that pair to the thermodynamic entropy of the $GGE$ describing the steady state after the quench from the homogeneous Néel state. The quasiparticle picture holds in the hydrodynamic limit $\ell,t\to\infty$ with the ratio $t/\ell$ fixed.
Let us focus on the results for $\Delta=1$ in Fig. 11. The initial linear growth up to $t\approx 20$ is due to the entangled pairs that are initially produced in $A$ and at later times leave region $A$, and hence start to be shared with the rest (see Fig. 1). The saturation of $S$ at long times reflects that region $A$ is finite, and, as a consequence the number of entangled pairs produced in $A$ is finite, and it is proportional to the size $\ell$ of $A$. For $\Delta=1$, $S$ is essentially constant for $t\gtrsim 60$. The saturation of the entanglement entropy is the key to reach times long as $t\approx 120$ with $tDMRG$, because after the entropy saturates the computational cost is mild. One should also observe that the saturation of $S$ happens “smoothly”. This is somewhat reminiscent of the fact that for $\Delta=1$ the bound states are not confined. Crucially, for $\Delta=2$, $S$ exhibits a multistage linear growth before saturating. This reflects the bound-state confinement. Specifically, the linear growth up to $t\approx 20$ is due to the quasiparticles with $n=1$. In this early time regime, the slope of the linear growth should be the same as in the vacuum expansion of a semi-infinite chain prepared in the Néel state. The prediction of the quasiparticle picture is [14, 16]
[…]
where $s_{1,\zeta=0}^{YY}$ is the Yang-Yang entropy [2] of the macrostate with $\zeta=0$ that describes the interface between region $A$ and region $B$ (see Fig. 1). $s_{n,\zeta}^{YY}$ is readily obtained from the particle densities $\rho_{n,\zeta}(\lambda)$ describing the $GGE$ as
$s_{n,\zeta}(\lambda):=\rho_{t,n,\zeta}\ln(\rho_{t,n,\zeta})-\rho_{n,\zeta}\ln(\rho_{n,\zeta})-\rho_{h,n,\zeta}\ln(\rho_{h,n,\zeta}),$ (49)
where the densities $\rho_{n,\zeta},\rho_{h,n,\zeta},\rho_{t,n,\zeta}$ are defined in section 3 (see also section 4). Notice that in (48) only the quasiparticles with $n=1$ contribute, reflecting that all the bound states are confined in $A$, and cannot contribute to the von Neumann entropy. Now, the $tDMRG$ data in Fig. 11 exhibit deviations from (48). This is expected because (48) holds in the hydrodynamic limit $t\to\infty$. Finally, we should mention that it is possible to characterize the trajectories [16] of the the generic entangled pairs produced in $A$, which are responsible of the entanglement growth. This, in principle, could allow to determine also the secondary slopes visible in Fig. 11.
8 Conclusions
We investigated bound-state confinement after trap-expansion experiments in the $XXZ$ spin chain at $\Delta>1$. We showed that after the expansion in the vacuum of quite generic initial states the bound states of the $XXZ$ chain remain confined in the initial region, provided that $\Delta$ is large enough. Upon lowering $\Delta$ the bound states are allowed to leave region $A$, i.e., they are not confined anymore. The condition ensuring confinement is the presence of a finite density of magnon-like excitations in the bulk of the initial region. Physically, bound-state confinement is a genuine consequence of interactions, i.e., the scattering, between the bound states and the magnons. Interestingly, if the size of the initial region is finite, at long times the density of magnons vanishes, as they leave the initial region, and bound-state confinement breaks down. Thus, for the expansion of finite regions, bound-state confinement is present only at intermediate times $t={\mathcal{O}}(\ell)$.
We investigated bound-state confinement after the expansion of the $p$-Néel states (see Fig. 1). In the limit $p\to\infty$ one recovers the domain wall initial state, and transport is suppressed. We showed that for small $p\lesssim 4$, bound-state confinement happens for the bound states of size $n>p$. At larger $p$, even the bound states with $n=p$ are confined, which is consistent with the absence of transport in the limit $p\to\infty$. Finally, we showed that bound-state confinement is reflected in a multi-stage linear growth of the entanglement entropy.
Let us now discuss some directions for future research. First, we investigated bound-state confinement in a lattice model. It would be interesting to understand whether confinement occurs for systems in the continuum, such as the attractive Lieb-Liniger gas. This would pave the way for the experimental verification of bound-state confinement, for instance with atom-chip experiments [35, 36] or trapped ions experiments [8]. Interestingly, bound-state confinement could be potentially used to prepare thermodynamic states with high bound-state content. Specifically, the protocol to enrich the bound state content would consists of two steps. In a first step one can expand the initial state as in Fig. 1 (a). Then, after the magnon-like excitation are “evaporated” one could trap the remaining excitations by applying an external potential. As the bound states are typically associated with lower thermodynamic entropy, bound-state confinement can also be used to prepare low-entropy states. It has been pointed out recently that bound-states could be robust agains long-range interactions [37, 8]. It would be interesting to understand whether long-range spin chains exhibit bound-state confinement as well. Finally, it would be important to clarify whether bound-state confinement is also present in integrable quantum circuits [38, 39, 40].
Acknowledgements
This study was carried out within the National Centre on HPC, Big Data and Quantum Computing - SPOKE 10 (Quantum Computing) and received funding from the European Union Next-GenerationEU - National Recovery and Resilience Plan (NRRP) – MISSION 4 COMPONENT 2, INVESTMENT N. 1.4 – CUP N. I53C22000690001. This work has been supported by the project “Artificially devised many-body quantum dynamics in low dimensions - ManyQLowD” funded by the MIUR Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN) Bando 2022 - grant 2022R35ZBF.
This work has been partially funded by the ERC Starting Grant 101042293 (HEPIQ).

### v2

@@
 Bound-state confinement persists until the density of unbound excitations remains finite in the bulk of $A$.
-Since region $A$ is finite, at asymptotically long times bound states are “liberated” after the “evaporation” of all the unbound excitations.
+Since region $A$ is finite, at asymptotically long times bound states are “liberated” after the “evaporation” of the unbound excitations.
 Fingerprints of confinement are visible in the space-time profiles of local spin-projection operators.
@@
 Integrable quantum many-body systems provide an ideal playground to discover exotic behaviors both at equilibrium and out-of-equilibrium.
-As recognized already by Bethe [1], one of the hallmarks of interacting many-body systems is the formation of multiparticle bound states.
-In spin chains they are formed by elementary magnon-like excitations.
+As recognized already by Bethe [1], one of the hallmarks of interacting many-body systems is the formation of stable multiparticle bound states.
 Bound states play a pivotal role in the equilibrium and out-of-equilibrium integrable systems, as they enter prominently in the Thermodynamic Bethe Ansatz [2] approach.
@@
 Furthermore, bound states are key ingredients behind exotic transport behaviors, such as superdiffusion [9].
-Very recently, they have been observed in quantum-computing platforms [10].
-Here we investigate the dynamics of bound states after trap-expansion experiments in the Heisenberg spin chain, which is a prototypical integrable system [2].
+Very recently, they have been observed in quantum-computing platforms [10, 11].
+We should mention that while bound states should be present also in generic, i.e., non integrable models, integrability ensures that they are stable excitations.
+Here we investigate the dynamics of bound states after quenches from inhomogeneous initial states in the Heisenberg $XXZ$ spin chain, which is a prototypical integrable system [2].
+An important property of integrable systems is that they possess stable quasiparticles, which are labelled by a set of complex numbers called rapidities [2].
+For instance, the quasiparticles of the Heisenberg spin chain resemble standard magnon excitations and bound states of magnons.
+In contrast with standard magnons, however, the excitations of Bethe Ansatz systems are interacting.
+Crucially, quasiparticles and their rapidities are preserved during the dynamics because integrability allows only elastic scattering.
+This property allows to understand the dynamics of local observables, and of entanglement [12], in terms of the ballistic propagation of the quasiparticles.
+This observation is at the heart of the so-called Generalized Hydrodynamics ($GHD$) approach [13, 14].
 Our out-of-equilibrium protocol is depicted in Fig.
@@
 The full chain then evolves under the $XXZ$ Hamiltonian.
-Crucially, the $XXZ$ chain, for anisotropy $\Delta\geq 1$ possesses bound states of an arbitrary number of elementary excitations.
-During the dynamics, as the state of region $A$ is not an eigenstate of the model, region $A$ undergoes a nontrivial evolution, whereas the remaining part of the chain does not evolve because the ferromagnetic state is eigenstate of the $XXZ$ Hamiltonian.
-The dynamics of region $A$ is understood in terms of the thermodynamic excitations [2] of the model.
-In the limit $\ell\to\infty$, at long times local observables in the bulk of region $A$ are described by a thermodynamic macrostate, which is a Generalized Gibbs Ensemble [11] ($GGE$).
-The initial inhomogeneity that is present at the interface between region $A$ and the rest (see Fig.
+Since quasiparticles are initially present only in region $A$, whereas the remaining part of the chain is in the vacuum state, our protocol is akin to the trap-expansion protocol routinely employed in cold atom experiments.
+For this reason, in the following we often refer to the protocol in Fig.
+1 as trap-expansion or expansion protocol.
+The initial state is not globally an eigenstate of the $XXZ$ model, implying that the system evolves with time.
+Specifically, the bulk of region $A$ undergoes a nontrivial evolution from time $t=0^{+}$, whereas the bulk of the remaining part of the chain initially does not evolve, because it is in the vacuum state.
+The dynamics of region $A$ is understood in terms of the quasiparticle excitations [2] of the model.
+For times tmuch-less-$t\ll\ell$, provided that $t$ is long enough for the system to reach local equilibrium, local observables in the bulk of $A$ are described by a thermodynamic macrostate, which is a Generalized Gibbs Ensemble [15] ($GGE$).
+Moreover, the initial inhomogeneity that is present at the interface between region $A$ and the rest (see Fig.
 1) “melts” during the dynamics.
-This is due to excitations leaving region $A$, where they are initially produced.
-At long times and large distances from the interface, the state of the system is described by a space-time dependent $GGE$, which can be obtained within the Generalized Hydrodynamics ($GHD$) approach [12, 13].
-In contrast with free-fermion systems, the thermodynamic excitations of the $XXZ$ chain are interacting.
-However, only elastic scattering is allowed, due to integrability.
-During the scattering the identity of the quasiparticles is preserved.
-The net effect of the scatterings is to dress the “bare” velocities of the quasiparticles.
-The dressed velocities depend on the space-time density of quasiparticles, and hence on the initial state.
-Crucially, the dressing can be so dramatic that the sign of the velocities change.
-Typically, this is more striking for the bound states, because their bare velocities (see section 5) are typically smaller than the magnons, implying that the bound states are more affected by scatterings.
-As we will show, this leads to bound-state confinement, meaning that the bound states remain trapped in region $A$ where they are produced (see Fig.
-1).
+This is due to the fact that the quasiparticles that initially are present only in $A$ spread in the rest of the chain.
+At long times, different regions around the edges of $A$ reach different equilibrium macrostates, giving rise to a space-time dependent $GGE$.
+This inhomogeneous $GGE$ can be obtained within the $GHD$ approach [13, 14].
+The main idea of $GHD$ (see section 4) is that the space-time dependent densities of quasiparticles satisfy a continuity equation, reflecting that quasiparticles do not decay and preserve their rapidity.
+Moreover, since the scattering between different quasiparticles is elastic, its net effect is to dress the “bare” velocities of the quasiparticles.
+In turn, the dressed velocities, which appear in the continuity equation, depend on the space-time-dependent densities.
+This means that the continuity equation has to be solved self-consistently.
+Crucially, the effect of the scatterings can be so dramatic that the sign of the velocities changes.
+Typically, this is more striking for the bound states, because their bare velocities (see section 5) are generically smaller than those of the magnons.
+As we will show, this leads to bound-state confinement, meaning that the transport of bound states from region $A$ to region $B$ is completely suppressed at ballistic level, i.e.
+at long times $t$ and large distances $x$ from the interface between $A$ and the rest, with the ratio $\zeta:=x/t$ fixed.
+In the rest of the paper, the confinement of bound states must always be intended in this limit.
+Intuitively, as the bound states approach the interface between $A$ and the rest, they scatter with the magnons, and, as a consequence, they are “pushed back” in $A$.
+This leads to an accumulation of bound states in $A$.
+From now on we refer to this behavior, i.e., to the fact that the bound states remain trapped in region $A$ at ballistic level, as bound state confinement.
 This was observed already in Ref.
@@
 […]
-Here by combining the time-dependent Density Matrix Renormalization Group [19] ($tDMRG$) with state-of-the-art integrability methods [20, 21], we provide a thorough investigation of bound-state confinement in the $XXZ$ chain at anisotropy $\Delta>1$.
-We show that for generic initial states of region $A$, multiparticle bound states remain confined in $A$.
-Precisely, this happens provided that interactions, i.e., the anisotropy $\Delta$, are strong enough, and there is a finite density of magnon-like quasiparticles, i.e., unbound ones, in region $A$.
+Finally, we should mention that bound-state confinement gives rise to the so-called quantum distillation [21, 22, 23].
+It has been observed in Ref.
+[21] that after the sudden expansion of interacting fermions if the initial state has a strong admixture of double occupancies, the doublons group together forming a band-insulating state trapped at the center of the system.
+This state has low-entropy.
+This protocol has been investigated in cold-atom experiments in [22, 23].
+In this work we provide a thorough investigation of bound-state confinement in the $XXZ$ chain at anisotropy $\Delta>1$ by combining the time-dependent Density Matrix Renormalization Group [24] ($tDMRG$) with state-of-the-art integrability methods [25, 26].
+For $\Delta>1$ the $XXZ$ chain possesses bound states of an arbitrary number of magnons.
+We show that all the multiparticle bound states remain confined in $A$, provided that interactions, i.e., the anisotropy $\Delta$, are strong enough, and there is a finite density of magnonic quasiparticles, i.e., unbound ones, in region $A$.
 Confinement is present at large $\Delta$, and it can be understood perturbatively in the limit $\Delta\to\infty$.
-Upon lowering $\Delta$ there is a threshold below which the lighter bound states, i.e., formed by two elementary quasiparticles, are not confined.
+Upon lowering $\Delta$ there is a threshold below which the lighter bound states, i.e., formed by two magnons, are not confined.
 This means that two-particle bound states are allowed to leave region $A$ (see Fig.
@@
 This is in agreement with the results of Ref.
-[14] and [15] for the case of Néel and Majumdar-Ghosh initial states.
+[16] and [17] for the specific cases of Néel and Majumdar-Ghosh initial states.
 Moreover, for typical initial states of region $A$ there is a hierarchical emission of bound states, meaning that heavier bound states are “liberated” at lower $\Delta$.
-While confinement can be understood perturbatively, the de-confinement happens near $\Delta=1$, i.e., away from the perturbative regime.
-This means that the deconfinement pattern of the different bound states depends on the details of the initial state.
-Finally, it is important to stress that bound-state confinement happens at ballistic level, i.e., in the hydrodynamic limit $x,t\to\infty$ with $\zeta:=x/t$, with $x$ the distance from the interface between $A$ and $B$ (see Fig.
-1).
-Crucially, since the size $\ell$ of $A$ (see Fig.
-1) is finite, at long times tmuch-greater-$t\gg\ell$ the density of magnon quasiparticles in $A$ vanishes.
-In turn, this implies that after the “evaporation” of all the magnons from region $A$, the system cannot support bound-state confinement.
-Thus, even for large $\Delta$ bound states cannot remain confined at arbitrarily long times (see Fig.
-1 (c)) if region $A$ is finite.
-Still, the short-time behavior of the spin projectors $P_{\uparrow^{\otimes n}}$ on $n$ consecutive sites are reminiscent of confinement.
+While confinement can be understood perturbatively, the deconfinement happens near $\Delta=1$, i.e., away from the perturbative regime.
+This means that the deconfinement pattern of the different bound states depends on the details of the initial state, and it cannot be addressed perturbatively.
+Moreover, it is important to stress that if the size $\ell$ of region $A$ is finite, confinement can be observed only at intermediate times tmuch-less-$t\ll\ell$.
+As it will explained in Section 5, intermediate time tmuch-less-$t\ll\ell$ are both a requirement for the applicability of the bipartite protocol in GHD and to ensure that the central region has a finite density of excitations.
+Indeed, at times tsimilar-$t\sim\ell$ the density of magnons in the bulk of $A$ starts to be depleted.
+This implies that the effect of the scattering between magnons and bound states, and hence the dressing of bound-state velocity, weakens.
+At long-enough times, after the magnons “evaporated” from $A$, deconfinement of the bound states happens (see Fig.
+1 (c)).
 Following Ref.
-[22], we use $P_{\uparrow^{\otimes n}}$ as witness for $n$-particles bound-states.
-We observe that if interactions are not strong enough to support confinement in the hydrodynamic limit, at short times the projectors exhibit a “halo” signal around the interface between $A$ and its complement.
-On the other hand, if confinement is present, the projectors exhibit a “bottleneck” at the interface between $A$ and $B$ before the “deconfinement” of the bound states happens at longer times.
-Finally, we confirm the confinement scenario by considering a “cut” quench (see Fig.
-1 (b)) in which the expansion is stopped after a time $t^{*}$.
+[27], we employ the space-time profiles of the spin projectors $P_{x,\uparrow^{\otimes n}}$ to witness bound-state confinement.
+$P_{x,\uparrow^{\otimes n}}$ are sensitive to the presence of a string of $n$ up spins on consecutive sites starting from $x$.
+We also confirm the confinement scenario by considering a “cut” quench (see Fig.
+1 (b)) in which the expansion is stopped after a time $t^{*}$, and for $t>t^{*}$ only part $B$ of the chain evolves.
 The stopping time is chosen such that the density of magnons in $A$ is finite, i.e., before the bound-state deconfinement starts.
-For $t>t^{*}$ only part $B$ of the chain is further evolved.
-If interactions are strong enough to permit bound-state confinement, the projectors $P_{\uparrow^{\otimes n}}$ do not show any signal at $t>t^{*}$ in region $B$.
-In this work we focus on the vacuum expansion of the $p$-Néel states, which are “fattened” version of the Néel state.
+If interactions are strong enough to permit bound-state confinement, the expectation values of the projectors $P_{\uparrow^{\otimes n}}$ do not show any signal at any $t>t^{*}$ in region $B$, since there are no bound states outside of $A$.
+In this work, we focus on the expansion of the $p$-Néel states, which are “fattened” version of the Néel state.
 They are defined by repetition of a unit cell containing $p$ up spins followed by $p$ down spins (see Fig.
 1 (a)).
-In the limit $p\to\infty$ the $p$-Néel states converge to the domain-wall state.
-Interestingly, for $\Delta>1$ transport is suppressed after the quench from the domain-wall state [23].
-At $\Delta=1$ transport is neither ballistic nor diffusive, but a superdiffusive dynamics is expected [24].
 An interesting feature of the $p$-Néel state is that by changing $p$, one can change the bound-state content.
-Precisely, states with larger $p$ contain more bound states.
-For $p=1$ one recovers the standard Néel state, which is somewhat special because for $p=1$ the $GGE$ describing the bulk of $A$ can be determined analytically [25].
-This is not the case for $p>2$ [26], although the $GGE$ can be obtained at modest numerical cost by using the results of Ref.
+Precisely, states with larger $p$ contain larger bound states.
+We observe the following pattern for bound state confinement.
+If region $A$ (see Fig.
+1) is prepared in the Néel state (with $p=1$), at large enough $\Delta\gtrsim 1.4$, all the bound states are confined, as observed in Ref.
 [17].
-If region $A$ (see Fig.
-1) is prepared in the Néel state (with $p=1$), at large enough $\Delta\gtrsim 1.4$, bound states are confined, as observed in Ref.
-[15].
-Interestingly, for $2\leq p\lesssim 4$, only bound states with size $n>p$ are confined, whereas for $p\gtrsim 4$ bound states of size $n=p$ are also confined.
-Since bound states with $p=n$ dominate the behavior of local observables after the quench from the $p$-Néel states, this is compatible with the absence of transport in the $p\to\infty$ limit.
-Finally, we show that bound-state confinement has striking consequences on the dynamics of entanglement entropy.
+Interestingly, for $42\leq p<$ 4, only bound states with size $n>p$ are confined, whereas for $p\geq 4$ bound states of size $n=p$ are also confined.
+Finally, we show that bound-state confinement and deconfinement are sharply reflected in the dynamics of the entanglement entropy.
 First, within the quasiparticle picture for entanglement dynamics [28, 29, 12], entanglement is created in region $A$.
@@
 In section 2 we introduce the $XXZ$ spin chain and the out-of-equilibrium protocols.
-In section 3 we review the Thermodynamic Bethe Ansatz of the model.
+In section 3 we review the Thermodynamic Bethe Ansatz ($TBA$) of the model.
 In section 4 we discuss the Generalized Hydrodynamics ($GHD$) approach.
@@
 In particular, in section 5.1 we provide numerical $GHD$ results confirming bound-state confinement after the expansion of the $p$-Néel states.
-In section 6, we discuss numerical $tDMRG$ results for the trap-expansion dynamics.
+In section 6, we discuss numerical $tDMRG$ results for the expansion dynamics.
 In particular, in section 6.1 we discuss the validity of the $GHD$ approach.
@@
 We conclude and discuss future directions in section 8.
+In A we discuss the $TBA$ derivation of the $GGE$ for homogeneous quenches.
+In B we discuss the convergence of the $tDMRG$ method.
 2 Model and quantum transport protocols
@@
 The generic $p$-Néel state reads as
-$|N_{p}\rangle:=|\uparrow^{\otimes p}\downarrow^{\otimes p}\rangle^{\otimes L/(2p)},$ (2)
+$|N_{p}\rangle:=|\uparrow^{\otimes p}\downarrow^{\otimes p}\rangle^{\otimes\ell/(2p)},$ (2)
 where $\otimes$ denotes the tensor product.
+In the limit $p\to\infty$ the $p$-Néel states converge to the domain-wall state.
+Interestingly, for $\Delta>1$ transport is suppressed after the quench from the domain-wall state [30].
+At $\Delta=1$ transport is neither ballistic nor diffusive, but a superdiffusive dynamics is expected [31].
 […]
@@
 As anticipated, in both protocols $(a)$ and $(b)$ of Fig.
-1, at $t>0$ part $A$ of the chain “melts” because $|N_{p}\rangle$ is not an eigenstate of (1).
+1, since $|N_{p}\rangle$ is not an eigenstate of (1), the part $A$ of the chain starts to “melt” from arbitrarily small times $t=\epsilon>0$.
 First, the dynamics of an infinite homogeneous chain prepared in $|N_{p}\rangle$ gives rise to a Generalized Gibbs Ensemble [15] ($GGE$) with extensive thermodynamic entropy.
@@
 Oppositely, at the interface between region $A$ and the rest, the dynamics leads to the formation of a nontrivial profile.
-Specifically, in the long-time limit, provided that tmuch-less-$t\ll\ell$, to ensure that the central region has a finite density of excitations, ballistic transport sets in.
+Specifically, in the long-time limit ballistic transport sets in.
+As it will be explained in Section 5, provided that tmuch-less-$t\ll\ell$ to ensure that the central region has a finite density of excitations, $GHD$ description is valid.
 This means that expectation values of generic local observables ${\mathcal{O}}(x,t)$ (with $x$ measured from the interface between $A$ and rest) are functions of $\zeta:=x/t$, and not of space and time separately.
@@
 Crucially, since the $XXZ$ chain is interacting, for sufficiently large $\Delta>1$ transport of bound states (of arbitrary size) of elementary excitations from region $A$ to $B$ is inhibited in the ballistic regime.
-This means that the expectation values ${\mathcal{O}}(\zeta)$ for $\zeta>0$ are fully determined by magnon-like excitations.
+This means that the expectation values ${\mathcal{O}}(\zeta)$ for $\zeta>0$ are fully determined by magnonic excitations.
 This bound-state confinement happens for both the protocols of Fig.
-1.
+1 at times tmuch-less-$t\ll\ell$.
 As already stressed, for the protocol in Fig.
-1 $(a)$ the dynamics at asymptotically long times leads to zero density of magnon-like excitations in the bulk of $A$.
+1 $(a)$ the dynamics at asymptotically long times leads to zero density of magnonic excitations in the bulk of $A$.
 As we will discuss in section 5 this will lead to the deconfinement of bound states.
@@
 (2)), it is necessary to construct the $GGE$ describing the steady state after the quench from the homogeneous $p$-Néel states.
-Here we provide a self-contained discussion of the derivation of the $GGE$ for a generic pre-quench initial state.
-We follow Ref.
-[20, 21].
-We employ the framework of the Thermodynamic Bethe Ansatz [2] ($TBA$).
+A self-contained discussion of the derivation of the $GGE$ for generic prequench initial states is reported in A.
+Here, we employ the framework of the Thermodynamic Bethe Ansatz [2] ($TBA$).
 The eigenstates of the $XXZ$ chain are identified by a set of complex numbers $\lambda_{j}$, which are called rapidities.
@@
 The corresponding rapidities form string patterns in the complex plane [1].
-Then, any thermodynamic macrostate of the system is characterized by a set of rapidity densities $\{\rho_{j}(\lambda)\}_{j=1}^{\mathcal{N}}$, where the number of families ${\mathcal{N}}$ depends on the model under consideration.
-For the $XXZ$ chain at $\Delta>1$ one has ${\mathcal{N}}\to\infty$, meaning that bound states of arbitrary sized are present.
+Then, any thermodynamic macrostate of the system is characterized by a set of rapidity density $\{\rho_{j}(\lambda)\}_{j=1}^{\mathcal{N}}$, where ${\mathcal{N}}$ depends on the model under consideration and where, for $j>1$, $\rho_{j}=\rho_{j}(\mathrm{Re}(\lambda))$ with $\mathrm{Re}(\lambda)$ the center of the string.
+From now on, $\lambda$ will denote the string center, and we will drop the real part in $\mathrm{Re}(\lambda)$.
+For the $XXZ$ chain at $\Delta>1$ one has ${\mathcal{N}}\to\infty$, meaning that bound states of arbitrary size are present.
 Similar to the densities $\rho_{j}(\lambda)$, one defines the hole densities $\rho_{h,j}(\lambda)$.
@@
 $v^{b}_{n}=\frac{\partial\varepsilon_{n}(\mu)}{\partial\mu}\left(\frac{\partial p_{n}}{\partial\mu}\right)^{-1},$ (8)
-where the energy $\varepsilon_{n}$ and the quasimomentum $p_{n}$ are defined as
-$\varepsilon_{n}(\mu)=\frac{\sinh(\eta)\sinh(n\eta)}{\cosh(n\eta)-\cos(2\mu)},\quad p_{n}(\mu)=-i\sum_{j=1}^{n}\ln\left[\frac{\sin\left(\mu+i\frac{\eta}{2}(n+1-2j)+i\frac{\eta}{2}\right)}{\sin\left(\mu+i\frac{\eta}{2}(n+1-2j)-i\frac{\eta}{2}\right)}\right].$ (9)
+where the bare energy $\varepsilon^{b}_{n}$ and the quasimomentum $p^{b}_{n}$ are defined as
+$\varepsilon^{b}_{n}(\mu)=\frac{\sinh(\eta)\sinh(n\eta)}{\cosh(n\eta)-\cos(2\mu)},\quad p^{b}_{n}(\mu)=-i\sum_{j=1}^{n}\ln\left[\frac{\sin\left(\mu+i\frac{\eta}{2}(n+1-2j)+i\frac{\eta}{2}\right)}{\sin\left(\mu+i\frac{\eta}{2}(n+1-2j)-i\frac{\eta}{2}\right)}\right].$ (9)
 Notice that Eq.
@@
 Crucially, to uniquely identify a thermodynamic macrostate, one has to determine both the particle and the hole densities.
-In the following, instead of $\rho_{n}(\lambda)$ and $\rho_{h,n}(\lambda)$ we will often consider the ratios $\eta_{n}(\lambda)$ and $\vartheta_{n}(\lambda)$ defined as
+The information about the expectation values of local and quasilocal conserved quantities over the initial state allows to fix the $TBA$ densities identifying the $GGE$ that describes the steady state after the quench [25, 26].
+In particular, the hole density is fixed by the initial-state expectation values of quasilocal charges $X_{n}(\lambda)$ by the relationship
+$\rho_{h,n}(\lambda)=a_{n}-X_{n}\left(\lambda+i\frac{\eta}{2}\right)-X_{n}\left(\lambda-i\frac{\eta}{2}\right),$ (10)
+where $a_{n}$ is the same as in  (5) and $\eta$ is defined in  (6).
+An explicit expression for $X_{n}$ (49) is given in A.
+Hence, Eq.
+(3) and Eq.
+(10) form a system of equations that can be solved to univocally determine $\rho_{n}(\lambda)$ and $\rho_{h,n}(\lambda)$.
+In the following, instead of $\rho_{n}(\lambda)$ and $\rho_{h,n}(\lambda)$, we will often consider the ratios $\eta_{n}(\lambda)$ and $\vartheta_{n}(\lambda)$ defined as
 $\eta_{n}:=\frac{\rho_{h,n}}{\rho_{n}},\quad\vartheta_{n}(\lambda):=\frac{\rho_{n}(\lambda)}{\rho_{t,n}(\lambda)}.$ (11)
-We now show that the information about the expectation values of local and quasilocal conserved quantities over the pre-quench initial state allows to fix uniquely the $TBA$ densities identifying the $GGE$ that describes the steady state after the quench.
-A central object in the Algebraic Bethe Ansatz treatment [30] of the $XXZ$ chain is the spin-$j/2$ Lax operator $L_{j}(\mu)$ defined as
+Finally, the expectation value of any local observable $\mathcal{O}$ on a generic thermodynamic macrostate $GGE$ is computed from the quasiparticle densities $\rho_{n}(\lambda)$ as [2]
+[…]
+where $o_{n}(\lambda)$ depends on the observable under consideration. For instance, for the density of quasiparticles one has $o_{n}(\lambda)=n$.
+4 Generalized Hydrodynamics ($GHD$) for quenches from inhomogeneous initial states
+Here we introduce the framework of Generalized Hydrodynamics [14, 13] ($GHD$) to describe the expansion dynamics (see Fig.
+1).
+Generalized Hydrodynamics is an extension of standard hydrodynamics to integrable systems.
+Integrable systems possess an extensive number of conserved charges.
+A crucial consequence of this property is that the many-body scattering is factorized into two-body elastic processes.
+Scattering processes preserve the bound states and the rapidities.
+This means that quasiparticles (i.e., magnons and bound states of magnons) are stable.
+Also the fact that the total number of quasiparticles for any given value of the rapidity is preserved suggests that the spatial densities of quasiparticles satisfy a continuity equation, which is the main equation of $GHD$.
+In the standard $GHD$ setup, two semi-infinite homogeneous but different initial states are joined together at $t=0$ at the origin $x=0$.
+At $t>0$ the full system evolves under a homogeneous integrable Hamiltonian.
+This is the so-called bipartite quench protocol [19].
+Notice that although the setup of the bipartite quench is different from the one described in Fig.
+1 (a), the two protocols give the same dynamics provided that $\ell$ (see Fig.
+1) is large enough to ensure that the bulk of region $A$ is in the hydrodynamic limit and $1\ll v_{\mathrm{max}}t\ll\ell$ is large, but short enough to neglect the fact that region $A$ is finite.
+Here $v_{\mathrm{max}}$ is the maximum velocity in the system, i.e, $v_{\mathrm{max}}=\max_{n,\zeta,\lambda}v_{n,\zeta}(\lambda)$.
+Under these conditions the dynamics around the two edges of region $A$ happen independently and coincide with the dynamics in the bipartite protocol.
+For typical initial states in the bipartite setup, at long times and large distances with the ratio $\zeta:=x/t$ fixed, the behavior of the system is described by a $\zeta$-dependent $GGE$.
+This $GGE$ is fully characterized by a set of $TBA$ densities $\rho_{n,\zeta}(\lambda)$ and $\rho_{h,n,\zeta}(\lambda)$.
+The $TBA$ densities satisfy the $TBA$ equations (3) for any $\zeta$ and the $GHD$ equation, which reads as [13, 14, 19]
+$[\zeta-v_{n,\zeta}(\lambda)]\partial_{\zeta}\vartheta_{n,\zeta}(\lambda)=0,\quad\mathrm{with}\,\zeta:=x/t.$ (13)
+Here $\vartheta_{n,\zeta}$ are the filling functions (cf. (14)) defined as
+$\vartheta_{n,\zeta}(\lambda)=\frac{\rho_{n,\zeta}(\lambda)}{\rho_{t,n,\zeta}(\lambda)},$ (14)
+with $\rho_{t,n,\zeta}=\rho_{n,\zeta}+\rho_{h,n,\zeta}$ the total density. In (13), $v_{n,\zeta}$ are the dressed velocities, which are defined as
+$v_{n,\zeta}=\frac{[\partial_{\lambda}\varepsilon_{n,\zeta}]^{\mathrm{dr}}(\lambda)}{[\partial_{\lambda}p_{n,\zeta}]^{\mathrm{dr}}(\lambda)},$ (15)
+where $\varepsilon_{n}(\lambda)$ and $p_{n}(\lambda)$ are respectively the bare energy and quasimomentum of the quasiparticles, defined in (9).
+Given a generic “bare” function of the rapidity $\tau^{\mathrm{b}}_{n}(\lambda)$, its dressed version $\tau_{n,\zeta}^{\mathrm{dr}}$ is obtained by solving the integral equation as [32, 33]
+[…]
+where $\partial_{\lambda}\Theta_{n,m}(\lambda)$ is defined in (7), and the filling function $\vartheta_{n,\zeta}(\lambda)$ is given in (14).
+It is clear from the definition of the momentum $p_{n}$ as [2]
+$[\partial_{\lambda}p_{n,\zeta}(\lambda)]^{\rm dr}=2\pi\rho_{t,n,\zeta}(\lambda),$ (17)
+that the dressing of $\partial_{\lambda}p_{n}$ is obtained by considering $\tau^{\mathrm{dr}}_{n,\zeta}=2\pi\rho_{t,n,\zeta}(\lambda)$ in (16).
+To solve (13) one has to provide the boundary conditions at $\zeta=\pm\infty$.
+Let us focus on the bipartite protocol denoting by $\vartheta_{n,L}$ and $\vartheta_{n,R}$ the two thermodynamic macrostates, i.e., the $TBA$ filling functions, describing the steady state arising at infinite time after the quench from the left and right homogeneous states.
+The functions $\vartheta_{n,L},\vartheta_{n,R}$, provide the boundary conditions in (13).
+Before proceeding, one can observe that Eq.
+(13) can be solved for each $\zeta$ separately.
+It is straightforward to check that the solution of (13) can be written as
+$\vartheta_{n,\zeta}(\lambda)=\Theta_{\mathrm{H}}(v_{n,\zeta}-\zeta)[\vartheta_{n,L}(\lambda)-\vartheta_{n,R}(\lambda)]+\vartheta_{n,R}(\lambda).$ (18)
+where $\Theta_{\mathrm{H}}(x)$ is the Heaviside step function.
+The dressed velocity $v_{n,\zeta}$ in (18) is obtained by using (16) and (15) (17).
+Notice that since $v_{n,\zeta}$ depends on $\vartheta_{n,\zeta}$ via (16), Eq.
+(18) is not an explicit solution of (13).
+To obtain $\vartheta_{n,\zeta}$ one can use an iterative strategy, starting from (18) fixing $v_{n,\zeta}=v^{b}_{n}$, with $v^{b}_{n}$ the “bare” velocities (cf.
+(8)).
+Then, one substitutes Eq.
+(18) in (16) to obtain the dressed $\partial_{\lambda}\varepsilon_{n,\zeta}$.
+By using the $TBA$ equation (3) one obtains $\rho_{t,n,\zeta}$ and the new dressed velocities via (15).
+Finally, these steps are iterated until convergence.
+In the following we focus on the macrostate with $\zeta=0$, which describes the interface between the two chains.
+We also consider the situation in which the right chain is prepared in the vacuum state, while the left chain is prepared in the same product state defined in the region A (see Fig.
+1).
+This corresponds to $\vartheta_{n,R}=0$ for any $n$.
+Now, Eq.
+(18) becomes
+$\vartheta_{n,0}(\lambda)=\Theta_{\mathrm{H}}(v_{n,0})\vartheta_{n,L}(\lambda).$ (19)
+In the following we neglect the subscript $\zeta$ in $\vartheta_{n,\zeta}$ and $v_{n,\zeta}$ since we always consider $\zeta=0$.
+5 Bound-state confinement after trap-expansion experiments
+Here we discuss the confinement of the bound states for the expansion protocol of Fig.
+1 for generic initial state of part $A$.
+The state of $A$ corresponds to fixing $\vartheta_{n,L}$ in (19).
+Let us summarize our main results.
+We show that for moderately large $\Delta$ ballistic transport of all the bound states is inhibited, for quite generic initial states of part $A$.
+Specifically, in the hydrodynamic limit only quasiparticles with $n=1$, i.e., unbound quasiparticles, are present at $\zeta\geq 0$.
+This means that all the bound states with $n>1$ remain confined within region $A$ in Fig.
+1 (a).
+In the following, we will employ the results of section 4, which are valid for quenches from bipartite initial states.
+Again, although our initial state is tripartite, for times tmuch-less-$t\ll\ell$, the bulk of $A$ is not affected by the presence of region $B$.
+Thus, the dynamics around the edges of $A$, which is relevant to address confinement, are exactly the same as the dynamics that one has in the bipartite protocol.
+As a consequence, for tmuch-less-$t\ll\ell$ and $|x|\ll\ell$ the state of the system is well approximated by a $\zeta$-dependent GGE (see Section 6.1 for some numerical checks).
+At times $t\approx\ell$ when information about the interface between $A$ and the rest reaches the bulk of $A$ one could still apply $GHD$ although one should take into account that the bulk of $A$, and hence the filling functions $\vartheta_{n,L}$, evolve with time.
+Bound-state confinement can be understood perturbatively in the large $\Delta$ limit, and it relies on the behavior of the scattering phases (7).
+Indeed, generic initial states of region $A$ that give rise to $TBA$ macrostates with ${\mathcal{O}}(1)$ density of quasiparticles with $n=1$ will lead to confinement of all the bound states with $n\geq 2$.
+Although it is established perturbatively, bound-state confinement survives up to $\Delta\approx 1$.
+Precisely, for the Néel state bound state confinement survives up to [17] $\Delta=\Delta^{*}\approx 1.4$.
+Upon lowering $\Delta$ below $\Delta^{*}$, bound states with $n=2$ start to leave region $A$.
+For $\Delta\ll\Delta^{*}$ bound states with $n>2$ are not confined.
+However, while the threshold $\Delta^{*}$ for confinement of the two-particle bound states is understood perturbatively, this is not the case for the larger bound states.
+Indeed, the deconfinement of bound states with $n>2$ happens in the nonperturbative regime at $\Delta\approx 1$, and is sensitive to the details of the initial state.
+Bound-state confinement manifests itself in the fact that the dressed velocities $v_{n}(\lambda)$ at $\zeta=0$ are negative for any $\lambda$ for $n>1$.
+To proceed, let us assume that only bound states with $n>n_{a}$, for some $n_{a}>1$, are confined.
+This implies that Eq.
+(16) for $\tau_{n}=\partial_{\lambda}\varepsilon_{n}$ becomes
+$\partial_{\lambda}\varepsilon^{\mathrm{dr}}_{n}(\lambda)=\partial_{\lambda}\varepsilon^{\mathrm{b}}_{n}(\lambda)-\sum_{m=1}^{n_{\mathrm{a}}}\int^{\frac{\pi}{2}}_{-\frac{\pi}{2}}\!\!\frac{d\mu}{2\pi}\partial_{\lambda}\Theta_{n,m}(\lambda-\mu)\vartheta_{m,L}(\mu)\Theta_{\mathrm{H}}(\partial_{\mu}\varepsilon_{m}^{\mathrm{dr}}(\mu))\partial_{\mu}\varepsilon_{m}^{\mathrm{dr}}(\mu),$ (20)
+where $n_{\mathrm{a}}$ is such that $\partial_{\lambda}\varepsilon^{\mathrm{dr}}_{n}(\lambda)<0$ for any $\lambda$ and for any $n>n_{\mathrm{a}}$.
+Notice that in (20) we replaced $v_{n,0}$ with $\partial_{\mu}\varepsilon_{n}^{\mathrm{dr}}$ in the argument of the Heaviside theta function.
+This is possible because $\rho_{t,n}(\lambda)$ (cf.
+(15)) is positive.
+Also, notice that in the integrand in (20) one has that $\partial_{\lambda}\Theta_{n,m}$ and $\vartheta_{n,L}$ are both positive functions.
+For (20) to be consistent with confinement of the bound states with $n>n_{a}$, we require that the solutions of (20) satisfy the condition
+$\partial_{\lambda}\varepsilon_{n}^{\mathrm{dr}}(\lambda)<0,\forall n>n_{\mathrm{a}},\forall\lambda\,.$ (21)
+Clearly, while the dressing (16) for the $XXZ$ with $\Delta>1$ is a system of infinite equations since $n\in[1,\infty)$, the system (20) is finite.
+Still, given the solutions of (20), one has to check the validity of (21) for any $n>n_{a}$.
+Before proceeding, it is interesting to observe that not all the species of quasiparticles can be confined at the same time, at least in the hydrodynamic limit.
+Indeed, if all the quasiparticles are confined Eq.
+(20) implies that $\partial_{\lambda}\varepsilon_{n}^{\mathrm{dr}}(\lambda)=\partial_{\lambda}\varepsilon_{n}^{\mathrm{b}}(\lambda)$.
+However, $\partial_{\lambda}\varepsilon_{n}^{\mathrm{b}}(\lambda)$ is not negative for all values of $\lambda$ (cf.
+(9)), which is in contraddiction with the assumption that all the quasiparticles are confined.
+This implies that the suppression of ballistic transport observed for instance in the domain-wall quench [30] cannot be attributed to the bound-state confinement.
+The conditions under which the systems (20) and (21) admit a solution can be derived in the large $\Delta$ limit.
+First, let us observe that $0\leq\vartheta_{n,L}(\lambda)\leq 1$ for any $\lambda$.
+Moreover, in the large $\Delta$ limit, at the leading order in powers of $\Delta$, one has that $\partial_{\lambda}\Theta_{n,m}(\lambda-\mu)={\mathcal{O}}(1)$ for any $n,m$.
+Precisely, for large $\Delta$ one has that
+$\partial_{\lambda}\Theta_{n,m}(\lambda)\to\frac{1}{\pi}(n+m-1),\quad\mathrm{for}\,\,n,m\neq 1,$ (22)
+and $\partial_{\lambda}\Theta_{n,m}(\lambda)\to(n+m)/\pi$ otherwise.
+Finally, it is straightforward to check that the velocities $v_{n}^{\mathrm{b}}$ (cf.
+(8)) at the leading order in the large $\Delta$ limit are
+$v_{n}^{\mathrm{b}}(\lambda)=2z^{n-1}\sin(2\lambda)+{\mathcal{O}}(z^{n+1}),\quad\mathrm{with}\,\,z:=e^{-\eta}=\Delta-\sqrt{\Delta^{2}-1}\,.$ (23)
+Notice that at the leading order in $z$, $v_{n}^{\mathrm{b}}(\lambda)=\partial_{\lambda}\varepsilon_{n}^{\mathrm{b}}(\lambda)$, with $\varepsilon_{n}^{\mathrm{b}}(\lambda)$ the bare energy in (9).
+For large $\Delta$ one has $z\approx 1/(2\Delta)$.
+Moreover, unlike the dressed velocities, the bare ones do not depend on the pre-quench initial state, so (23) holds irrespective of the initial state of $A$ (see Fig.
+1).
+Crucially, while the bare velocities of the quasiparticles with $n=1$ are ${\mathcal{O}}(1)$, the bound-state velocities, in the regime $\Delta>1$, decay exponentially with the bound-state size $n$.
+This hierarchy in the bare velocities is responsible for the bound-state confinement.
+Let us consider the case with $\vartheta_{1,L}(\lambda)={\mathcal{O}}(1)$, i.e., the situation in which in region $A$ there is a finite density of unbound quasiparticles.
+This happens for typical states, such as the Néel state and the Majumdar-Ghosh state [17].
+Let us also assume that $n_{\mathrm{a}}=1$, i.e., bound states with $n\geq 2$ remain confined within $A$ during the expansion dynamics.
+Now, let us consider Eq.
+(20) with $n=2$.
+It is clear that the second term in the right-hand side dominates for large $\Delta$ because $\partial_{\lambda}\varepsilon_{n}^{\mathrm{b}}={\mathcal{O}}(z)$ for $n=2$, whereas the second term in (20) is ${\mathcal{O}}(1)$.
+Clearly, the same argument applies also for $n>2$, implying that all the bound states with $n\geq 2$ are confined for sufficiently large $\Delta$.
+Interestingly, as the only dependence on $\lambda$ in the second term in (20) is in $\Theta_{n,m}$, which becomes independent of $\lambda$ for large $\Delta$, the dressed velocities of the confined bound states become “flat” as functions of $\lambda$.
+Upon lowering $\Delta$, the two terms in (20) become of the same order, and quasiparticles with $n=2$ start to be transported in $B$ (see Fig.
+1).
+Notice that it is difficult to characterize analytically the confinement scenario for $\Delta\to 1$.
+Typically, we observe that there is a “cascade” effect.
+Precisely, upon lowering $\Delta$, transport of bound states with progressively larger $n$ become allowed (see Section 5.1).
+We now illustrate two situations in which the system can escape bound-state confinement.
+The key observation is that bound-state confinement relies on the fact that at any time during the expansion the bulk of $A$ is described by $\vartheta_{1,L}={\mathcal{O}}(1)$.
+This condition remains true at arbitrarily long times if region $A$ (cf.
+Fig.
+1) is infinite.
+Oppositely, if region $A$ is finite, at long times tmuch-greater-$t\gg\ell$ the bulk of $A$ starts to be depleted and the filling functions $\vartheta_{n,L}$ vary with time.
+One could still apply (20) after taking into account the time dependence in $\vartheta_{n,L}$.
+The main effect is that the second term in (20) becomes smaller.
+Indeed, one should observe that in Eq.
+(20), $\vartheta_{1,L}\Theta_{H}(\partial_{\mu}\varepsilon^{\mathrm{dr}}_{1}(\mu))$ accounts for the scattering with the magnons arriving at $\zeta=0$ from the bulk of $A$.
+Now, the fact that at long times the bulk of $A$ cannot supply the magnons, because they left $A$, suggests that $\vartheta_{1,L}$ should be small.
+The reasoning above and Eq.
+(20) hold irrespective of the precise dynamics in the bulk $A$.
+Notice that in principle one could apply the $GHD$ directly to the tripartite protocol, following, for instance, Ref.
+[34].
+The reasoning above allows us to conclude that for tmuch-greater-$t\gg\ell$ the bound states with $n=2$ can leave region $A$ after the “evaporation” of the unbound particles.
+As for the second scenario in which there is no bound-state confinement, we observe that the condition $\vartheta_{1,L}={\mathcal{O}}(1)$ is not valid for generic initial states of $A$.
+For instance, it is violated by all the $p$-Néel states with generic $p>1$.
+By using the framework of section 3, one can check that $\vartheta_{n,L}={\mathcal{O}}(z^{2})$ for $n<p$, whereas $\vartheta_{n,L}={\mathcal{O}}(1)$ for $n\geq p$.
+In turn, by solving the $TBA$ equations (3) for the particle densities $\rho_{n}(\lambda)$, this implies that $\rho_{p}={\mathcal{O}(1)}$, whereas $\rho_{n\neq p}$ are suppressed in the large $\Delta$ limit.
+This reflects that since the $p$-Néel state contains $p$ down spins on consecutive sites it has overlap mostly with eigenstates of the $XXZ$ chain that contain bound states of $p$ magnons.
+This has striking consequences on transport properties and bound-state confinement.
+Indeed, we anticipate that the fact that $\vartheta_{n,L}={\mathcal{O}}(z^{2})$ implies a strong suppression of bound-state transport for $p\geq 4$.
+Let us first consider the case with $p=2$.
+Now, Eq.
+(20) suggests that bound states with $n=2$ are not confined.
+Indeed, the term with $m=1$ in the right-hand-side of (20) is ${\mathcal{O}}(z^{2})$ and it is negligible compared to $\partial_{\lambda}\varepsilon_{2}^{\mathrm{b}}={\mathcal{O}}(z)$.
+This means that unlike the case with $p=1$ the two-particle bound states cannot be confined by the magnons, whereas all the bound states with $n>2$ are confined.
+We observe a similar behavior for $p=3$ (see section 5.1), i.e., that the bound states with $n\leq p$ are not confined, whereas the ones with $n>p$ are all confined.
+Now, a different scenario emerges for $p\geq 4$.
+Indeed, for $p=4$ we observe that since $\partial_{\lambda}\varepsilon_{4}^{\mathrm{b}}={\mathcal{O}}(z^{3})$, it is clear that the second term in (20) dominates at large $\Delta$ because of the scatterings with the quasiparticles with $n\leq 4$, which contribute with terms with smaller powers of $z$.
+This means that for $p=4$ all the bound states with $n<p$ are not confined, whereas the ones with $n\geq p$ are confined.
+This implies that upon increasing $p$ the bound states that dominate the thermodynamic properties, i.e., the bound states with $n=p$, become confined, and transport is suppressed.
+5.1 Bound-state confinement: Numerical GHD results
+Figure 2: Group velocities $v_{n}$ of the $n$-particle bound states calculated over the $GGE$ with $\zeta:=x/t=0$, which in the hydrodynamic limit describes the interface between region $A$ and the rest (Fig.
+1).
+At $t=0$ region $A$ is prepared in the $p$-Néel states (cf.
+(2)).
+The different panels correspond to different values of $p$ and $\Delta$.
+Panel (a) and (b) are for $p=1$.
+Panels (c-f) show results for $p=2-5$.
+In all the panels one has that $v_{1}(\lambda)$ is positive in an extended region in $\lambda$, signaling that magnonic excitations are never confined.
+(a) The case of $p=1$, i.e., the Néel state.
+At $\Delta\lesssim 1.4$ $v_{2}<0$ for any $\lambda$, i.e., the two-particle bound states are confined.
+(b) Same as in (a) for $\Delta=1.01$.
+While bound states with $n=2$ are not confined, $v_{3}\leq 0$ for any $\lambda$, which means that three-particle bound states are confined.
+(c) The $p$-Néel state with $p=2$.
+Bound states with $n\geq 3$ are confined.
+At $\Delta\approx 1.09$ three-particle bound states are not confined anymore.
+(d) The case with $p=3$.
+At $\Delta\approx 1.02$ the quasiparticles with $n=4$ are not confined.
+(e) The case with $p=4$.
+Now, even quasiparticles with $n=4$ remain confined, at least up to $\Delta\approx 1.07$.
+(f) The Néel state with $p=5$.
+Similar to (e), quasiparticles with $n=p$ are confined.
+Although qualitative features of bound-state confinement are understood perturbatively, to characterize quantitatively bound-state confinement it is necessary to determine numerically the dressed velocities $v_{n}(\lambda)$.
+Here we discuss numerical results confirming the scenario outlined in section 5.
+Precisely, we numerically solve the $GHD$ equation (13) to obtain the dressed velocities $v_{n}$ for $\zeta=0$.
+The numerical solution of (13) is standard.
+By using a Gauss discretization, Eq.
+(13) becomes an infinite system of nonlinear equations, which can be truncated considering only bound states of size $n\leq n_{\mathrm{max}}$.
+Finally, the truncated system can be solved with standard methods.
+By studying how the velocities change by varying $n_{\mathrm{max}}$, we observe that $n_{\mathrm{max}}=40$ is sufficient to have converged results.
+Our results are reported in Fig.
+2.
+We plot $v_{n}(\lambda)$ for $1\leq n\leq 5$ versus $\lambda$.
+We consider as initial states the $p$-Néel states (cf.
+(2)) with $1\leq p\leq 5$.
+For $p=1$ the quench is the popular Néel quench.
+At large $\Delta$ one has that $v_{n}<0$ for any $\lambda$ and for any $n>1$, meaning that all the bound states are confined within region $A$.
+Fig.
+2 (a) shows that for $\Delta\approx 1.4$ the bound states with less negative velocity $n=2$ start to be transmitted [17].
+Upon lowering $\Delta$, at $\Delta\approx 1.01$, the bound states with $n=3$ are not confined, as shown in Fig.
+2 (b).
+For $p=2$, as anticipated in section 5, the bound states with $n=2$ are not confined.
+Figure 3: Group velocities $v_{n}$ of the $n$-particle bound states calculated over the $GGE$ with $\zeta:=x/t=0$.
+At $t=0$ region $A$ is prepared in the $p$-Néel state (cf.
+(2)) with $p=2$.
+The remaining part of the chain is prepared in the ferromagnet.
+We show results for the $XXZ$ chain with $\Delta=5$.
+Notice that the two-particle bound states are not confined because $v_{2}(\lambda)>0$ in the region with $\lambda\approx 0.5$.
+On the other hand $v_{3}(\lambda)$ and $v_{4}(\lambda)$ are negative for all values of $\lambda$, signaling confinement.
+Moreover, both $v_{3}(\lambda)$ and $v_{4}(\lambda)$ exhibit a “flat” behavior as functions of $\lambda$, as expected in the large $\Delta$ limit.
+On the other hand, Fig.
+2 (c) shows that for $p=2$ at $\Delta\approx 1.1$ the three-particle bound states are still confined in $A$.
+For $p=2$, we checked that absence of confinement for the two-particle bound states persists also at large $\Delta$.
+In Fig.
+3 we report numerical results for the group velocities $v_{n}$ up to $n=4$ and the quench from the $p$-Néel state with $p=2$.
+The results are for $\Delta=5$.
+The fact that $v_{2}(\lambda)$ is positive in the region with $\lambda\approx 0.5$ signals that the bound states with $n=2$ are not confined.
+Moreover, we observe that larger bound states with $n=3,4$ are confined.
+The group velocities $v_{3}(\lambda)$ and $v_{4}(\lambda)$ are “flat”, i.e., they exhibit a weak dependence on $\lambda$, in agreement with the discussion in section 5.
+A similar scenario occurs for $p=3$, i.e., the bound states with $n<4$ are not confined, whereas the bound states with $n\geq 4$ are all confined up to $\Delta\approx 1.02$.
+However, the scenario changes at $p=4$.
+As it is clear from Fig.
+2 (e), for $p\geq 4$ the group velocities $v_{p}(\lambda)$ are negative for any $\lambda$.
+This holds true for $p=5$ as well (see Fig.
+2 (f)).
+Since the thermodynamic properties of the system are dominated by the bound states with $n=p$, this means that transport becomes more and more suppressed upon increasing $p$.
+6 Numerical tDMRG results
+Here we discuss $tDMRG$ data [24] for the expansion protocols introduced in Fig.
+1.
+Our results support the bound-state confinement scenario outlined in section 5.
+As a preliminary check, in section 6.1 we discuss $tDMRG$ data for the full-expansion protocol $(a)$ in Fig.
+1, considering times tmuch-less-$t\ll\ell$, which allows us to neglect the effects of the finite size of $A$ and use the $GHD$ results for the bipartite protocol.
+In our $tDMRG$ simulations we employ a Trotter decomposition of the evolution operator.
+During the dynamics the bond dimension $\chi$ of the Matrix Product State ($MPS$) representation of the wavefunction grows.
+At each time step we compress the $MPS$ by performing a Singular Value Decomposition, keeping the largest $\chi$ singular values.
+We increase the bond dimension $\chi$ until convergence of physical observables (see, for instance, B).
+Typically, in our simulations we use $\chi\approx 10^{3}$.
+Here we benchmark the results of section 4 for several $p$-Néel initial states.
+In section 6.2 we focus on the full-expansion protocol (see Fig.
+1 (a)) considering times tmuch-greater-$t\gg\ell$.
+In section 6.3 we discuss the cut quench protocol (see Fig.
+1 (b)).
+Figure 4: Releasing the $p$-Néel states in the vacuum (see Fig.
+1).
+The figure shows $tDMRG$ data for the local magnetization $\langle\sigma_{x}^{z}\rangle$ plotted as a function of $\zeta=x/t$.
+Here $x$ is distance from the interface between region $A$ and the rest (Fig.
+1).
+The dashed line is the $GHD$ result valid in the limit $x,t\to\infty$ at fixed $\zeta$.
+We show results for $p\leq 3$ and $\Delta=1,2$.
+The different symbols denote different times $t$.
+Notice that for $p=3$ and $\Delta=2$ $t,x$ are not large enough to reach the hydrodynamic limit.
+Figure 5: Releasing the Néel states in the vacuum (see Fig.
+1).
+We plot the expectation value $\langle P_{x,\uparrow\uparrow}\rangle$ (cf.
+(25)) versus the rescaled time $\zeta=x/t$, with $x$ the distance from the interface between region $A$ and the rest (see Fig.
+1).
+The size of region $A$ is $\ell=40$ sites.
+Panels (a) and (b) show the expansion dynamics for $\Delta=1$ and $\Delta=2$, respectively.
+The different symbols denote different times.
+We only show results up to $t\lesssim 15$, to ensure that in the bulk of region $A$ one has $\vartheta_{1,L}={\mathcal{O}}(1)$, which implies bound-state confinement (see section 5).
+Notice that at $\zeta=0$ for $\Delta=1$ one has a finite value for $\langle P_{x,\uparrow\uparrow}\rangle$, whereas at $\Delta=2$ one has $\langle P_{x,\uparrow\uparrow}\rangle\approx 0$, reflecting bound-state confinement.
+6.1 Validity of GHD
+Before investigating bound-state confinement, it is crucial to assess the validity of the $GHD$ description for the protocols discussed in Fig.
+1.
+This is summarized in Fig.
+4.
+We consider the expansion of the $p$-Néel states with $1\leq p\leq 3$.
+The different panels show $tDMRG$ data for the local magnetization $\sigma_{x}^{z}$ as a function of $\zeta=x/t$.
+For each $p$ we show results for $\Delta=1,2$.
+Here $x$ is the distance from the right edge of region $A$ (see Fig.
+1).
+In our $tDMRG$ simulations we consider times tmuch-less-$t\ll\ell$ to avoid effects due to the finite size of $A$.
+In Fig.
+4 we provide $tDMRG$ data for several times up to $t\approx 20$ (different symbols in the Figure).
+The $tDMRG$ data exhibit collapse when plotted as function of $x/t$, confirming that $t\approx 20$ is large enough to ensure the hydrodynamic limit for all the values of $p$ and $\Delta$ considered, except for $p=3$ with $\Delta=2$ (see Fig.
+4 (f)).
+For $p=3$ and $\Delta=2$ the evolution in the bulk of region $A$ is slower compared to the other cases.
+Indeed, at $\zeta\leq-1$, the magnetization profile exhibits strong oscillations, which are reminiscent of the structure of the initial state.
+In all the panels, the dashed curves are the $GHD$ results obtained by solving numerically the $GHD$ equation (13).
+The numerical results were obtained by truncating the infinite system of equations (13) keeping only the bound states with $n\leq n_{\mathrm{max}}=40$.
+The $GHD$ equation (13) allows to obtain the occupation functions $\vartheta_{\zeta=0,n}(\lambda)$.
+The numerical results for $\vartheta_{0,n}(\lambda)$ can be substituted in the $TBA$ equations (3), after using that $\rho_{t,n}(\lambda)=\vartheta_{n}(\lambda)\rho_{n}(\lambda)$.
+Eq.
+(3) allows one to obtain the quasiparticle densities $\rho_{n}(\lambda)$, which encode information about local observables.
+For instance, the local magnetization is given as
+[…]
+For all the quenches that we explore in panels (a-e) of Fig.
+4 the $GHD$ results are in satisfactory agreement with the $tDMRG$ data, despite the limitation to times $t\lesssim 20$.
+For the expansion of the $p$-Néel state with $p=3$ and $\Delta=2$ (panel (f) in Fig.
+4) deviations from the $GHD$ prediction are large.
+Also, notice that for $\zeta>0$ one has that $\langle\sigma_{x}^{z}\rangle\approx-1$, signaling that transport in general is strongly suppressed for $p=3$.
+As anticipated in section 5, for the $p$-Néel state with $p=1$, the bound states with $n=2$ remain confined in region $A$, as long as the system is in the hydrodynamic limit, and the $GHD$ description is valid.
+Specifically, this holds for $x,t\to\infty$ with $x/t$ fixed, with $x$ being the distance from the right edge of region $A$ (see Fig.
+1).
+Moreover, for the bound-state confinement to happen, the condition $\vartheta_{1,L}={\mathcal{O}}(1)$ has to be satisfied.
+This is the case for $p=1$, at least if tmuch-less-$t\ll\ell$, i.e., before the magnons leave region $A$.
+Confinement of the two-particle bound states is discussed in Fig.
+5.
+The Figure shows $P_{x,\uparrow\uparrow}$ (cf.
+(25)) plotted versus $\zeta=x/t$.
+The projectors $P_{x,\uparrow^{\otimes m}}$ are defined as
+$P_{x,\uparrow^{\otimes m}}:=\frac{1}{2^{m}}\prod_{j=0}^{m-1}(\mathds{1}+\sigma_{x+j}^{z}).$ (25)
+Although the projectors (25) are not bona fide witnesses of bound states, we anticipate that $P_{x,\uparrow^{\otimes m}}$ are sensitive to the presence of a $m$-particle bound state at position $x$ [10, 27].
+Panels (a) and (b) show $tDRMG$ data for $P_{x,\uparrow\uparrow}$ for the $XXZ$ Hamiltonian with $\Delta=1$ and $\Delta=2$, respectively.
+In each panel the different symbols correspond to different times.
+For both $\Delta=1$ and $\Delta=2$ the $tDMRG$ data exhibit scaling as a function of $\zeta$.
+Crucially, while for $\Delta=1$ (Fig.
+5 (a)) $\langle P_{x,\uparrow\uparrow}\rangle$ is nonzero for $\zeta>0$, one has $\langle P_{x,\uparrow\uparrow}\rangle\approx 0$ for any $\zeta>0$ for $\Delta=2$ (see Fig.
+5).
+This reflects that while for $\Delta=1$ the bound states with $n=2$ are not confined, for $\Delta=2$ they remain confined in region $A$ during the expansion.
+Still, at $\zeta=0$, $\langle P_{x,\uparrow\uparrow}\rangle\approx 0.025$ for $\Delta=1$, which suggests that the density of two-particle bound states that are not confined is “small”.
+This means that it could be challenging to verify confinement, for instance experimentally.
+Figure 6: Releasing the Néel state in the vacuum.
+Region $A$ in Fig.
+1 of size $\ell=20$ is prepared in the Néel state at $t=0$.
+The remaining part of the chain is in the vacuum state, i.e., the ferromagnetic state.
+At $t>0$ the chain is let to evolve under the $XXZ$ chain Hamiltonian respectively with $\Delta=1$ in the top row and $\Delta=2$ in bottom one.
+Panels (a,d) show $tDMRG$ data for the local magnetization $\langle\sigma^{z}_{x}\rangle$ at generic position $x$ and at different times.
+Notice the reflection at the boundary of the chain at long times.
+Panels (b,e) and (c,f) show the projectors $P_{x,\uparrow\uparrow}$ and $P_{x,\uparrow\uparrow\uparrow}$.
+6.2 Full expansion dynamics
+In section 6.1 we established the validity of the $GHD$ description for the quenches from the $p$-Néel states.
+Moreover, Fig.
+4 suggests that $t\approx 20$ is enough to access the hydrodynamic limit, which is essential to observe bound-state confinement.
+We now focus on the full expansion dynamics, i.e., considering times tmuch-greater-$t\gg\ell$.
+We show $tDMRG$ data for the expansion of the $p$-Néel states with $p=1$ for $\Delta=1$ and $\Delta=2$ in Fig.
+6.
+The protocol is reported in Fig.
+1.
+Region $A$ at center of the chain is prepared in the Néel state.
+Here we consider $\ell=20$.
+The reason for choosing $\ell=20$ is that it allows to simulate the dynamics using $tDMRG$ up to times $t\approx 100$.
+Indeed, entanglement, which is the major limitation of $tDMRG$, is produced in the bulk of region $A$ only, whereas the expansion does not contribute significantly to the entanglement growth.
+This also reflects that the $GHD$ dynamics preserves the thermodynamic entropy [16, 18].
+As a consequence, the largest entanglement produced during the dynamics is bounded by $\ell$.
+The different panels in Fig.
+6 show different observables.
+Precisely, we consider the local magnetization $\sigma_{x}^{z}$, and the local projectors $P_{x,\uparrow\uparrow},P_{x,\uparrow\uparrow\uparrow}$ (cf.
+(25)).
+As anticipated, and has already observed in Ref.
+[27] and Ref.
+[10], $P_{x,\uparrow^{\otimes m}}$ is sensitive to bound states of $m$ spins starting from position $x$.
+Let us first discuss the dynamics with the $XXZ$ chain at $\Delta=1$.
+In Fig.
+6 (a,b,c) we plot the density of local magnetization $\langle\sigma_{x}^{z}\rangle$ as a function of position $x$ and time $t$.
+After the expansion starts, the central region “melts” as quasiparticles are emitted at its edges.
+First, several “jets” of excitations are visible, traveling at different velocities.
+The fastest quasiparticles correspond to the magnonic excitations, i.e., with $n=1$, whereas the jets moving at smaller velocities correspond to bound states with $n=2$ and $n=3$ magnons.
+The dashed lines are the bare maximum group velocities of the quasiparticles (cf.
+(8)).
+Since the initial state is expanded in the vacuum, the density of quasiparticles at the edges of the jets is small and interactions are “weak”, which explains why the bare velocities describe the dynamics of the jets.
+In Fig.
+6 (b) we show the dynamics of the projector $P_{x,\uparrow\uparrow}$.
+Now, the outermost jet in Fig.
+6 (a) is not present, confirming that it is due to the quasiparticles with $n=1$.
+Oppositely, the two innermost jets of Fig.
+6 (a), corresponding to $n=2,3$ are still present.
+Similar behavior was observed in the expansion of few-spin initial states in Ref.
+[27].
+Importantly, one should observe that at short times $t\lesssim 20$, $P_{x,\uparrow\uparrow}$ exhibits a “halo” feature at the edges of region $A$.
+This can be attributed to the fact that for $\Delta=1$ the two-particle bound states are not confined in the hydrodynamic limit.
+Notice, however, that the halo is quite weak, reflecting that the density of two-particle bound states at the interface between $A$ and the rest is small.
+A similar scenario is observed for $P_{x,\uparrow\uparrow\uparrow}$ (see Fig.
+6 (c)).
+Now, the three-particle bound states behave as if they were confined up to $t\approx 20$.
+Notice that the fact (see Fig.
+2 (b)) that $v_{3}\leq 0$ at $\Delta\approx 1.01$ suggests that the density $\rho_{3}(\lambda)$ of three-particle bound states at the interface between $A$ and the vacuum could be small.
+This implies that even though bound states with $n=3$ are not confined at $\Delta=1$, this could be difficult to observe.
+Let us now discuss the same protocol in the $XXZ$ chain with $\Delta=2$ (see Fig.
+6, bottom row).
+The qualitative behavior is the same as in the top row in Fig.
+6 .
+However, the projector $P_{x,\uparrow\uparrow}$ at short times $t\lesssim 20$ exhibits a “bottleneck” feature.
+In contrast with $\Delta=1$, at $\Delta=2$ the two-particle bound states are confined (see Fig.
+2 (f)).
+Moreover, in the hydrodynamic regime the density $\rho_{2}(\lambda)$ is zero for $\zeta=0$, and it is presumably small also at $\zeta\lesssim 0$.
+This explains the bottleneck in Fig.
+6 (d,e,f).
+At long times after all the magnons have left region $A$, the two-particle bound states are allowed to leave region $A$, as discussed in section 5, and as it is confirmed in Fig.
+6 (e).
+A similar dynamics happens for the three-particle bound states, as shown in Fig.
+6 (f).
+Figure 7: Releasing the Néel state in the vacuum.
+Same setup as in Fig.
+6.
+Panels (a) and (b) show respectively the dynamics with $\Delta=2$ and $\Delta=1$.
+We show the magnetization profile $\langle\sigma_{x}^{z}\rangle$ as a function of the position $x$ in the chain.
+The different curves correspond to different times.
+More precisely, the blue (orange; green) curve shows the magnetization profile at time $t=20$ ($t=40$; $t=80$).
+Notice the different peaks, which correspond to different families of excitations.
+Precisely, the outermost peaks correspond to magnonic excitations, whereas the innermost ones to the bound states with $n\geq 2$.
+Let us now discuss the space-time profiles of local observables.
+In Fig.
+7 (a) and (b) we show the profile of the local magnetization $\langle\sigma_{x}^{z}\rangle$ after the expansion of the Néel state for the $XXZ$ chain with $\Delta=1$ and $\Delta=2$, respectively.
+Both panels in Fig.
+7 confirm that for $\Delta=1$ and $\Delta=2$ the bound states with $n=2$ and $n=3$ are not confined at long times, as already observed in Fig.
+6.
+The $tDMRG$ data for $\Delta=1$ (left panel in Fig.
+7) show a rapid melting of the profile.
+As the dynamics proceeds, $\langle\sigma_{x}^{z}\rangle$ exhibits a multi-peak structure, each peak corresponding to a different bound-state family.
+For $\Delta=1$ the peaks are blended together, reflecting that the bound states of different types start to leave region $A$ at the same time.
+In particular, already at $t=40$ the bound states with $n=2$ are moving towards the edges of region $A$.
+In the inset we show the dynamics of $P_{x,\uparrow\uparrow}$, which is sensitive to quasiparticles with $n\geq 2$ only.
+The inset shows a rapid melting of $P_{x,\uparrow\uparrow}$.
+In contrast, for $\Delta=2$ (right panel in Fig.
+7) the dynamics is slower, and the peaks corresponding to different $n$ are well separated.
+Notice that at $t=80$ the lump around $x=100$ corresponds to the three-particle bound states, which remain almost locked at the center of the chain.
+This observation suggests that for $\Delta=2$ the trap-expansion protocol can be used to “distil” states with large bound-state content.
+The inset shows the dynamics of $P_{x,\uparrow\uparrow}$.
+Notice that for $t\lesssim 10$ the bound states with $n=2$ are still within region $A$.
+Importantly, for $\ell=20$ the density of quasiparticles with $n=1$ in the bulk of $A$ starts to be small already at $t\approx 10$, implying that bound-state confinement is not supported anymore (see section 5).
+This behavior does not violate GHD analytical prediction.
+In fact, for times $t\gtrsim 20$, the bipartite GHD formulation is not anymore predictive for the tripartite protocol due to the finite length of the region $A$.
+Figure 8: Releasing the Néel state in the vacuum.
+At $t=0$ the central part $A$ (see Fig.
+1) of a spin chain containing $\ell=20$ sites is prepared in the Néel state.
+At $t>0$ the chain is let to evolve under the $XXZ$ chain Hamiltonian with different values of $\Delta=\{1,1.2,1.4,2\}$.
+For $0<t\leq 20$ the full chain evolves.
+At $t>20$ only the $B$ part (see Fig.
+1) is evolved.
+Panels (a)-(d) show $tDMRG$ data for the local magnetization $\langle\sigma_{x}^{z}\rangle$ as a function of the position $x$ in the chain.
+Panels (e)-(f) show results for the projector $P_{x,\uparrow\uparrow}$.
+Notice that at $\Delta=2$, $P_{x,\uparrow\uparrow}$ vanishes, signalling bound-state confinement.
+The sudden jumps around $t\approx 20$ are numerical artifacts due to the fact that we truncate the $MPS$ representation of the system for $t>t^{*}$.
+6.3 Dynamics after the cut protocol
+To investigate bound-state confinement, it is also convenient to employ the protocol in Fig.
+1 (b).
+The full chain is now evolved with the $XXZ$ Hamiltonian up to a time $t^{*}$.
+Here $t^{*}$ is chosen large enough to ensure the validity of the hydrodynamic limit.
+As it is clear from section 6.1, $t\approx 20$ is sufficient to ensure the hydrodynamic limit.
+Moreover, $t^{*}$ is chosen short enough compared to $\ell$ to have $\vartheta_{1,L}={\mathcal{O}}(1)$ in the bulk of $A$, which is essential to have bound-state confinement.
+Here we choose $t^{*}=20$ and $\ell=40$.
+We consider the $p$-Néel states with $p=1-3$.
+The cut protocol allows to accurately study the nature of the excitation that at time $t>t^{*}$ are in the region on the right of $A$.
+In fact, all the quasiparticles that are confined in region $A$ for $0<t<t^{*}$, are ”frozen” by the cut protocol and do not contribute to the dynamics.
+6.3.1 $p$-Néel state with $p=1$
+Our results for the Néel state (i.e., for $p=1$) are shown in Fig.
+8.
+Panels $(a-d)$ show the local magnetization $\sigma_{x}^{z}$, as a function of position $x$ and time $t$.
+In the Figure we show $tDMRG$ data.
+The $tDMRG$ algorithm is the same as in section 6.1 and 6.2.
+In our simulations we keep the bond dimension $\chi\approx 1000$ for $t\leq t^{*}$.
+We observe that typically for $x>0$, i.e., in region $B$, the bond dimension needed to accurately describe local observables is significantly smaller as compared with $x<0$ (see B).
+This is expected since entanglement is produced in region $A$ only.
+Thus, since at $t>t^{*}$ we evolve only region $B$ we further reduce the bond dimension to $\chi\approx 800$.
+As it is clear from the Figure, for $t>t^{*}$ (see Fig.
+1), under the evolution with the Hamiltonian (1) restricted to $x>0$ the quasiparticles emitted from $A$ propagate ballistically in the vacuum.
+The dependence on $\Delta$ is weak.
+In the bottom row in Fig.
+8 (panels (e-f) in the Figure) we report the dynamics of $P_{x,\uparrow\uparrow}$.
+For $\Delta=1$, $\langle P_{x,\uparrow\uparrow}\rangle$ exhibits a ballistic spreading.
+The nonzero signal reflects that the bound states with $n=2$ are not confined.
+$P_{x,\uparrow\uparrow}$ becomes smaller upon increasing $\Delta$.
+For instance, it is barely visible at $\Delta=2$.
+Notice that in the hydrodynamic limit one should expect that $\langle P_{x,\uparrow\uparrow}\rangle$ is exactly zero for $\Delta>\Delta^{*}\approx 1.4$ (see Fig.
+2).
+6.3.2 $p$-Néel state with $p=2$
+Let us now discuss the cut protocol (see Fig.
+1 (b)) for the $p$-Néel state with $p=2$.
+Our results are reported in Fig.
+9.
+Panels (a)-(c) show $tDMRG$ data for the dynamics with the $XXZ$ chain Hamiltonian and $\Delta=1,2,4$.
+In the Figure we plot the local magnetization $\langle\sigma_{x}^{z}\rangle$ as a function of position $x$ and time $t$.
+For $\Delta=1$ (Fig.
+9 (a)) two distinct signals are visible.
+The two “jets” correspond to quasiparticles with $n=1$, i.e., free ones, and to the two-particle bound states.
+The two-particle jet is due to the fact that for $p=2$ the bound-state content of the chain is enhanced as compared to $p=1$, and two-particle bound states are not confined at $\Delta=1$.
+Notice also that the bound states exhibit smaller group velocities as compared with the quasiparticles with $n=1$.
+The dashed lines reported in Fig.
+9 (a) are the maximum bare velocities $v_{n,\mathrm{max}}^{\mathrm{b}}=\max_{\lambda}(v_{n}^{\mathrm{b}}(\lambda))$ (cf.
+(8)).
+In Fig.
+9 (b) we show data for $\Delta=2$.
+Crucially, the jet corresponding to the two-particle bound states is still visible.
+This is in agreement with the results of section 5.
+Indeed, for $p=2$ one should expect that the quasiparticles with $n=2$ are never confined.
+On the other hand, as it is clear from Fig.
+2 (c), the three-particle bound states are confined for $\Delta\gtrsim 1.1$.
+We should mention that, however, as the density of quasiparticles with $n=3$ is “small”, it is numerically challenging to observe the three-particle bound states confinement.
+Finally, in Fig.
+9 (c) we show results for $\Delta=4$.
+Now, the jet corresponding to $n=1$ is weak, in contrast with the case with $p=1$.
+This is due to the fact that for $p=2$ the density $\rho_{1}(\lambda)$ even in the bulk of $A$ vanishes in the limit of $\Delta$ large, reflecting that thermodynamic properties of the system are dominated by the bound states with $n=2$.
+On the other hand, although one has $\rho_{2}(\lambda)=\mathcal{O}(1)$ in the bulk of $A$, the flux of two-particle bound states through the interface between $A$ and the rest is suppressed upon increasing $\Delta$ because the bound-state velocities, despite the dressing, decrease with $\Delta$.
+This explains why the jet associated with the two-particle bound states is not visible in Fig.
+9 (c).
+Notice that this is a “trivial” suppression of transport, which depends in a continuous way on $\Delta$.
+In particular, the $1/\Delta$ dependence of $v_{2}^{b}$ would suggest a $1/\Delta$ suppression of two-particle bound states upon increasing $\Delta$.
+Finally, in Fig.
+9 (d,e) we report the projector $P_{x,\uparrow\uparrow\uparrow}$ after the expansion of the $p$-Néel state with $p=2$.
+In Fig.
+9 (d) and (e) we show results for $\Delta=1$ and $\Delta=2$, respectively.
+While for $\Delta=1$ ballistic spreading is visible, for $\Delta=2$, the projector $P_{x,\uparrow\uparrow\uparrow}$ is “frozen”.
+Figure 9: Releasing the “fattened” $p$-Néel state with $p=2$ in the vacuum in the cut protocol (see Fig.
+1 (b)).
+The figure shows $tDMRG$ data for the local magnetization $\langle\sigma_{x}^{z}\rangle$ as a function of $x$ and the time $t$.
+The different panels (a-c) show results for different values of $\Delta$.
+For $\Delta=1$ and $\Delta=2$ the two jets correspond to the magnonic excitations and the two-particle bound states.
+Panels (d,e) show $\langle P_{x,\uparrow\uparrow\uparrow}\rangle$ as a function of time and position $x$ for $\Delta=1$ and $\Delta=2$.
+Notice that for $\Delta=2$ the quasiparticles with $n=3$ are confined in $A$.
+The origin of the sudden jumps visible at $t\approx 20$ in (a-c) is the same as explained in the caption of Fig.
+8.
+Figure 10: Releasing the Néel state in the vacuum.
+Dynamics of the entanglement entropy.
+The curve shows $tDMRG$ data for a chain with $L=200$ sites.
+The size of region $A$ (see Fig.
+1) is $\ell=20$.
+The plot shows the entropy $S$ of subsystem $B$ (see Fig.
+1).
+Since subsystem $A$ includes the Néel part, the growth of $S$ with time is due to the entanglement between the excitations that leave the Néel part and those that are left behind.
+At long times $S$ measures the entanglement between the “jets” of excitations in Fig.
+6.
+Notice the multistage growth of the entropy.
+At short times the growth is dominated by the entangled magnonic excitations.
+The second slope visible in the figure is attributed to the two-excitations bound states.
+The last stage of the growth is due to the three-particle bound states.
+7 Bound-state confinement and dynamics of the entanglement entropy
+Here we show that bound-state confinement is sharply reflected in the dynamics of the von Neumann entropy.
+We focus on the full expansion of the Néel state (see Fig.
+1 (a)).
+Fig.
+10 shows the von Neumann entropy $S$ of region $B$ (see Fig.
+1 (a)).
+The dotted curve is the result for $\Delta=1$, whereas the full line shows the result for $\Delta=2$.
+First, for both values of $\Delta$ the von Neumann entropy saturates at long time.
+This reflects that the entanglement dynamics is due to the creation of pairs of entangled quasiparticles in region $A$ [12].
+This is the content of the so-called quasiparticle picture [35, 29, 12, 36].
+The two quasiparticles forming the entangled pairs travel with opposite group velocities.
+As they propagate, they entangle larger regions.
+Precisely, the von Neumann entropy of a region is proportional to the number of entangled pairs that are shared between the region and the rest.
+The contribution of a shared pair characterized by rapidities $(-\lambda,\lambda)$ is the contribution of that pair to the thermodynamic entropy of the $GGE$ describing the steady state after the quench from the homogeneous Néel state.
+The quasiparticle picture holds in the hydrodynamic limit $\ell,t\to\infty$ with the ratio $t/\ell$ fixed.
+Notice that the saturation of the von Neumann entropy at long times is due to the finite-size of $A$.
+Indeed, since $A$ is finite, the number of entangled pairs that can generate entanglement between $A$ and $B$ is bounded by $\ell$.
+For a global quench from two semi-infinite chains prepared in nontrivial initial states, the entanglement entropy of each chain would grow indefinitely.
+Let us focus on the results for $\Delta=1$ in Fig.
+10.
+The initial linear growth up to $t\approx 20$ is due to the entangled pairs that are initially produced in $A$ and at later times leave region $A$, and hence start to be shared with the rest (see Fig.
+1).
+Again, the saturation of $S$ at long times reflects that region $A$ is finite.
+For $\Delta=1$, $S$ is essentially constant for $t\gtrsim 60$.
+The saturation of the entanglement entropy is the key to reach times long as $t\approx 120$ with $tDMRG$, because after the entropy saturates the computational cost is mild.
+One should also observe that the saturation of $S$ happens “smoothly”.
+This is somewhat reminiscent of the fact that for $\Delta=1$ the bound states are not confined.
+Crucially, for $\Delta=2$, $S$ exhibits a multistage linear growth before saturating.
+This reflects the bound-state confinement.
+Specifically, the linear growth up to $t\approx 20$ is due to the quasiparticles with $n=1$.
+In this early time regime, the slope of the linear growth should be the same as in the vacuum expansion of a semi-infinite chain prepared in the Néel state.
+The prediction of the quasiparticle picture is [16, 18]
+[…]
+where $s_{1,\zeta=0}^{YY}$ is the Yang-Yang entropy [2] of the macrostate with $\zeta=0$ that describes the interface between region $A$ and region $B$ (see Fig.
+1).
+$s_{n,\zeta}^{YY}$ is readily obtained from the particle densities $\rho_{n,\zeta}(\lambda)$ describing the $GGE$ as
+$s_{n,\zeta}(\lambda):=\rho_{t,n,\zeta}\ln(\rho_{t,n,\zeta})-\rho_{n,\zeta}\ln(\rho_{n,\zeta})-\rho_{h,n,\zeta}\ln(\rho_{h,n,\zeta}),$ (27)
+where the densities $\rho_{n,\zeta},\rho_{h,n,\zeta},\rho_{t,n,\zeta}$ are defined in section 3 (see also section 4).
+Notice that in (26) only the quasiparticles with $n=1$ contribute, reflecting that all the bound states are confined in $A$, and cannot contribute to the von Neumann entropy.
+Now, the $tDMRG$ data in Fig.
+10 exhibit deviations from (26).
+This is expected because (26) holds in the hydrodynamic limit $t\to\infty$.
+Finally, we should mention that it is possible to characterize the trajectories [18] of the generic entangled pairs produced in $A$, which are responsible for the entanglement growth.
+This, in principle, could allow to determine also the secondary slopes visible in Fig.
+10.
+8 Conclusions
+We investigated bound-state confinement after trap-expansion experiments in the $XXZ$ spin chain at $\Delta>1$.
+We showed that after the expansion in the vacuum of quite generic initial states the bound states of the $XXZ$ chain remain confined in the initial region, provided that $\Delta$ is large enough.
+Upon lowering $\Delta$ the bound states are allowed to leave region $A$, i.e., they are not confined anymore.
+The condition ensuring confinement is the presence of a finite density of magnonic excitations in the bulk of the initial region.
+Physically, bound-state confinement is a genuine consequence of interactions, i.e., the scattering, between the bound states and the magnons.
+Interestingly, if the size of the initial region is finite, at long times the density of magnons vanishes, as they leave the initial region, and bound-state confinement breaks down.
+Thus, for the expansion of finite regions, bound-state confinement is present only at intermediate times $t={\mathcal{O}}(\ell)$.
+We investigated bound-state confinement after the expansion of the $p$-Néel states (see Fig.
+1).
+In the limit $p\to\infty$ one recovers the domain wall initial state, and transport is suppressed.
+We showed that for small $p\lesssim 4$, bound-state confinement happens for the bound states of size $n>p$.
+At larger $p$, even the bound states with $n=p$ are confined, which is consistent with the absence of transport in the limit $p\to\infty$.
+Finally, we showed that bound-state confinement is reflected in a multi-stage linear growth of the entanglement entropy.
+Let us now discuss some directions for future research.
+First, we investigated bound-state confinement in a lattice model.
+It would be interesting to understand whether confinement occurs for systems in the continuum, such as the attractive Lieb-Liniger gas.
+This would pave the way for the experimental verification of bound-state confinement, for instance with atom-chip experiments [37, 38].
+It would be also interesting to investigate confinement in lattice models that can be realized with trapped ions [8] and cold-atom experiments [39].
+Interestingly, bound-state confinement could be potentially used to prepare thermodynamic states with high bound-state content.
+Specifically, the protocol to enrich the bound state content would consists of two steps.
+In a first step one can expand the initial state as in Fig.
+1 (a).
+Then, after the magnonic excitation are “evaporated” one could trap the remaining excitations by applying an external potential.
+As the bound states are typically associated with lower thermodynamic entropy, bound-state confinement can also be used to prepare low-entropy states.
+It has been pointed out recently that bound-states could be robust against long-range interactions [40, 8].
+It would be interesting to understand whether long-range spin chains exhibit bound-state confinement as well.
+Finally, it would be important to clarify whether bound-state confinement is also present in integrable quantum circuits [41, 42, 43].
+Acknowledgements
+This study was carried out within the National Centre on HPC, Big Data and Quantum Computing - SPOKE 10 (Quantum Computing) and received funding from the European Union Next-GenerationEU - National Recovery and Resilience Plan (NRRP) – MISSION 4 COMPONENT 2, INVESTMENT N.
+1.4 – CUP N.
+I53C22000690001.
+This work has been supported by the project “Artificially devised many-body quantum dynamics in low dimensions - ManyQLowD” funded by the MIUR Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN) Bando 2022 - grant 2022R35ZBF.
+This work has been partially funded by the ERC Starting Grant 101042293 (HEPIQ).
+Appendix A $GGE$ thermodynamic macrostate in the $XXZ$ spin chain
+Here we provide a self-contained discussion of the derivation of the $GGE$ for a generic pre-quench initial state. We follow Ref. [25, 26].
+In particular, we show that the information about the expectation values of local and quasilocal conserved quantities over the pre-quench initial state allows to fix univocally the $TBA$ densities identifying the $GGE$ that describes the steady state after the quench.
+We start from the Algebraic Bethe Ansatz treatment of the $XXZ$ chain. A central object in the Algebraic Bethe Ansatz [33] for the $XXZ$ chain is the spin-$j/2$ Lax operator $L_{j}(\mu)$ defined as
 $L_{j}(\mu)=\frac{1}{\sinh(\eta)}\left(\begin{array}[]{cc}\sin(\mu+i\eta s^{z})&i\sinh(\eta)s^{-}\\ i\sinh(\eta)s^{+}&\sin(\mu-i\eta s^{z})\end{array}\right).$ (28)
@@
 $\mathbb{L}_{j}^{(k)}(x,\mu)=\frac{L_{j}^{(k),[-]}(\mu)L_{j}^{(k),[+]}(\mu+x)}{L_{0}^{(k),[-j-1]}(\mu)L_{0}^{(k),[j+1]}(\mu+x)}.$ (51)
-4 Generalized Hydrodynamics ($GHD$) for quenches from inhomogeneous initial states
-Here we introduce the framework of Generalized Hydrodynamics [13, 12] ($GHD$) to describe the expansion dynamics (see Fig.
-1).
-This will allow us to characterize the bound-state confinement in section 5.
-In the standard $GHD$ setup two semi-infinite homogeneous but different initial states are joined together at $t=0$ at the origin $x=0$.
-At $t>0$ the full system evolves under a homogeneous integrable Hamiltonian.
-For typical initial states, at long times and large distances with the ratio $\zeta:=x/t$ fixed, the behavior of the system is described by a $\zeta$-dependent $GGE$.
-This $GGE$ is fully characterized by a set of $TBA$ densities $\rho_{n,\zeta}(\lambda)$ and $\rho_{h,n,\zeta}(\lambda)$.
-The $TBA$ densities satisfy the $TBA$ equations (3) for any $\zeta$ and the $GHD$ equation, which reads as [12, 13, 17]
-$[\zeta-v_{n,\zeta}(\lambda)]\partial_{\zeta}\vartheta_{n,\zeta}(\lambda)=0,\quad\mathrm{with}\,\zeta:=x/t.$ (35)
-Here $\vartheta_{n,\zeta}$ are the filling functions (cf. (36)) defined as
-$\vartheta_{n,\zeta}(\lambda)=\frac{\rho_{n,\zeta}(\lambda)}{\rho_{t,n,\zeta}(\lambda)},$ (36)
-with $\rho_{t,n,\zeta}=\rho_{n,\zeta}+\rho_{h,n,\zeta}$ the total density. In (35), $v_{n,\zeta}$ are the dressed velocities, which are defined as
-$v_{n,\zeta}=\frac{[\partial_{\lambda}\varepsilon_{n,\zeta}]^{\mathrm{dr}}(\lambda)}{[\partial_{\lambda}p_{n,\zeta}]^{\mathrm{dr}}(\lambda)},$ (37)
-where $\varepsilon_{n,\zeta}(\lambda)$ and $p_{n,\zeta}(\lambda)$ are respectively the bare energy and quasimomentum of the quasiparticles, defined in (9).
-Given a generic “bare” quantity $\tau^{\mathrm{b}}_{n,\zeta}(\lambda)$, its dressed version $\tau_{n,\zeta}^{\mathrm{dr}}$ is obtained by solving the integral equation as [32, 30]
-[…]
-where $\partial_{\lambda}\Theta_{n,m}(\lambda)$ is defined in (7), and the filling function $\vartheta_{n,\zeta}(\lambda)$ is given in (36).
-It is clear from the definition of the momentum $p_{n,\zeta}$ as [2]
-$\partial_{\lambda}p_{n,\zeta}(\lambda)=2\pi\rho_{t,n,\zeta}(\lambda),$ (39)
-that the dressing of $\partial_{\lambda}p_{n}$ is obtained by considering $\tau^{\mathrm{dr}}_{n,\zeta}=2\pi\rho_{t,n,\zeta}(\lambda)$ in (38).
-To solve (35) one has to provide the boundary conditions at $\zeta=\pm\infty$.
-Let us consider the prototypical setup in which the initial state is obtained by joining two semi-infinite chains prepared in two macroscopically different states.
-This is the so-called bipartite quench protocol [17].
-Let us denote by $\vartheta_{n,-\infty}$ and $\vartheta_{n,+\infty}$ the two thermodynamic macrostates, i.e., the $TBA$ filling functions, describing the steady state arising at infinite time after the quench from the left and right homogeneous states.
-The functions $\vartheta_{n,\pm\infty}$, provide the boundary conditions in (35).
-Before proceeding, one can observe that Eq.
-(35) can be solved for each $\zeta$ separately.
-It is straightforward to check that the solution of (35) can be written as
-$\vartheta_{n,\zeta}(\lambda)=\Theta_{\mathrm{H}}(v_{n,\zeta}-\zeta)[\vartheta_{n,-\infty}(\lambda)-\vartheta_{n,+\infty}(\lambda)]+\vartheta_{n,+\infty}(\lambda).$ (40)
-where $\Theta_{\mathrm{H}}(x)$ is the Heaviside step function.
-The dressed velocity $v_{n,\zeta}$ in (40) is obtained by using (38) and (37) (39).
-Notice that since $v_{n,\zeta}$ depends on $\vartheta_{n,\zeta}$ via (38), Eq.
-(40) is not an explicit solution of (35).
-To obtain $\vartheta_{n,\zeta}$ one can use an iterative strategy, starting from (40) fixing $v_{n,\zeta}=v^{b}_{n,\zeta}$, with $v^{b}_{n,\zeta}$ the “bare” velocities (cf.
-(8)).
-Then, one substitutes Eq.
-(40) in (38) to obtain the dressed $\partial_{\lambda}\varepsilon_{n,\zeta}$.
-By using the $TBA$ equation (3) one obtains $\rho_{t,n,\zeta}$ and the new dressed velocities via (37).
-Finally, these steps are iterated until convergence.
-Notice that although the setup of the bipartite quench is different from the one described in Fig.
-1 (a), the two protocols give the same dynamics provided that $\ell$ (see Fig.
-1) is large enough to ensure that the bulk of region $A$ is in the hydrodynamic limit and $1\ll v_{\mathrm{max}}t\ll\ell$ is large, but short enough to neglect the fact that region $A$ is finite.
-Here $v_{\mathrm{max}}$ is the maximum velocity in the system, i.e, $v_{\mathrm{max}}=\max_{n,\zeta,\lambda}v_{n,\zeta}(\lambda)$.
-In the following we focus on the macrostate with $\zeta=0$, which describes the interface between the two chains.
-We also consider the situation in which the right chain is prepared in the vacuum state (see Fig.
-1).
-This corresponds to $\vartheta_{n,+\infty}=0$ for any $n$.
-Now, Eq.
-(40) becomes
-$\vartheta_{n,0}(\lambda)=\Theta_{\mathrm{H}}(v_{n,0})\vartheta_{n,-\infty}(\lambda).$ (41)
-In the following we neglect the subscript $\zeta$ in $\vartheta_{n,\zeta}$ and $v_{n,\zeta}$ since we always consider $\zeta=0$.
-5 Bound-state confinement after trap-expansion experiments
-Here we discuss the confinement of the bound states for the expansion protocol of Fig.
-1 for generic initial state of part $A$ (see Fig.
-1).
-The state of $A$ corresponds to fixing $\vartheta_{n,-\infty}$ in (41).
-Let us summarize our main results.
-We show that for moderately large $\Delta$ ballistic transport of bound states is inhibited, for quite generic initial states of part $A$.
-Specifically, in the hydrodynamic limit only quasiparticles with $n=1$, i.e., unbound quasiparticles, can be present at $\zeta\geq 0$.
-This means that all the bound states with $n>1$ remain confined within region $A$ in Fig.
-1 (a).
-Bound-state confinement can be understood perturbatively in the large $\Delta$ limit, and it relies on the behavior of the scattering phases (7).
-Indeed, the only assumption on the initial state of region $A$ is that it gives rise to $TBA$ macrostates with ${\mathcal{O}}(1)$ density of quasiparticles with $n=1$.
-Although bound-state confinement is established in the large $\Delta$ limit, it survives up to $\Delta\approx 1$.
-Precisely, for the Néel state bound state confinement survives up to [15] $\Delta=\Delta^{*}\approx 1.4$.
-Upon lowering $\Delta$ below $\Delta^{*}$, bound states with $n=2$ start to leave region $A$.
-For $\Delta\ll\Delta^{*}$ bound states with $n>2$ are not confined.
-However, while the threshold $\Delta^{*}$ for confinement of the two-particle bound states is understood perturbatively, this is not the case for the larger bound states.
-Indeed, the deconfinement of bound states with $n>2$ happens in the nonperturbative regime at $\Delta\approx 1$, and is sensitive to the details of the initial state.
-Crucially, since region $A$ is finite (see Fig.
-1), the condition that the density of quasiparticles with $n=1$ is finite in $A$, which ensures the bound-state confinement, is necessarily violated at long times.
-This implies that at times tproportional-$t\propto\ell$ excitations with $n>1$ can leave region $A$, i.e., they are not confined anymore.
-Bound-state confinement manifests itself in the fact that the dressed velocities $v_{n}(\lambda)$ at $\zeta=0$ are negative for any $\lambda$ and for any $n>1$.
-To proceed, let us assume that only bound states with $n>n_{a}$, for some $n_{a}>1$, are confined.
-This implies that Eq.
-(38) for $\tau_{n}=\partial_{\lambda}\varepsilon_{n}$ becomes
-$\partial_{\lambda}\varepsilon^{\mathrm{dr}}_{n}(\lambda)=\partial_{\lambda}\varepsilon^{\mathrm{b}}_{n}(\lambda)-\sum_{m=1}^{n_{\mathrm{a}}}\int\!\!\frac{d\mu}{2\pi}\partial_{\lambda}\Theta_{n,m}(\lambda-\mu)\vartheta_{m,-\infty}(\mu)\Theta_{\mathrm{H}}(\partial_{\mu}\varepsilon^{\mathrm{dr}}(\mu))\partial_{\lambda}\varepsilon_{m}^{\mathrm{dr}}(\mu),$ (42)
-where $n\leq n_{\mathrm{a}}$, and $n_{\mathrm{a}}$ is such that $\partial_{\lambda}\varepsilon^{\mathrm{dr}}_{n}(\lambda)<0$ for any $\lambda$ and for any $n>n_{\mathrm{a}}$.
-Notice that in (42) we replaced $v_{n,0}$ with $\partial_{\lambda}\varepsilon_{n}^{\mathrm{dr}}$ in the argument of the Heaviside theta function.
-This is possible because $\rho_{t,n}(\lambda)$ (cf.
-(37)) is positive.
-Also, notice that in the integrand in (42) one has that $\partial_{\lambda}\Theta_{n,m}$ and $\vartheta_{n,-\infty}$ are both positive functions.
-For (42) to be consistent with confinement of the bound states with $n>n_{a}$, we require that the solutions of (42) satisfy the condition
-$\partial_{\lambda}\varepsilon_{n}^{\mathrm{dr}}(\lambda)<0,\forall n>n_{\mathrm{a}},\forall\lambda,$ (43)
-Clearly, while the dressing (38) for the $XXZ$ with $\Delta>1$ is a system of infinite equations since $n\in[1,\infty)$, the system (42) is finite.
-Still, given the solutions of (42), one has to check the validity of (43) for any $n>n_{a}$.
-Before proceeding, it is interesting to observe that $n_{\mathrm{a}}>0$, meaning that not all the species of quasiparticles can be confined at the same time, at least in the hydrodynamic limit.
-Indeed, $n_{\mathrm{a}}=0$ implies (cf.
-(42)) that $\partial_{\lambda}\varepsilon^{\mathrm{dr}}(\lambda)=\partial_{\lambda}\varepsilon^{\mathrm{b}}(\lambda)$, and $\partial_{\lambda}\varepsilon^{\mathrm{b}}(\lambda)$ is not negative for all values of $\lambda$ (cf.
-(8)).
-This implies that the suppression of ballistic transport observed for instance in the domain-wall quench [23] cannot be attributed to the bound-state confinement.
-The conditions under which the systems (42) and (43) admit a solution can be derived in the large $\Delta$ limit.
-First, let us observe that $0\leq\vartheta_{n,-\infty}(\lambda)\leq 1$ for any $\lambda$.
-Moreover, in the large $\Delta$ limit, at the leading order in powers of $\Delta$, one has that $\partial_{\lambda}\Theta_{n,m}(\lambda-\mu)={\mathcal{O}}(1)$ for any $n,m$.
-Precisely, for large $\Delta$ one has that
-$\partial_{\lambda}\Theta_{n,m}(\lambda)\to\frac{1}{\pi}(n+m-1),\quad\mathrm{for}\,\,n,m\neq 1,$ (44)
-and $\partial_{\lambda}\Theta_{n,m}(\lambda)\to(n+m)/\pi$ otherwise.
-Finally, it is straightforward to check that the bare velocities $v_{n}^{\mathrm{b}}$ (cf.
-(8)) at the leading order in the large $\Delta$ limit are
-$v_{n}^{\mathrm{b}}(\lambda)=2z^{n-1}\sin(2\lambda)+{\mathcal{O}}(z^{n+1}),\quad\mathrm{with}\,\,z:=e^{-\eta}.$ (45)
-Here we defined $z=\Delta-\sqrt{\Delta^{2}-1}$, and for large $\Delta$ one has $z\approx 1/(2\Delta)$.
-Moreover, unlike the dressed velocities, the bare ones do not depend on the pre-quench initial state, so (45) holds irrespective of the initial state of $A$ (see Fig.
-1).
-Crucially, while the bare velocities of the quasiparticles with $n=1$ are ${\mathcal{O}}(1)$, the bound-state velocities for $\Delta>1$ decay exponentially with the bound-state size $n$.
-This hierarchy in the bare velocities is responsible for the bound-state confinement.
-Let us consider the case with $\vartheta_{1,-\infty}(\lambda)={\mathcal{O}}(1)$, i.e., the situation in which in region $A$ there is a finite density of unbound quasiparticles.
-This happens for typical states, such as the Néel state and the Majumdar-Ghosh state [15].
-We anticipate that the $p$-Néel states with $p>1$ do not satisfy this condition.
-Let us also assume that $n_{\mathrm{a}}=1$, i.e., bound states with $n\geq 2$ remain confined within $A$ during the expansion dynamics.
-Now, let us consider Eq.
-(42) with $n=2$.
-It is clear that the second term in the right-hand side dominates for large $\Delta$ because $\partial_{\lambda}\varepsilon_{n}^{\mathrm{b}}={\mathcal{O}}(z)$ for $n=2$, whereas the first term in (42) is ${\mathcal{O}}(z)$.
-Clearly, the same argument applies also for $n>2$, implying that all the bound states with $n\geq 2$ are confined for sufficiently large $\Delta$.
-Interestingly, as the only dependence on $\lambda$ in the second term in (42) is in $\Theta_{n,m}$, which becomes independent of $\lambda$ for large $\Delta$, the dressed velocities of the confined bound states become “flat” as functions of $\lambda$.
-Upon lowering $\Delta$, the two terms in (42) become of the same order, and quasiparticles with $n=2$ start to be transported in $B$ (see Fig.
-1).
-Notice that while bound-state confinement can be established perturbatively, and even the threshold $\Delta^{*}$ at which bound states with $n=2$ are not confined can be calculated from the large $\Delta$ expansion, it is difficult to characterize analytically the scenario for $\Delta\to 1$.
-Typically, we observe that there is a “cascade” effect.
-Precisely, upon lowering $\Delta$, transport of bound states with progressively larger $n$ become allowed (see section 5.1).
-The scenario outlined so far relies on the fact that at any time during the expansion $\vartheta_{1,-\infty}={\mathcal{O}}(1)$.
-This condition remains true at arbitrarily long times if region $A$ (cf.
-Fig.
-1) is infinite.
-Oppositely, if region $A$ is finite, at long times one has that $\vartheta_{1,-\infty}$ vanishes.
-This implies that the second term in (42) becomes smaller.
-Thus, at long times the bound states with $n=2$ can leave region $A$ after the “evaporation” of the unbound particles.
-Moreover, the condition $\vartheta_{1,-\infty}$ is not valid for generic initial states.
-For instance, it is violated by all the $p$-Néel states with generic $p>1$.
-By using the framework of section 3, one can check that $\vartheta_{n,-\infty}={\mathcal{O}}(z^{2})$ for $n<p$, whereas $\vartheta_{n,-\infty}={\mathcal{O}}(1)$ for $n\geq p$.
-This has striking consequences on transport properties and bound-state confinement.
-Indeed, we anticipate that the fact that $\vartheta_{n,-\infty}={\mathcal{O}}(z^{2})$ implies a strong suppression of bound-state transport for $p\geq 4$.
-Let us first consider the case with $p=2$.
-Now, Eq.
-(42) suggests that bound states with $n=2$ are not confined.
-Indeed, the term with $m=1$ in the right-hand-side of (42) is ${\mathcal{O}}(z^{2})$ and it is negligible compared to $\partial_{\lambda}\varepsilon_{2}^{\mathrm{b}}={\mathcal{O}}(z)$.
-This means that unlike the case with $p=1$ the two-particle bound states cannot be confined by the magnons.
-Now, if we consider the case with $p=4$ we observe that since $\partial_{\lambda}\varepsilon_{4}^{\mathrm{b}}={\mathcal{O}}(z^{3})$, it is clear that the second term in (42) dominates because of the scatterings with the quasiparticles with $n\leq 4$, which contribute with terms with smaller powers of $z$.
-This implies that upon increasing $p$ the bound states that dominate the thermodynamic properties, i.e., the bound states with $n=p$, become confined, and transport is suppressed.
-5.1 Bound-state confinement: Numerical GHD results
-Figure 2: Group velocities $v_{n}$ of the $n$-particle bound states calculated over the $GGE$ with $\zeta:=x/t=0$, which in the hydrodynamic limit describes the interface between region $A$ and the rest (Fig.
-1).
-At $t=0$ region $A$ is prepared in the $p$-Néel states (cf.
-(2)).
-The different panels correspond to different values of $p$ and $\Delta$.
-Panel (a) and (b) are for $p=1$.
-Panels (c-f) show results for $p=2-5$.
-In all the panels one has that $v_{1}(\lambda)$ is positive in an extended region in $\lambda$, signaling that magnon-like excitations are never confined.
-(a) The case of $p=1$, i.e., the Néel state.
-At $\Delta\lesssim 1.4$ $v_{2}<0$ for any $\lambda$, i.e., the two-particle bound states are confined.
-(b) Same as in (a) for $\Delta=1.01$.
-While bound states with $n=2$ are not confined, $v_{3}\leq 0$ for any $\lambda$, which means that three-particle bound states are confined.
-(c) The $p$-Néel state with $p=2$.
-Bound states with $n\geq 3$ are confined.
-At $\Delta\approx 1.09$ three-particle bound states are not confined anymore.
-(d) The case with $p=3$.
-At $\Delta\approx 1.02$ the quasiparticles with $n=4$ are not confined.
-(e) The case with $p=4$.
-Now, even quasiparticles with $n=4$ remain confined, at least up to $\Delta\approx 1.07$.
-(f) The Néel state with $p=5$.
-Similar to (e), quasiparticles with $n=p$ are confined.
-Although qualitative features of bound-state confinement are understood perturbatively, to characterize quantitatively bound-state confinement it is necessary to determine numerically the dressed velocities $v_{n}(\lambda)$.
-Here we discuss numerical results confirming the scenario outlined in section 5.
-Precisely, we numerically solve the $GHD$ equation (35) to obtain the dressed velocities $v_{n}$ for $\zeta=0$.
-The numerical solution of (35) is standard.
-By using a Gauss discretization, Eq.
-(35) becomes an infinite system of nonlinear equations, which can be truncated considering only bound states of size $n\leq n_{\mathrm{max}}$.
-Finally, the truncated system can be solved with standard methods.
-Here we consider $n_{\mathrm{max}}=40$.
-Our results are reported in Fig.
-2.
-We plot $v_{n}(\lambda)$ for $1\leq n\leq 5$ versus $\lambda$.
-We consider as initial states the $p$-Néel states (cf.
-(2)) with $1\leq p\leq 5$.
-For $p=1$ the quench is the popular Néel quench.
-At large $\Delta$ one has that $v_{n}<0$ for any $\lambda$ and for any $n>1$, meaning that all the bound states are confined within region $A$.
-Fig.
-2 (a) shows that for $\Delta\approx 1.4$ the faster bound states with $n=2$ start to be transmitted [15].
-Upon lowering $\Delta$, at $\Delta\approx 1.01$, the bound states with $n=3$ are not confined, as shown in Fig.
-2 (b).
-For $p=2$, as anticipated in section 5, the bound states with $n=2$ are not confined.
-We checked the absence of confinement up to $\Delta=5$, although we do not report the results in Fig.
-2.
-On the other hand, Fig.
-2 (c) shows that for $p=2$ at $\Delta\approx 1.1$ the three-particle bound states are still confined in $A$.
-A similar scenario occurs for $p=3$, i.e., the bound states with $n\leq 4$ are not confined, whereas the bound states with $n>4$ are all confined up to $\Delta\approx 1.02$.
-However, the scenario changes at $p=4$.
+Appendix B Convergence of $tDMRG$
+Here we discuss the convergence of the $tDMRG$ algorithm.
+We consider the expansion of the Néel state, i.e., the $p$-Néel state with $p=1$ focussing on the $XXZ$ chain with $\Delta=1$.
+Typically, we observe that the choice $\Delta=1$, $p=1$ gives rise to the strongest entanglement growth.
+We discuss the convergence of the projector $P_{x,\uparrow\uparrow}$, as it is relevant to assess the confinement of the two-particle bound states.
+In Fig.
+11 we plot the time-dependent expectation value $\langle P_{x,\uparrow\uparrow}\rangle$ versus $\zeta=x/t$.
+The region $A$ prepared in the Néel state has $\ell=40$ sites.
+We show data at fixed time $t=16$.
+In our $tDMRG$ simulations we employ a seventh-order Trotter decomposition (see [45]) of the evolution operator, with time step $\delta t=0.2$.
+The different symbols in Fig.
+11 correspond to different values of the bond dimension $\chi$ up to $\chi=800$.
 As it is clear from Fig.
-2 (e), for $p\geq 4$ the group velocities $v_{p}(\lambda)$ are negative for any $\lambda$.
-This holds true for $p=5$ as well (see Fig.
-2 (f)).
-Since the thermodynamic properties of the system are dominated by the bound states with $n=p$, this means that transport becomes more and more suppressed upon increasing $p$.
-6 Numerical tDMRG results
-Here we discuss $tDMRG$ data [19] for the vacuum expansion protocols introduced in Fig.
-1.
-Our results support the bound-state confinement scenario outlined in section 5.
-As a preliminary check, in section 6.1 we discuss $tDMRG$ data for full-expansion protocol $(a)$ in Fig.
-1, considering times tmuch-less-$t\ll\ell$, such that the fact that region $A$ is of finite size can be neglected.
-We benchmark the results of section 4 for several $p$-Néel initial states.
-In section 6.2 we focus on the confinement of bound states for the full-expansion protocol (see Fig.
-1 (a)).
-In section 6.3 we discuss the cut quench protocol (see Fig.
-1 (b)).
-Figure 3: Releasing the $p$-Néel states in the vacuum (see Fig.
-1).
-The figure shows $tDMRG$ data for the local magnetization $\langle\sigma_{x}^{z}\rangle$ plotted as a function of $\zeta=x/t$.
-Here $x$ is distance from the interface between region $A$ and the rest (Fig.
-1).
-The dashed line is the $GHD$ result valid in the limit $x,t\to\infty$ at fixed $\zeta$.
-We show results for $p\leq 3$ and $\Delta=1,2$.
-The different symbols denote different times $t$.
-Notice that for $p=3$ and $\Delta=2$ $t,x$ are not large enough to reach the hydrodynamic limit.
-Figure 4: Releasing the Néel states in the vacuum (see Fig.
-1).
-We plot the expectation value $\langle P_{x,\uparrow\uparrow}\rangle$ (cf.
-(47)) versus the rescaled time $\zeta=x/t$, with $x$ the distance from the interface between region $A$ and the rest (see Fig.
-1).
-The size of region $A$ is $\ell=40$ sites.
-Panels (a) and (b) show the expansion dynamics for $\Delta=1$ and $\Delta=2$, respectively.
-The different symbols denote different times.
-We only show results up to $t\lesssim 15$, to ensure that in the bulk of region $A$ one has $\vartheta_{1,-\infty}={\mathcal{O}}(1)$, which implies bound-state confinement (see section 5).
-Notice that at $\zeta=0$ for $\Delta=1$ one has a finite value for $\langle P_{x,\uparrow\uparrow}\rangle$, whereas at $\Delta=2$ one has $\langle P_{x,\uparrow\uparrow}\rangle\approx 0$, reflecting bound-state confinement.
-6.1 Validity of GHD
-Before investigating bound-state confinement, it is crucial to assess the validity of the $GHD$ description for the protocols discussed in Fig.
-1.
-This is summarized in Fig.
-3.
-We consider the expansion of the $p$-Néel states with $1\leq p\leq 3$.
-The different panels show $tDMRG$ data for the local magnetization $\sigma_{x}^{z}$ as a function of $\zeta=x/t$.
-For each $p$ we show results for $\Delta=1,2$.
-Here $x$ is the distance from the right edge of region $A$ (see Fig.
-1).
-In our $tDMRG$ simulations we consider times tmuch-less-$t\ll\ell$ to avoid effects due to the finite size of $A$.
-In Fig.
-3 we provide $tDMRG$ data for several times up to $t\approx 20$ (different symbols in the Figure).
-The $tDMRG$ data exhibit collapse when plotted as function of $x/t$, confirming that $t\approx 20$ is large enough to ensure the hydrodynamic limit for all the values of $p$ and $\Delta$ considered, except for $p=3$ with $\Delta=2$ (see Fig.
-3 (f)).
-For $p=3$ and $\Delta=2$ the evolution in the bulk of region $A$ is slower compared to the other cases.
-Indeed, at $\zeta\leq-1$, the magnetization profile exhibits strong oscillations, which are reminiscent of the structure of the initial state.
-In all the panels, the dashed curve are the $GHD$ results obtained by solving numerically the $GHD$ equation (35).
-The numerical results were obtained by truncating the infinite system of equations (35) keeping only the bound states with $n\leq n_{\mathrm{max}}=40$.
-The $GHD$ equation (35) allows to obtain the occupation functions $\vartheta_{\zeta=0,n}(\lambda)$.
-The numerical results for $\vartheta_{0,n}(\lambda)$ can be substituted in the $TBA$ equations (3), after using that $\rho_{t,n}(\lambda)=\vartheta_{n}(\lambda)\rho_{n}(\lambda)$.
-Eq.
-(3) allows one to obtain the quasiparticle densities $\rho_{n}(\lambda)$, which encode information about local observables.
-For instance, the local magnetization is given as
-[…]
-For all the quenches that we explore in panels (a-e) of Fig.
-3 the $GHD$ results are in satisfactory agreement with the $tDMRG$ data, despite the limitation to times $t\lesssim 20$.
-For the expansion of the $p$-Néel state with $p=3$ and $\Delta=2$ (panel (f) in Fig.
-3) deviations from the $GHD$ prediction are large.
-Also, notice that for $\zeta>0$ one has that $\langle\sigma_{x}^{z}\rangle\approx-1$, signaling that transport in general is strongly suppressed for $p=3$.
-As anticipated in section 5 for the $p$-Néel state with $p=1$, the bound states with $n=2$ remain confined in region $A$, as long as the system is in the hydrodynamic limit, and the $GHD$ description is valid.
-Specifically, this holds for $x,t\to\infty$ with $x/t$ fixed, with $x$ being the distance from the right edge of region $A$ (see Fig.
-1).
-Moreover, for the bound-state confinement to happen, the condition $\vartheta_{1,-\infty}={\mathcal{O}}(1)$ has to be satisfied.
-This is the case for $p=1$, at least if tmuch-less-$t\ll\ell$.
-Confinement of the the two-particle bound states is discussed in Fig.
-4.
-The Figure shows $P_{x,\uparrow\uparrow}$ (cf.
-(47)) plotted versus $\zeta=x/t$.
-The projectors $P_{x,\uparrow^{\otimes m}}$ are defined as
-$P_{x,\uparrow^{\otimes m}}:=\frac{1}{2^{m}}\prod_{j=0}^{m-1}(\mathds{1}+\sigma_{x+j}^{z}).$ (47)
-Although the projectors (47) are not bona fide witnesses of bound states, we anticipate that $P_{x,\uparrow^{\otimes m}}$ is sensitive to the presence of a $m$-particle bound state at position $x$.
-Panels (a) and (b) show $tDRMG$ data for $P_{x,\uparrow\uparrow}$ for the $XXZ$ Hamiltonian with $\Delta=1$ and $\Delta=2$, respectively.
-In each panel the different symbols correspond to different times.
-For both $\Delta=1$ and $\Delta=2$ the $tDMRG$ data exhibit scaling as a function of $\zeta$.
-Crucially, while for $\Delta=1$ (Fig.
-4 (a)) $\langle P_{x,\uparrow\uparrow}\rangle$ is nonzero for $\zeta>0$, one has $\langle P_{x,\uparrow\uparrow}\rangle\approx 0$ for any $\zeta>0$ for $\Delta=2$ (see Fig.
-4).
-This reflects that while for $\Delta=1$ the bound states with $n=2$ are not confined, for $\Delta=2$ they remain confined in region $A$ during the expansion.
-Still, at $\zeta=0$, $\langle P_{x,\uparrow\uparrow}\rangle\approx 0.025$ for $\Delta=1$, which suggests that the density of two-particle bound states that are not confined is “small”, meaning that it could be challenging to observe it, for instance experimentally.
-Figure 5: Releasing the Néel state in the vacuum.
-Regionn $A$ in Fig.
-1 of size $\ell=20$ is prepared in the Néel state at $t=0$.
-The remaining part of the chain is in the vacuum state, i.e., the ferromagnetic state.
-At $t>0$ the chain is let to evolve under the $XXZ$ chain Hamiltonian with $\Delta=2$.
-Panel (a) shows $tDMRG$ data for the local magnetization $\langle\sigma^{z}_{x}\rangle$ at generic position $x$ and at different times.
-Notice the reflection at the boundary of the chain at long times.
-Panels (b) and (c) show the projectors $P_{x,\uparrow\uparrow}$ and $P_{x,\uparrow\uparrow\uparrow}$.
-Figure 6: Releasing the Néel state in the vacuum.
-Region $A$ (with $\ell=20$ sites) of a chain is prepared in the Néel state at $t=0$.
-The remaining part of the chain is in the vacuum state, i.e., the ferromagnetic state.
-At $t>0$ the chain is let to evolve under the $XXZ$ chain Hamiltonian with $\Delta=2$.
-Panel (a) shows $tDMRG$ data for the local magnetization $\langle\sigma^{z}_{x}\rangle$ at generic position $x$ and at different times.
-Notice the reflection at the boundary of the chain at long times.
-Panels (b) and (c) show the projectors $P_{x,\uparrow\uparrow}$ and $P_{x,\uparrow\uparrow\uparrow}$, respectively (cf.
-(47)).
-The black lines in the panels are the maximum bare velocities $\max_{\lambda}v_{n}^{b}(\lambda)$ (cf.
-(8)) for $n=1$ (in panel (a)), $n=2$ (panel (b)), and $n=3$ (panel (c)).
-6.2 Full expansion dynamics
-In section 6.1 we established the validity of the $GHD$ description for the quenches from the $p$-Néel states.
-Moreover, Fig.
-3 suggests that $t\approx 20$ is enough to access the hydrodynamic limit, which is essential to observe bound-state confinement.
-We now focus on the full expansion dynamics, i.e., considering times tmuch-greater-$t\gg\ell$.
-We show $tDMRG$ data for the expansion of the $p$-Néel states with $p=1$ for $\Delta=1$ and $\Delta=2$ in Fig.
-5 and Fig.
-6, respectively.
-The protocol is reported in Fig.
-1.
-Region $A$ at center of the chain is prepared in the Néel state.
-Here we consider $\ell=20$.
-The reason for choosing $\ell=20$ is that it allows to simulate the dynamics using $tDMRG$ up to times $t\approx 100$.
-Indeed, entanglement, which is the major limitation of $tDMRG$, is produced in the bulk of region $A$ only, whereas the expansion does not contribute significantly to the entanglement growth.
-This also reflects that the $GHD$ dynamics preserves the thermodynamic entropy [14, 16].
-As a consequence, the largest entanglement produced during the dynamics is bounded by $\ell$.
-The different panels in Fig.
-5 and Fig.
-6 show different observables.
-Precisely, we consider the local magnetization $\sigma_{x}^{z}$, and the local projectors $P_{x,\uparrow\uparrow},P_{x,\uparrow\uparrow\uparrow}$ (cf.
-(47)).
-As anticipated, and has already observed in Ref.
-[22] and Ref.
-[10], $P_{x,\uparrow^{\otimes m}}$ is sensitive to bound states of $m$ spins starting from position $x$.
-Let us first discuss the dynamics with the $XXZ$ chain at $\Delta=1$ (see Fig.
-5).
-In Fig.
-5 we plot the density of local magnetization $\langle\sigma_{x}^{z}\rangle$ as a function of position $x$ and time $t$.
-After the expansion starts, the central region “melts” as quasiparticles are emitted at its edges.
-First, several “jets” of excitations are visibile, traveling at different velocities.
-The fastest quasiparticles correspond to the magnon-like excitations, i.e., with $n=1$, whereas the jets moving at smaller velocities correspond to bound states with $n=2$ and $n=3$ quasiparticles.
-The dashed lines are the bare maximum group velocities of the quasiparticles (cf.
-(8)).
-Since the initial state is expanded in the vacuum, the density of quasiparticles at the edges of the jets is small and interactions are “weak”, which explains why the bare velocities describe the dynamics of the jets.
-In Fig.
-5 (b) we show the dynamics of the projector $P_{x,\uparrow\uparrow}$.
-Now, the outermost jet in Fig.
-5 (a) is not present, confirming that it is due to the quasiparticles with $n=1$.
-Oppositely, the two innermost jets of Fig.
-5, corresponding to $n=2,3$ are still present.
-Similar behavior was observed in the expansion of few-spin initial states in Ref.
-[22].
-Importantly, one should observe that at short times $t\lesssim 20$, $P_{x,\uparrow\uparrow}$ exhibits a “halo” feature at the edges of region $A$.
-This can be attributed to the fact that for $\Delta=1$ the two-particle bound states are not confined in the hydrodynamic limit.
-Notice, however, that the halo is quite weak, reflecting that the density of two-particle bound states at the interface between $A$ and the rest is small.
-A similar scenario is observed for $P_{x,\uparrow\uparrow\uparrow}$ (see Fig.
-5 (c)).
-Now, the three-particle bound states behave as if they were confined up to $t\approx 20$.
-Notice that the fact (see Fig.
-2 (b)) that $v_{3}\leq 0$ at $\Delta\approx 1.01$ suggests that the density $\rho_{3}(\lambda)$ of three-particle bound states at the interface between $A$ and the vacuum could be small.
-This implies that even though bound states with $n=3$ are not confined at $\Delta=1$, this could be difficult to observe.
-Let us now discuss the same protocol in the $XXZ$ chain with $\Delta=2$ (see Fig.
-6).
-The qualitative behavior is the same as in Fig.
-5.
-However, the projector $P_{x,\uparrow\uparrow}$ at short times $t\lesssim 20$ exhibits a “bottleneck” feature.
-In contrast with $\Delta=1$, at $\Delta=2$ the two-particle bound states are confined (see Fig.
-2 (c)).
-Moreover, in the hydrodynamic regime the density $\rho_{2}(\lambda)$ is zero for $\zeta=0$, and it is presumably small also at $\zeta\lesssim 0$.
-This explains the bottleneck in Fig.
-6 At long times after all the magnons have left region $A$, the two-particle bound states are allowed to leave region $A$, as discussed in section 5, and as it is confirmed in Fig.
-6 (b).
-A similar dynamics happens for the three-particle bound states, as shown in Fig.
-6 (c).
-Figure 7: Releasing the Néel state in the vacuum.
-Same setup as in Fig.
-5 and Fig.
-6.
-We show the magnetization profile $\langle\sigma_{x}^{z}\rangle$ as a function of the position $x$ in the chain.
-The different curves correspond to different times.
-Notice the different peaks, which correspond to different families of excitations.
-Precisely, the outermost peaks correpsond to magnon-like excitations, whereas the innermost ones to the bound states with $n\geq 2$.
-Let us now discuss the space-time profiles of local observables.
-In Fig.
-7 we show the profile of the local magnetization $\langle\sigma_{x}^{z}\rangle$ after the expansion of the Néel state for the $XXZ$ chain with $\Delta=1$ and $\Delta=2$, respectively.
-Both panels in Fig.
-7 confirm that for $\Delta=1$ and $\Delta=2$ the bound states with $n=2$ and $n=3$ are not confined at long times, as already observed in Fig.
-5 and Fig.
-6.
-The $tDMRG$ data for $\Delta=1$ (left panel in Fig.
-7) show a rapid melting of the profile.
-As the dynamics proceeds, $\langle\sigma_{x}^{z}\rangle$ exhibits a multi-peak structure, each peak corresponding to a different bound-state family.
-For $\Delta=1$ the peaks are blended together, reflecting that the bound states of different types start to leave region $A$ at the same time.
-In particular, already at $t=40$ the bound states with $n=2$ are moving towards the edges of region $A$.
-In the inset we show the dynamics of $P_{x,\uparrow\uparrow}$, which is sensitive to quasiparticles with $n\geq 2$ only.
-The inset shows a rapid melting of $P_{x,\uparrow\uparrow}$.
-In contrast, for $\Delta=2$ (right panel in Fig.
-7) the dynamics is slower, and the peaks corresponding to different $n$ are well separated.
-Notice that at $t=80$ the lump around $x=100$ corresponds to the three-particle bound states, which remain almost locked at the center of the chain.
-This observation suggests that for $\Delta=2$ the trap-expansion protocol can be used to “distillate” states with large bound-state content.
-The inset shows the dynamics of $P_{x,\uparrow\uparrow}$.
-Notice that for $t\lesssim 10$ the bound states with $n=2$ are still within region $A$.
-Importantly, for $\ell=20$ the density of quasiparticles with $n=1$ in the bulk of $A$ starts to be small already at $t\approx 10$, implying that bound-state confinement is not supported anymore (see section 5).
-Figure 8: Releasing the Néel state in the vacuum.
-At $t=0$ the central part $A$ (see Fig.
-1) of a spin chain containing $\ell=20$ sites is prepared in the Néel state.
-At $t>0$ the chain is let to evolve under the $XXZ$ chain Hamiltonian with $\Delta=2$.
-For $0<t\leq 20$ the full chain evolves.
-At $t>20$ only the $B$ part (see Fig.
-1) is evolved.
-Panels (a)-(d) show $tDMRG$ data for the local magnetization $\langle\sigma_{x}^{z}\rangle$ as a function of the position $x$ in the chain.
-Panels (e)-(h) show results for the projector $P_{x,\uparrow\uparrow}$.
-Notice that at $\Delta\approx 1.4$, $P_{x,\uparrow\uparrow}$ vanishes, signalling bound-state confinement.
-6.3 Dynamics after the cut protocol
-To investigate bound-state confinement, it is also convenient to employ the protocol in Fig.
-1 (b).
-The full chain is now evolved with the $XXZ$ Hamiltonian up to a time $t^{*}$.
-Here $t^{*}$ is chosen large enough to ensure the validity of the hydrodynamic limit.
-As it is clear from section 6.1, $t\approx 20$ is sufficient to ensure the hydrodynamic limit.
-Moreover, $t^{*}$ is chosen short enough compared to $\ell$ to have $\vartheta_{1,-\infty}={\mathcal{O}}(1)$ in the bulk of $A$, which is essential to have bound-state confinement.
-Here we choose $t^{*}=20$ and $\ell=40$.
-We consider the $p$-Néel states with $p=1-3$.
-6.3.1 $p$-Néel state with $p=1$
-Our results for the Néel state (i.e., for $p=1$) are shown in Fig.
-8.
-Panels $(a-d)$ shows the local magnetization $\sigma_{x}^{z}$, as a function of position $x$ and time $t$.
-As it is clear from the Figure, for $t>t^{*}$ (see Fig.
-1), under the evolution with the Hamiltonian (1) restricted to $x>0$ the quasiparticles emitted from $A$ propagte ballistically in the vacuum.
-The dependence on $\Delta$ is weak.
-In the bottom row in Fig.
-8 (panels (e-h) in the Figure) we report the dynamics of $P_{x,\uparrow\uparrow}$.
-For $\Delta=1$, $\langle P_{x,\uparrow\uparrow}\rangle$ exhibits a ballistic spreading.
-The nonzero signal reflects that the bound states with $n=2$ are not confined.
-$P_{x,\uparrow\uparrow}$ becomes smaller upon increasing $\Delta$.
-For instance, it is barely visible at $\Delta=2$.
-Notice that in the hydrodynamic limit one should expect that $\langle P_{x,\uparrow\uparrow}\rangle$ is exactly zero for $\Delta>\Delta^{*}\approx 1.4$ (see Fig.
-2).
-6.3.2 $p$-Néel state with $p=2$
-Let us now discuss the cut protocol (see Fig.
-1 (b)) for the $p$-Néel state with $p=2$.
-Our results are reported in Fig.
-9.
-Panels (a)-(c) show $tDMRG$ data for the dynamics with the $XXZ$ chain Hamiltonian and $\Delta=1,2,4$.
-In the Figure we plot the local magnetization $\langle\sigma_{x}^{z}\rangle$ as a function of position $x$ and time $t$.
-For $\Delta=1$ (Fig.
-9 (a)) two distinct signals are visible.
-The two “jets” correspond to quasiparticles with $n=1$, i.e., free ones, and to the two-particle bound states.
-The two-particle jet is due to the fact that for $p=2$ the bound-state content of the chain is enhanced as compared to $p=1$, and two-particle bound states are not confined at $\Delta=1$.
-Notice also that the bound states exhibit smaller group velocities as compared with the quasiparticles with $n=1$.
-The dashed lines reported in Fig.
-9 (a) are the maximum bare velocities $v_{n,\mathrm{max}}^{\mathrm{b}}=\max_{\lambda}(v_{n}^{\mathrm{b}}(\lambda))$ (cf.
-(8)).
-In Fig.
-9 (b) we show data for $\Delta=2$.
-Crucially, the jet corresponding to the two-particle bound states is still visible.
-This is in agreement with the results of section 5.
-Indeed, for $p=2$ one should expect that the quasiparticles with $n=2$ are never confined.
-On the other hand, as it is clear from Fig.
-2 (c), the three-particle bound states are confined for $\Delta\gtrsim 1.1$.
-We should mention that, however, as the density of quasiparticles with $n=3$ is “small”, it is numerically challenging to observe the three-particle bound states confinement.
-Finally, in Fig.
-9 (c) we show results for $\Delta=4$.
-Now, the jet corresponding to $n=1$ is weak, in contrast with the case with $p=1$.
-This is due to the fact that for $p=2$ the density $\rho_{1}(\lambda)$ vanishes, reflecting that thermodynamic properties of the system are dominated by the bound states with $n=2$.
-On the other hand, for $\Delta=4$ the maximum group velocity of the two-particle bound states is $v_{2,\mathrm{max}}^{\mathrm{b}}\approx 0.1$, and their expansion is barely visible.
-Finally, in Fig.
-10 we report the projector $P_{x,\uparrow\uparrow\uparrow}$ after the expansion of the $p$-Néel state with $p=2$.
-In Fig.
-10 (a) and (b) we show results for $\Delta=1$ and $\Delta=2$, respectively.
-While for $\Delta=1$ ballistic spreading is visible, for $\Delta=2$, the projector $P_{x,\uparrow\uparrow\uparrow}$ is “frozen”.
-Figure 9: Releasing the “fattened” $p$-Néel state with $p=2$ in the vacuum.
-The figure shows $tDMRG$ data for the local magnetization $\langle\sigma_{x}^{z}\rangle$ as a function of $x$ and the time $t$.
-The different panels show results for different values of $\Delta$.
-For $\Delta=1$ and $\Delta=2$ the two jets correspond to the magnon-like excitations and the two-particle bound states.
-Figure 10: Releasing the $p$-Néel state with $p=2$ in the vacuum.
-The figure shows $\langle P_{x,\uparrow\uparrow\uparrow}\rangle$ as a function of time and position $x$.
-Panel (a) and (b) show $tDMRG$ data for the $XXZ$ chain with $\Delta=1$ and $\Delta=2$, respectively.
-Notice that for $\Delta=2$ the quasiparticles with $n=3$ are confined in $A$.
-Figure 11: Releasing the Néel state in the vacuum.
-Dynamics of the entanglement entropy.
-The curve shows $tDMRG$ data for a chain with $L=200$ sites.
-The size of region $A$ (see Fig.
-1) is $\ell=20$.
-The plot shows the entropy $S$ of subsystem $B$ (see Fig.
-1).
-Since subsystem $A$ includes the Néel part, the growth of $S$ with time is due to the entanglement between the excitations that leave the Néel part and those that are left behind.
-At long times $S$ measures the entanglement between the “jets” of excitations in Fig.
-5 and Fig.
-6.
-Notice the multistage growth of the entropy.
-At short times the growth is dominated by the entangled magnon-like excitations.
-The second slope visible in the figure is attributed to the two-excitations bound states.
-The last stage of the growth is due to the three-particle bound states.
-7 Bound-state confinement and dynamics of the entanglement entropy
-Here we show that bound-state confinement has striking signatures in the dynamics of the von Neumann entropy.
-We focus on the full expansion of the Néel state (see Fig.
-1 (a)).
-Fig.
-11 shows the von Neumann entropy $S$ of region $B$ (see Fig.
-1 (a)).
-The dotted curve is the result for $\Delta=1$, whereas the full line shows the result for $\Delta=2$.
-First, for both values of $\Delta$ the von Neumann entropy saturates at long time.
-This reflects that the entanglement dynamics is due to the creation of pairs of entangled quasiparticles in region $A$ [29].
-This is the content of the so-called quasiparticle picture [33, 28, 29, 34].
-The two quasiparticles forming the entangled pairs travel with opposite group velocities.
-As they propagate, they entangle larger regions.
-Precisely, the von Neumann entropy of a region is proportional to the number of entangled pairs that are shared between the region and the rest.
-The contribution of a shared pair characterized by rapidities $(-\lambda,\lambda)$ is the contribution of that pair to the thermodynamic entropy of the $GGE$ describing the steady state after the quench from the homogeneous Néel state.
-The quasiparticle picture holds in the hydrodynamic limit $\ell,t\to\infty$ with the ratio $t/\ell$ fixed.
-Let us focus on the results for $\Delta=1$ in Fig.
-11.
-The initial linear growth up to $t\approx 20$ is due to the entangled pairs that are initially produced in $A$ and at later times leave region $A$, and hence start to be shared with the rest (see Fig.
-1).
-The saturation of $S$ at long times reflects that region $A$ is finite, and, as a consequence the number of entangled pairs produced in $A$ is finite, and it is proportional to the size $\ell$ of $A$.
-For $\Delta=1$, $S$ is essentially constant for $t\gtrsim 60$.
-The saturation of the entanglement entropy is the key to reach times long as $t\approx 120$ with $tDMRG$, because after the entropy saturates the computational cost is mild.
-One should also observe that the saturation of $S$ happens “smoothly”.
-This is somewhat reminiscent of the fact that for $\Delta=1$ the bound states are not confined.
-Crucially, for $\Delta=2$, $S$ exhibits a multistage linear growth before saturating.
-This reflects the bound-state confinement.
-Specifically, the linear growth up to $t\approx 20$ is due to the quasiparticles with $n=1$.
-In this early time regime, the slope of the linear growth should be the same as in the vacuum expansion of a semi-infinite chain prepared in the Néel state.
-The prediction of the quasiparticle picture is [14, 16]
-[…]
-where $s_{1,\zeta=0}^{YY}$ is the Yang-Yang entropy [2] of the macrostate with $\zeta=0$ that describes the interface between region $A$ and region $B$ (see Fig.
-1).
-$s_{n,\zeta}^{YY}$ is readily obtained from the particle densities $\rho_{n,\zeta}(\lambda)$ describing the $GGE$ as
-$s_{n,\zeta}(\lambda):=\rho_{t,n,\zeta}\ln(\rho_{t,n,\zeta})-\rho_{n,\zeta}\ln(\rho_{n,\zeta})-\rho_{h,n,\zeta}\ln(\rho_{h,n,\zeta}),$ (49)
-where the densities $\rho_{n,\zeta},\rho_{h,n,\zeta},\rho_{t,n,\zeta}$ are defined in section 3 (see also section 4).
-Notice that in (48) only the quasiparticles with $n=1$ contribute, reflecting that all the bound states are confined in $A$, and cannot contribute to the von Neumann entropy.
-Now, the $tDMRG$ data in Fig.
-11 exhibit deviations from (48).
-This is expected because (48) holds in the hydrodynamic limit $t\to\infty$.
-Finally, we should mention that it is possible to characterize the trajectories [16] of the the generic entangled pairs produced in $A$, which are responsible of the entanglement growth.
-This, in principle, could allow to determine also the secondary slopes visible in Fig.
-11.
-8 Conclusions
-We investigated bound-state confinement after trap-expansion experiments in the $XXZ$ spin chain at $\Delta>1$.
-We showed that after the expansion in the vacuum of quite generic initial states the bound states of the $XXZ$ chain remain confined in the initial region, provided that $\Delta$ is large enough.
-Upon lowering $\Delta$ the bound states are allowed to leave region $A$, i.e., they are not confined anymore.
-The condition ensuring confinement is the presence of a finite density of magnon-like excitations in the bulk of the initial region.
-Physically, bound-state confinement is a genuine consequence of interactions, i.e., the scattering, between the bound states and the magnons.
-Interestingly, if the size of the initial region is finite, at long times the density of magnons vanishes, as they leave the initial region, and bound-state confinement breaks down.
-Thus, for the expansion of finite regions, bound-state confinement is present only at intermediate times $t={\mathcal{O}}(\ell)$.
-We investigated bound-state confinement after the expansion of the $p$-Néel states (see Fig.
-1).
-In the limit $p\to\infty$ one recovers the domain wall initial state, and transport is suppressed.
-We showed that for small $p\lesssim 4$, bound-state confinement happens for the bound states of size $n>p$.
-At larger $p$, even the bound states with $n=p$ are confined, which is consistent with the absence of transport in the limit $p\to\infty$.
-Finally, we showed that bound-state confinement is reflected in a multi-stage linear growth of the entanglement entropy.
-Let us now discuss some directions for future research.
-First, we investigated bound-state confinement in a lattice model.
-It would be interesting to understand whether confinement occurs for systems in the continuum, such as the attractive Lieb-Liniger gas.
-This would pave the way for the experimental verification of bound-state confinement, for instance with atom-chip experiments [35, 36] or trapped ions experiments [8].
-Interestingly, bound-state confinement could be potentially used to prepare thermodynamic states with high bound-state content.
-Specifically, the protocol to enrich the bound state content would consists of two steps.
-In a first step one can expand the initial state as in Fig.
-1 (a).
-Then, after the magnon-like excitation are “evaporated” one could trap the remaining excitations by applying an external potential.
-As the bound states are typically associated with lower thermodynamic entropy, bound-state confinement can also be used to prepare low-entropy states.
-It has been pointed out recently that bound-states could be robust agains long-range interactions [37, 8].
-It would be interesting to understand whether long-range spin chains exhibit bound-state confinement as well.
-Finally, it would be important to clarify whether bound-state confinement is also present in integrable quantum circuits [38, 39, 40].
-Acknowledgements
-This study was carried out within the National Centre on HPC, Big Data and Quantum Computing - SPOKE 10 (Quantum Computing) and received funding from the European Union Next-GenerationEU - National Recovery and Resilience Plan (NRRP) – MISSION 4 COMPONENT 2, INVESTMENT N.
-1.4 – CUP N.
-I53C22000690001.
-This work has been supported by the project “Artificially devised many-body quantum dynamics in low dimensions - ManyQLowD” funded by the MIUR Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN) Bando 2022 - grant 2022R35ZBF.
-This work has been partially funded by the ERC Starting Grant 101042293 (HEPIQ).
+11 for $\zeta<0$, the finite-$\chi$ effects are visible, decreasing with increasing $\chi$.
+In particular, already at $\chi=800$ the convergence is satisfactory, and even at $\chi=200$ the qualitative behavior of the data is correctly captured.
+Finally, for $\zeta>0$ finite-$\chi$ effects are significantly smaller.
+This is expected because entanglement, which limits the efficiency of $tDMRG$, is produced in region $A$, whereas it is simply “transported” [19] outside of $A$.
+This is the reason why with $\chi\approx 1000$, we can reach $t\approx 100$ in our simulations.
+Figure 11: Convergence of the $tDMRG$ for the full-expansion dynamics.
+We plot the projector $P_{\uparrow\uparrow}$ as a function of $\zeta=x/t$.
+We focus on the $XXZ$ chain with $\Delta=1$ and the expansion of the Néel state.
+The initial region $A$ (see Fig.
+1) contains $\ell=40$ sites.
+The different symbols corresponds to different values of the bond dimension $\chi$.
+All the results are for $t=16$.